A Complete Bibliography of Publications in Annals of Statistics (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2023 Version 1.05

Title word cross-reference

+t^{4} [GJW01a]. 2^{l−n} [Zhu03]. 2^{n−m} [CC09]. ^{2} [Joh04]. α [ACD09a]. β [BCV01]. c [DHL09]. C[0, 1] [EL06, dHL03]. χ [Joh04]. D [DF00, DF01, FH08, Pro00, MN09, ST07]. d + 2 [ST07]. D_{1} [DF00, DF01]. E [IS01]. E(s^{2}) [BC04]. ℓ_{1} [CP09, MN09]. ϵ [Zha06]. F [HC08, NWJ09]. Γ [NS01]. I [Bha06]. ∞ [RZ06]. K [LQS00, BW07]. L_{1} [Zha09, Gre06]. L_{2} [Nai04]. L_{k} [KL05]. M [AFR00, BMT07, CR01, FYZ01, JLW02, KT01b, Koz03, Ren03, RRY03, Sha00, TT00, Wu07]. L_{p} [Dur07]. Z^{+} [HS02]. MM [BMT07]. MTP_{2} [BF00]. n [Bic07, CL07a, CT07a, CT07b, EHT07, FS07, KM07, Rit07, SW07a]. p [Bic07, CL07a, CT07a, CT07b, EHT07, FS07, KM07, Rit07, SY00, SY03]. q [Sto03]. R [HOP06]. S [Ho06, TT00]. t [DSW09, JSZ04]. U [CLV08, GM07, JW03, SW04]. Z [Nis09].

-divergences [NWJ09]. -efficiencies [DF01]. -entropy [Zha06]. -error [Dur07, KL05]. -Estimates [FYZ01, KT01b, AFR00, BMT07, RRY03]. -estimation [CR01, HOP06, JLW02, Wu07]. -estimators [Koz03, Nis09, Ren03, Sha00]. -fitting [Nai04]. -functionals [DSW09, TT00]. -minimax [NS01]. -mixing [BCV01]. -monotone [MN09, BW07]. -norm [MN09]. -optimal [BC04, DF00, DHL09, FH08, IS01]. -optimum [Pro00]. -paths [Ho06]. -projections [Bha06]. -stable [ACD09a]. -statistic [GM07, JSZ04]. -statistics [CLV08, JW03, SW04]. -test [HC08]. -value [Sto03]. -values [SY00, SY03].

28 [IKL00a].

34 [Fuh10, Jen10].

aberration [CC09, CY04, CT05, WX01a, WX01b, WW02, Xu06]. absolute [BDT01, ZRY09]. Accelerated [DHM07, Dip03, NKY09]. according [Pon03]. Accumulated [Ing07]. accuracy [CH02, HPT02]. activation [ZY08a]. active [WW07b]. activity [ASJ09a]. acyclic [GH02, MW09, VR09]. AdaBoost [Jia04]. Adaptation [TvdG05, CL04a, CL05a, Lec07, SV09]. Adapting [ABDJ06]. Adaptive [ACDH06, BHR02, BKMP09a, BCV01, BHL03, BG03a, BGA07, CL06a, CW08, Efr08, FL06, GW08, GT01, HJPS01, KT01c, LM00, PT00, Pro00, RvdV06, Roh08, SSN09, Spo00, vdVvZ09, CL05c, CL06b, CZ09a, CY09, Dal05, DGMR07, DL03, GBS09, HM03, HZ04, HZH09, Hua04, LBM09, LM03, MF09, PS03, VvS08, VW01, Zha05b, ZHCC07, ZZ09b]. Addendum [DG06]. Additive [FHT00, HM04, JvdM09, LNV03, MP05a, MP06, MS00a, MvdGB09, NBPW04, Sch01, WY07, YPM08, ZW03]. additivity [SSG05]. adequacy [HP04b]. adjusted [CGLZ09, ŞM06a, ZHCC07]. adjusting [Zen04]. admissibility [BST05]. Admissible [BGX08]. Affinely [RS06b]. after [Lah05]. against [CH09, LT05]. aggregating [Gol09]. Aggregation [BTW07, Aud09, BS07, JN00, Tsy04]. aided [KP03]. algebra [GMS06]. algorithm [BD08, MP07c, NS01]. algorithmic [GG08]. Algorithms [Kün05, BCDD08, GKM06, HM08, Hun04, HL05b, HN07, RSD07, Zhu04]. aliasing [Zhu03]. alignment [PSS09]. alignments [SY00, SY03]. all-pass [ADB07, BDT01]. Allocation [YZ02]. almost [LR06, LR08]. alternative [Ste00, Zha08b]. among [MR06]. AMP [LPM03, AP06, LMP01]. Analysis [GEGMMI08, Gel05, HC08, LC05, RSD07, Zha06, AW09, And02, BRT09, Bri02b, BK01, CB06, DMM03, DN01, DMP05, DR06b, Dry05, GR05, GZH09, GX00, HMW06, HL07, Hös03, Joh01, Joh08, KL03a, KN04, Kün05, LR02, LR03, LP05, MT04, MR01, Nad08, Owe09, RGN02, Sie03, Spe02, Ste00, TW08, VW01, Wal01, WM07, YMW05, ZC05a, ZH05, ZILZ07]. Analyzing [BY02]. Ancestral [RS02, ARS09]. angle [EHJT04]. angles [Pan09]. angular [Mar06]. anisotropic [HL02]. Ann [Fuh10, Jen10]. Annals [IKL00a]. annealed [FF08]. ANOVA [GKK^{+}00, HKST00, Lin00, MZ06]. Antithetic [FGR00, CM05, Owe08]. application [CWTG08, CR08, Cho04, EdHL06, Jam05a, LWH^{+}07, LQS00, MS06, SSSW03, Ye03, ZH05, BHL05]. applications [BM04, BS07, BG01, BC03, CY09, CT05, DMM03, DM08, IK06, KM07, KZW06b, KY05, KT01d, Li02, Lia09, Nis09, Spo09, TW08, Wal01, Reb05]. Applied [Bic08, LP05]. Approach [HJPS01, AR09b, CZ09b, CL09, CFRC07, DS02, FHS09, GEGMMI08, GW04a, GZH09, GX00, HM07b, KK09, Kol00, LZC05, LF09b, Nad08, ND03, PS03, SV09, SM06b, Xia07, Zen05]. approaches [BG09, Wal04]. Approximate [SY00, SY03, DGMS00]. Approximately [Shi04, NSY08]. Approximating [HY05a, RSV04]. Approximation [BCDD08, BW04, Car07, CLL08, Dah00, Fri01, JSZ04, JR04, Jon00, Lai03, MP07c, Nad08]. approximations [BW02, CL00, EdHL06, HS07, Kol03, RRY03]. arch [HT01, DR06a, FSR08, RZ06]. Archimedean [MN09]. areas [CFRC07]. argument [BW02]. arguments [EHJL08]. ARMA [LM03, MPY09]. armed [YZ02]. array [Loh08]. arrays [FPW01, ST07]. arrivals [JT09]. article [Fuh10, Jen10]. aspects [TV07, Zha00]. Assessing [HV06, Owe00]. associated [Zuo03]. association [Osi09]. assumption [Cat04]. asymmetrical [WX01a, WX01b]. Asymptotic [AC04, AW05, BHR02, BSK05, BW09, BCLZ02, BC07, BS09, CZ09a, Car07, CL00, DJ06, DMR04, DH02, DGM06, EL06, Fuh04, GQR07, GCLN02, GN06, GJ09, GJW01b, GK08, HVW06, HL07, Hös03, HZ04, HHM08, KL05, Lah07, Mar03, NKY09, Nis09, Osi09, Rei08, SW07b, Wan02, ZC05a, ZHCC07, ZZ09a, ACH04, AH06, AS00, BCD02, DZM09, HS07, IKL00a, IKL00b, JLW07, Lah03, LRZ03, Loh03, MM07, ZSW22]. Asymptotically [Asl06, Bar07, FH00, GG00, FDR09, SKF09, ST01]. Asymptotics [BKMP09b, BPP09, FDR07, FK00, LZ07a, LS03, MF09, Rei03, XY03, AGL05, BG01, BDN02, Hua03, KM07, Kou04, Loh05, Mar06, Rad08, SBZ04, Sha00]. Attributing [HO04]. augmentation [HM08]. authority [HM09]. authors [FHT00, HL02]. autocorrelations [MS00b]. autoregression [BCV01, GKLT09, GN06, WY07]. Autoregressive [KP03, AH03a, And02, ACD09a, Bel00, BBLR05, DMR04, GZ01, Ing04, IW05, Ing07, MPR05, MSW06]. Autoregressive-aided [KP03]. auxiliary [CHT08]. available [HY07]. average [Cha09, LZ07a]. averaged [FMS04]. averages [Jam05a]. averaging [JRT08].

backfitted [WY07]. backfitting [MP05a, MP06, YPM08]. background [ACDH06]. backward [CM05, Zhu04]. bagging [BY02]. Bahadur [CSW06, CSW07, Wu05]. balanced [BD04, HY03b, ZFLS05]. balls [Bar04, CL06a]. Banach [KP05a]. band [MR01]. bandit [YZ02]. bands [CvK03, GW08, SLM00, ZW08]. Bandwidth [HK05, MP05a, EM05, vdVvZ09]. bandwidths [HK01]. base [Jia02]. based [ADB07, Ban07, BC09, CWY09, DM01, FH08, GS09, GHR04, GX00, HP04a, HP02b, HP04b, HP06, HOP06, HP08, KMO03, KS07, LL08a, MS08, Pen04, RW02, RRY03, Roh08, RBK05, WZ07, Zha04, Zha08b, ZH08b, Zuo03, ZCY04, ZH06, Zuo06, vdVvZ08]. bases [Ber04]. basic [CT06]. Bayes [Asl06, BG09, CS05a, CL06c, Efr03, GvdV01, Ho06, JZ09, JS04, JS05, LM07a, Lia00, LV04, MS05a, TS06, Zha03, Zha05b]. Bayes-risk [LV04]. Bayesian [AC04, AGP07, ACH04, AK05, BG03a, BP04, BD07, CGMM09, CR08, Cho04, CY06, DGMS00, DR06b, GJ09, GD04, HW09, IR05, Jam05a, Jer05, Jia07, Joh04, KL03a, KL03b, KvdV06, Kom06, Lia09, PR08a, Pen06, Scr06, Ste00, Sto03, Wal04, Yua09, Zha00, vdVvZ09]. Bayesian-motivated [ACH04]. be [BR03]. behavior [AW05, KL06b, Zha04]. Behrens [BD08]. Benjamin [Mei06a]. Benjamini [FZ06]. Berkson [KS09, Wan04]. Bernoulli [GKK^{+}00]. Bernstein [Gho01, KL04, Kim06]. Best [Gre06, JLW02]. Beta [Jam05b, Sri03]. better [ACD09b]. between [BHKS06, Car02]. beyond [BDN02]. Bias [BZ01, BM04, BMT07, CT02, CHR00, CKS02, HP02a, ZH08a, ZCY04]. biased [AW05, Efr04, Gil00, HZ04]. BIC [CS00]. bilateral [BLBL02]. binary [BF00, CL06c, GR06, LC05]. binomial [BCD02, RW01]. bispectrum [Mar06]. bivariate [BS08, EdHL06, GS09]. black [Men08]. Blackwell [RSV02]. blending [FH08]. block [CZ09b, Lah07]. block-variables [Lah07]. Blocked [CM01, CM02, Xu06]. blocking [BM00]. blocks [MR07]. Blum [MP07c]. blurred [GZ06]. bodies [GKM06]. Boosting [Büh06, ZY05, BBB^{+}04, Bre04, FHT00, Fri01, Jia02, KY04, LV04, RSD07]. Bootstrap [CvK03, CB05, CLL08, DM01, HM09, HL05a, JP03, KL03b, KP03, Lah05, Nor09, Shi04]. Bootstrapping [HK01, SBZ02]. bound [Kar05, NS09]. boundaries [GZ06, HP02a, KL06b]. bounded [MP01]. Bounding [KP02, KP05b]. Bounds [CR08, Dre01, Tsa04, FMS04, GZ01, MN06, MM07, NEW04, Zha02, Zha05a, Zha09]. boxcar [JR04]. Bradley [Hun04]. brain [ZY08a]. branching [MT05]. Break [AHHR09]. Breakdown [DG05, DG06, Hen04, SBY08]. bridge [HHM08, Kou04]. Brownian [CD01, GJW01a]. bubbles [CWTG08]. Building [LL09b]. Bump [HMZ04]. burn [JH04]. burn-in [JH04].

calculating [NS01]. calculus [Jam05a, Jam06, Miz02, PR08b, TV07]. calibrating [HH00]. Cam [Ano02b, LZ07b, Yan02, vdV02]. Can [LP06, BR03]. Canonical [And02, GJW01a, ZH08b]. capacity [CM01, CM02]. CAR [Cat04]. carcinogenesis [LWH^{+}07]. cares [Men08]. Carlo [AR09b, BPR09, Cho04, CM05, DM08, FM03, Kün05, SG07]. cascades [OW00]. Case [GR01, GZH09, CR01, Che00, CSX09, PS09, dHP06]. Case-control [GZH09, Che00]. cases [PS04]. cat [Men08]. categorical [BR02, CCL02, Jae05, LCC03]. Causal [GR01, BW09, Lok08]. causation [VR09]. censored [AW05, PR06, Ren03, Ren08, SBY08, VJ03, WF09]. censoring [KL04, KS07, VJ03, Zen04, Zen05]. Central [CWTG08, Cho04, AGL05, Loh08]. certain [CY01, JM04]. Chain [CM05, LMP01, AP06, BB09, EHJL08, HS02, LPM03]. chains [DR06b, FGR00]. Change [LBM09, Fuh04, GTZ06, HWZ07, KS07, Lin07, Mei06a, Mei06b, Mou08, Pon03, Spo09]. Change-point [LBM09, GTZ06, HWZ07, KS07, Mei06b, Pon03]. changepoint [LP05]. changes [BHKS06]. characteristic [CWTG08, Efr08]. characteristics [Fuh04]. Characterization [CS05a, CY01, DHL09, GH02, GG08]. Characterizations [GJW01b]. Characterizing [AP06]. chart [HT04]. charts [WH06a]. checking [KS09]. checks [SZ05]. Cheng [Fuh10, Jen10]. Chernoff [Kar05, NS09]. Choice [HPS08, HK05]. chosen [HK01]. Cifarelli [Jam05b]. class [AS00, CW04, CFH07, CHC03, DMP04, FH08, GY02, HM07d, LZ06a, LPS08, LP10, MS05a]. classes [AP06, ML07]. classification [BS07, CHC03, FF08, HK05, HPS08, HI01, Jia02, JT08, KP05b, Lec07, TvdG05, Zha04]. Classifiers [KP02, AT07, FMS04, Tsy04]. Classifying [BO08, bCH09]. Clear [WW02]. Clinical [BHR02, HZ04]. Closed [AS08a, CY06]. Closed-form [AS08a]. CLT [AZ08]. cluster [GR05, GEGMMI08, Rob09, XY03, ZFS09]. clustering [Cha09, SBY09, vLBB08]. coarsened [DHM07]. coarsening [GG08]. codes [PX09]. coefficient [ALL05, HJS01, HC08, ŞM06a, WZZ09, ZL09]. coefficients [AH03a, CFZZ07, DM03, IK06, Kim07]. coherent [GW04b]. cohort [AW05]. coin [HZ04]. coincide [BLBL02]. cointegrating [CH06b]. cointegration [KMT07]. colon [LWH^{+}07]. color [MF09]. combinations [KP05b]. Combined [KP02]. Combining [SXS05]. Comments [Mei06a]. Common [BHK09, JT09, STT03]. compact [Jup05, Jup08]. companion [MP07c]. compared [LW02]. comparing [HY05b, Yan07]. comparison [CS05b, HT04, HM08, ND03, WW07a]. comparisons [BB02]. compartment [FH08]. competing [BM04, GMW08a, GMW08b]. complement [LZ07b]. complementary [XC08]. Complete [ST07]. Completeness [LMP01, LPM03, LC04]. Complex [GR01, GW04b, GLM03]. Complexities [KP05b, BBM05, BM06a, BM06c, CLV06, Kol06a, Kol06b, SW06, Tsy06, vdG06]. Complexity [LW04, BC09, JMP01, KN04, Lec07, Zha06]. Component [LZ06b, BMV06, CB06, HZ03, HV06, HMW06, Nad08]. components [AW09, BHK09, CHS09, HY03a, Joh01, Nob04, PP09, Ste00]. composed [FH08]. Composite [ZY08b, JLT09, ZRY09]. Compound [Zha03, BG09]. computational [BC09]. computations [AR09b]. computed [HPT02]. Computer [BWB^{+}07, MP07a, Loh08]. Computer-intensive [MP07a]. computerized [CY09]. Concave [CD01, BRW09, JvdM09]. concrete [CHC03]. condition [EdHL06, Lah03]. Conditional [Efr07a, Lee09, WH06a, CL02, HY05a, LP06, LCC03, SS00, Son09, TK03]. conditionally [SM06b]. Conditions [HS02, JW03, JSZ04, NS01]. cones [CV03]. Confidence [BM07, Bar04, BCD02, GW05, PQ06, SLM00, ZW08, CL04a, CL06a, CJL07, CvK03, DKM09, DG00, GW08, HL05a, Pen04, PT00, RvdV06, SXS05, Tsa04]. conjugate [CV03, MLD09]. connection [BW07]. Consistencies [BKL^{+}09]. Consistency [CGMM09, CL06c, GMW08a, Lah05, PP09, SA09, TMG09, Yan07, vLBB08, CS00, EM05, GR06, Jia04, JM09, KL01, LM07a, LV04, Sko00, Wal04, Zha04, ZY05, dHL03]. Consistent [AC09, BM04, Bun04, CT06, JMP01, AS00, Kar08a, SW07a]. Constrained [DF00, KT01b, BMT07, CKS02, Dur07, Sha00, ZH08b]. constraint [AS00, Gre06]. Constraints [Chi07, FH00, HH01, ST07]. constructed [PX09]. constructing [CC06]. Construction [BC04, QAW09, XW05, RBD09, LL08a]. constructive [Xia07]. contamination [IK03]. context [JM09]. contingency [BK01, Lan04]. Continuous [GR01, GK01, Lok08, Mou04, SGM09]. Contour [LZC05, LWT09]. contours [SZ00b]. contraction [vdVvZ08]. contrast [IK03, RBD09]. contributions [BB02, Bri02a, Coo04, Hub02, Spe02]. Control [RW07, VW01, BY01, CFH07, Che00, Chi07, GBS09, GW04a, GZH09, HT04, HY05b, RS06a, Wu08]. controlling [ABDJ06, Sar07]. Convergence [DGMR07, FM03, GKM06, GGvdV00, Gho01, GvdV07a, Hua04, Jia07, Scr06, ZY05, BKL^{+}09, BS09, DHM07, FPW01, GCL01, GW00, GvdV01, GvdV07b, GMW08a, HH07, HWZ07, JLW07, Joh08, LF09a, OGV09, SZ00b, SW01, WLP07]. convergent [BD08]. Convex [GJW01a, GJW01b, Pat01, BHL05, Bha06, GKM06, GZ06, Hof03, JN09, KP05b, Kol09, Zha04]. convexity [JN02]. convolution [SBY09]. copula [CWY09, EdHL06]. copula-based [CWY09]. copulas [GS09, MN09, OGV09]. corner [HM07c]. corner-preserving [HM07c]. Correction [CL10, CM02, DG12, FCZ02, IKL00a, LP10, LPM03, LR03, SY03, ZSW22, HP02a]. Corrections [BJYZ09]. Correlated [PA05]. correlation [And02, LL00, LWH^{+}07, SRB07, Zha08b, ZH08b]. correlations [AGL05]. corrupted [JvdM09]. cosmology [LQS00]. count [BZ09, WZ00, WZ07]. counting [WZ00]. counts [WW00]. coupling [CM05, FGR00]. course [TS06]. cousins [MRY07]. Covariance [BL08a, AZ08, AHHR09, BJYZ09, BL08b, DN01, Kar08a, Kar08b, LF09a, LW02, RMC08, STT03, SZT07]. covariances [JM04]. Covariate [CGLZ09, Bun04, DGM06, KS07, Pon03, ŞM06a, ZHCC07]. Covariate-adjusted [CGLZ09, ZHCC07]. covariates [WZ07, YZ02, Zen04, ZL09]. coverage [HPT02, Tsa04, ZZ09a]. Cox [DGM06, FL02, NEW04, Pon03]. Cramér [GQR07, HRW07]. Cramér-type [HRW07]. Cramér-von [GQR07]. criteria [Efr01, HN07, Ing07, Kou04]. criticism [DJ04, HJ08]. Cross [Leu05, WZ05, Yan07]. Cross-validation [Leu05, WZ05]. crossover [HY03b, HY05b, SBR05]. Current [GMW08a, GMW08b, VJ03, MK05]. curve [CM00a, FDR09, PT00]. curved [FM03]. curvelet [CD02]. Curves [BZ01, CvK03, HM03, KSC07, ND03]. CUSUM [Fuh03, HT04, Mou04].

Dantzig [Bic07, BRT09, CL07a, CT07a, CT07b, EHT07, FS07, MRY07, Rit07]. Data [BGA07, CHR00, GR01, GL05, Jup08, SBY09, ASJ09a, AVYZ09, AW05, BZ09, BSK05, BKMP09a, BR02, CFZZ07, CZ09b, bCH09, CHT08, DJ06, Efr04, FLZ06, GKK^{+}00, GH07, GHR04, GMW08a, GMW08b, GX00, HPT02, HMW06, HY07, HL07, HM08, Jae05, JT08, JvdM09, Kol00, KM07, LWH^{+}07, LL08a, LC05, MK05, MKB09, MS00a, MY09, NEW04, NKY09, PR06, RW02, Ren03, Ren08, SBY08, SSSW03, SG07, TS06, VJ03, WM07, WZ00, WZ07, YMW05, ZC07, ZFS09, ZCH04]. Data-driven [GL05, Jup08, CZ09b]. Decision [CS05b, AC04, BM07, BG09, FS02a, GD04, NS01, RSV02, Zha03]. decomposable [LM07b]. Decomposition [BB09, BG01]. Decompounding [BG03b]. Deconvolution [Joh09, DHM08, HM07b, HL08, JR04, Kol00, LS00, PS09]. deconvolutions [KM03]. decreasing [KL06b]. deep [Miz02]. Default [Pau05]. Deficiency [Car02]. defining [RBD09]. definition [BBS09]. deformations [AS08b, CM03]. deformed [AC09]. degree [ASJ09a, DHP00]. degrees [MW00, ZHT07]. delay [Bei00]. delayed [BHR02]. denoising [PS03]. Densities [CD01, DK04, Asl06, CWTG08, Efr08, EL00, GLX06, GvdV01, GvdV07b, GM07, Jia07, LPS08, LP10, RR05]. Density [Efr04, AFR00, BW07, BRW09, BKMP09a, BCLZ04, BGX08, Cav00, CH02, DW08, Efr05, Efr07a, FL06, GPPVW09, Gho01, GHR01, HK01, HO04, KL06b, Nai04, PKJ02, SW07a, Scr06, SSN09, WM02, Yan00, Zha06]. dependence [BHKS06, BC07, BDN02, CH02, CH09, HJ08, Lah03, Lah07, NL06, SRB07, Wu08]. Dependency [FDR07, BY01]. dependent [BC02, BMW03, CSX09, CSW06, CSW07, DGM06, HI01, KR09, LRZ03, SW07a, Wu05, Wu07, Zen04, Zen05]. Depth [ZC05b, LL08a, Miz02, SZ00a, SZ00b, Zuo03, ZCY04, ZCH04, ZH06, Zuo06]. depth-based [ZH06]. depth-weighted [ZCH04]. derivatives [CvK03, HY07, HMY09]. descriptors [DMP05]. design [BM00, BCLZ02, GK08, HM03, KR09, LM07a, Pro00, Rei08, RBK05, XC08, Zhu03]. design-adaptive [HM03]. design-based [RBK05]. design-consistency [LM07a]. designs [BB01, BHR02, BD04, BM06b, BG05, BD07, BC03, BC04, CY09, Che00, CC06, CC09, CM01, CM02, CY04, DF00, DF01, DM03, DMP04, DMP05, DHL09, DT09, DHP00, FH08, HY03b, HY05b, HZ04, HZH09, Imh01, IS01, KM00, LZ06a, Loh08, MS08, MF09, MR07, PX09, QAW09, RBD09, SBR05, ST01, VW01, WX01a, WX01b, WW07b, XW05, Xu06, YS09, Ye03, ZFLS05, ZHCC07]. detecting [DJ04]. Detection [Cha09, AH03a, ACDH06, AHHR09, Bei00, Fuh04, LP05, Mei06a, Mei06b, Mou08, Spo09, ZY08a]. determination [HC08, KM00]. Determining [CL04c]. deviation [BDT01]. deviations [HRW07, IK03, SSG05]. diagnostics [Jia01]. difference [BL07]. Differentiability [DSW09]. differentiation [Ols07]. diffusion [BPR09, CGT08, GCLN02, KSC07, Spo00]. diffusions [ASM04, AS08a, CHS09, Dal05, GHR04, MZ06, Nis09, Wan02]. Dimension [CL02, HJPS01, LCC03, LD09, ZH08b, CW04, CL04c, Coo04, FBJ09, GG00, HY05a, HR09, JM09, LW02, LZC05, LY07, LWT09, Xia07]. dimension-reduction [HR09]. dimensional [AW09, BJYZ09, BG03a, BG09, Büh06, bCH09, DMP05, Dry05, FF08, Gre06, HHM08, Jia07, JT08, Kar08a, Kar08b, KvdV06, MKB09, MvdGB09, MB06, MY09, PBHT08, Scr06, WLP07, WR09, XH09, ZH08a, vdG08]. dimensions [CCL02, Owe00]. Diophantine [JR04]. Direct [HJS01]. directed [GH02, MW09, VR09]. directional [BKMP09a, GKM06]. Directions [MP05b, Xia07]. Dirichlet [CM00b, GvdV07b, Jam05b, RGN02]. discontinuities [HM03]. Discovery [Sar02, ABDJ06, BY01, DJ06, Efr07b, FDR07, FDR09, GW04a, Sar06, SG09, Sto03, Wu08]. discrepancy [GD04]. discrete [ASM04, GH07, Lah03, MLD09, RR05]. discretely [ASJ07, ASJ09b, BPR09, JT09]. discriminant [RS06b]. discriminating [BHKS06]. discrimination [CHC03, DT09, MT04]. Discussion [AL06, BM06a, Bic07, BM06c, BM08, CL07a, CR06, CK06, CLV06, EHT07, FS07, MRY07, Men08, ML06, Per06, Rit07, SBZ06, SW06, Tho06, Tsy06, WZ06, Zha08a, vdG06, DG06, FHT00, HL02]. Discussions [BBB^{+}04]. dispersion [CAMMI08]. distance [Car02, KS09]. distances [LMR^{+}08, SRB07]. distributed [LD09]. Distribution [DHV05, KK09, AS00, BW07, BRW09, BG03a, CLL08, CK08a, EPdH01, ES09, GLM03, GMW08b, HY05a, Joh09, Joh01, JvdM09, KL06a, LP06, Loh03, Mei06a, Mei06b, NSY08, Nob04, Pen04, PQ06, Rob09]. Distributional [RLP03, BG01]. Distributions [KP02, BM04, EHJL08, GH02, GGvdV00, GvdV07a, HZ03, HL08, Hua04, HWH07, Jup05, LM07b, MN09, RGN02, RR05, RS06b, SW01, SXS05, Sri03, TT00, TMG09, WLP07, ZH06, vdVvZ08]. divergences [JW07, NWJ09]. diverging [FP04, LF08, MP07a, ZZ09b]. do [HN07, JP03]. Does [ACD09b]. domain [DZM09, Loh05, NL06]. Donoho [ZCH04]. Doubling [CC06, XC08]. doubly [HZ04, Ren03]. down [GBS09]. drift [BCLZ04, Dal05, GCLN02, PR08b, Spo00]. drifted [Bel00]. driven [CZ09b, GL05, Jup08, NBPW04]. dynamic [MS00a]. dynamical [SA09].

early [ZY05]. EBLUP [CLL08]. edge [TvdG05]. edges [ACD09b, CD02]. Edgeworth [GR03, JW03]. Effect [CH02, HPT02, WBCL08, GM05, Ste02]. effects [BD04, DS02, JH04, Lok08, MKB09, SZT07, WW07b, ZZ06]. efficiencies [DF01]. Efficiency [Rob05, BH04, CHT08, DGT06, GQR07, IKL00a, IKL00b, Kar05, RW00]. Efficient [ALL05, AH03a, Cav00, CB06, CWY09, Dah06, DL03, Fuh06, Fuh10, HZH09, Jen10, KP05a, MSW06, AR09b, BM06b, FH00, GY02, HP06, HOP06, HY05b]. eigen [RMSE08]. eigen-inference [RMSE08]. Eigenspaces [SBY09]. eigenvalue [Joh01, Joh08]. eigenvalues [SMT08]. eigenvectors [SMT08]. elastic [ZZ09b]. elastic-net [ZZ09b]. ellipsoidally [RS06b]. elliptic [HP04b]. elliptical [BZ01, CAMMI08]. Emax [FH08]. Empirical [CLV08, HT01, JS04, JS05, Kol00, KP02, Pen04, RW02, WC09, BG09, CL08, CL10, DJR04, Efr03, ES09, FZ04, FHS09, GCL01, GJ09, HV06, HM09, HMV09, JLW02, JZ09, Lia00, MP04, ND03, NL06, Ren08, Sch07, TS06, Tsa04, Zha03, Zha05b]. Empirical-likelihood-based [Pen04]. empirically [HK01]. Endogenous [BO08]. endpoint [CS05a]. endpoints [HP02a]. energy [AL06, CK06, KZW06b, ML06, WZ06]. Enriched [CV03]. ensembles [Bre04, Joh08]. Entropies [GvdV01]. entropy [GD04, Kol09, SDG06, Zha06]. enumeration [ST07]. equations [BSK05, CB05, JLW07, LC05, PP00, SM06b, WC09, XY03]. Equi [AL06, CK06, KZW06b, ML06, WZ06]. Equi-energy [AL06, CK06, KZW06b, ML06, WZ06]. Equivalence [BCLZ04, ARS09, AP06, BCLZ02, CZ09a, GCLN02, GN06, GJ09, Rei08, Rom05]. ergodic [Dal05, Nis09, SA09]. Error [KP02, BS09, CL00, CSW06, CSW07, DJR04, Dur07, HM06, HR08, Joh09, KP05b, KL06a, KL05, LR05, LY07, RS06a, RW07, VW01, ZL09]. error-prone [ZL09]. Errors [Tau01, DH02, Efr05, FR02, GK08, HM07a, Ing07, KK09, KS09, KR09, LM03, Wan04, Wu07]. Errors-in-Variables [Tau01, HM07a]. estimate [KMO03, LP06, TMG09]. Estimates [FYZ01, KT01b, AY02, AFR00, AC09, BS07, BMT07, BHZ06, BBLR05, BM08, DG00, Gil00, JS04, Men08, RRY03, SBZ02, SBZ04, Zha08a, ZL08a, ZL08b]. Estimating [ASJ09a, AS08b, CM03, JvdM09, LL00, LNV03, LQS00, MKB09, ML07, MR06, Mül09, PSS09, SW04, Zen04, BSK05, Can03, CB05, CKS02, DM03, GCLN02, JLW07, LC05, Mer07, WC09, XY03]. Estimation [BW07, BS09, BGA07, CJL07, CW04, Cha07, Efr05, Fuh10, GLT06, GH07, GT01, GJW01a, GJW01b, HL08, HMY09, JM04, Jen10, KT01c, KMS00, LR06, LR08, LSV08, Pon03, Reb05, SZT07, Tau01, Wan04, Zha05a, ALL05, AR09a, ADB07, ACD09a, AK05, BRW09, BKMP09a, BCV01, BM04, Bel00, BPR09, Bic07, BL08b, BMV06, BDT01, BCLZ04, BL07, BGX08, BCZ08, BG09, BG03b, But07, CL03, CL04b, CL05c, CL05a, CL06b, CL07a, CW08, CZ09a, CZ09b, CT07a, CT07b, Cav00, CR01, CM00a, Che04, CH06b, CZ07, CWY09, CM01, CM02, Coe08, CT06, CAMMI08, CFRC07, Dah06, DP06, Dal05, Efr04, Efr07a, Efr08, EHT07, EL00, EPdH01, ES09, FLZ06, FS07, FSR08, Fuh06, GK01, Gho01, GvdV01, GM07, GHR01, GR03, GHR04, GTZ06, GS08, HZ03, HM03, HP04a]. estimation [HM06, HQ07, HY07, HOP06, Hid05, HS02, Hof03, HR08, HM04, HM07d, Hös03, HJS01, Hua04, JMP01, Jer05, JLW02, JZ09, JLT09, JN09, KT01a, KMT07, Kar08a, Kar08b, KP05a, KN04, KSC07, LF09a, LBM09, LM00, LZ07a, LWH^{+}07, LM03, LP05, MK05, MM09, MP07a, MM07, MS06, MPR05, MPY09, Nai04, NL04, OGV09, OW00, PKJ02, PS04, PT00, PP00, PR08b, RMC08, Rit07, RZ06, Rob08, RR05, Sch07, Scr06, SP05, SSN09, Spo00, SM06b, TvdG05, VvS08, WBCL08, WZ07, WM02, Wu07, Xia07, Yan00, YZ02, ZW03, Zha05b, Zha06, ZH05, ZW09, vdVvZ09]. estimations [CH02]. estimator [BW09, CS00, DMR04, GZ04, KL05, Lah05, LRZ03, MRT08, ZZ09a, ZCH04]. Estimators [ASM04, ASJ07, AS00, BC09, Ber04, BH04, BR03, BKL^{+}09, BO00, BC07, CL05b, CL06c, CKS09, DZM09, Dur07, EM05, EL06, FK00, GY02, Gol09, GLdW03, HK01, HWZ07, Hen04, HHM08, IK03, KM03, Kim04, Koz03, LM07a, LP06, LPS08, LP10, MS05a, Nis09, PR06, PP09, Ren03, SW07a, Sha00, SZ01, WZ00, ZC05a, ZCY04, ZC05b, dHL03]. Euler [CWTG08]. Evaluation [EHJL08]. events [DMM03]. ever [Gel05]. evidence [ACCHZ08]. evolution [SA09]. evolutionary [VvS08]. EWMA [HT04]. Exact [SP05, RGN02]. exactly [Zha05b]. example [Li02]. Examples [Zha05a]. Existence [RBD09]. expansion [JW03]. expansions [AS08a, BCD02, GR03, Lah07, LRZ03]. expectation [FM03]. expected [FR02]. experimental [Pro00]. experiments [BB09, Car02, Car07, LWH^{+}07, Loh08, MS08]. Explicit [IK06]. exponent [Coe08]. exponential [Bei00, CL00, DJ06, FM03, GHKM01, Imh01, Scr06]. exponentially [Sch07]. expressions [CY06]. Extended [Hel06]. Extending [HMV09]. extrapolation [HPT02]. extrema [HV06]. Extremal [Che05]. Extreme [Dre01, EPdH01, EL06, EdHL06, ES09, GS09, GLdW03, Rob09, dHL03]. extreme-value [ES09, GS09]. Extremes [DK01, GS08, MS00b, dHP06]. extrinsic [BP03, BP05].

factor [DS02, WW02]. factorial [BC03, CM01, CM02, CY04, DS02, PX09, RBD09, WX01a, WX01b, Xu06, Ye03]. factors [BM00, CY04, Zhu03]. failure [CFZZ07, NKY09]. False [Sar02, Sar06, ABDJ06, BY01, DJ06, Efr07b, FDR07, FDR09, GW04a, MR06, SG09, Sto03, Wu08]. families [CL00, FPW01, FM03, GHKM01, Scr06]. family [CV03, ZRY09]. familywise [LR05, RS06a]. Fast [AT07, Aud09, SS07]. fault [HR00]. FDR [Chi07, GBS09, Sar07]. feature [PBHT08]. features [FF08]. Fence [JRGN08]. field [GPPVW09, LL00, vdVvZ09]. fields [AS08b, AC09, CWTG08, CW04, CT06, KSC07, LB06, Loh05, Mar06, NSY08, SSSW03, TW08]. filamentary [ACDH06]. filling [QAW09]. filtering [ACD09b, ZH08b]. filters [Kün05]. Financial [Myk03]. Finite [Kim04, Nad08, RW00, BD08, BG01, HRW07, IK06, Kou04, LM07a, Nob04]. finite-sample [BD08]. first [Cha07]. Fisher [BD08, Lua04]. fit [ACH04, BRS06, BM06b, But07, DHV05, FL06, JW07, Jan00, Jia01, Joh04, Jup05, KK04, KK09, OGV09, Roh08]. fitted [BO08, Jia07]. Fitting [KL06a, Nai04, RW01]. Fixed [DZM09, Loh05, CL00, PA05]. Fixed-domain [DZM09, Loh05]. Flexible [RMC08]. fMRI [SSSW03, ZY08a]. forecasting [Ing07, SA09]. form [AS08a, CY06]. formal [BBS09, EHJL08]. forms [KT01d]. formulation [HR09]. forward [CM05, Zhu04]. foundation [LMR^{+}08]. four [DHP00]. Fourier [DM03, Lah03, MM09]. fractal [CW04]. fraction [PX09]. fractional [BBLR05, BC03, CH06b, CM01, CM02, DS02, Rob05, SP05, TV07, WX01a, WX01b, Xu06]. Frailty [FL02, GZH09, KLF04]. frames [CD02]. framework [RBK05, ZFLS05]. free [DHV05, KK09]. freedom [MW00, ZHT07]. frequency [ASJ09a, GHR04, NL06]. Frequentist [IR05, Pen06, ACH04, CK08a, CK08b, DGMS00, Yua09]. Fuh [Jen10]. full [GZH09]. fully [GY02]. Function [BGA07, GJW01b, ACH04, AK05, AW05, BHL05, CW08, CT02, CL06c, Dal05, EM05, Fri01, HQ07, HM04, JvdM09, LQS00, PR06, Reb05, RSD07, SZ00a, SZ01, WBCL08, Ye03, Zen04, Zhu03, ZCY04]. Functional [HKST00, JWZ09, JN00, PS09, YMW05, BWB^{+}07, BHK09, But07, CL05b, CL06b, CH06a, CKS09, HMW06, HH07, HMY09, LM00, MS05b, ZC07]. Functionality [AH05]. Functionals [Jam05b, KT01c, CL03, CL04b, CL05c, DSW09, Mül09, RGN02, TT00]. Functions [BW01, GJW01a, BMT07, BW02, BW04, Can03, Efr08, FZ04, GM07, HH00, HY05a, HV06, HM07d, Jan00, JLT09, LWH^{+}07, LNV03, Mar03, MN09, Mer07, NWJ09, SZ00b, Zuo03]. fused [Rin09]. FWER [Sar07].

GACV [GKK^{+}00]. Game [GD04]. Gamma [Jam05b, vdVvZ09, BLBL02]. GARCH [BH04, KY05, LM03, MS00b, MS06, Wan02]. Gaussian [AS08b, AC09, AH03b, AH05, BLBL02, Bar04, BHL05, BGH09, BCLZ04, BTW07, CW04, Coe08, DP06, GW00, GR06, GHR01, HMZ04, Jia05, LL00, LRZ03, LB04, Loh05, MP04, MRT08, NSY08, Pau05, Pen06, PR08b, RMC08, SMT08, SS07, vdVvZ08, vdVvZ09]. General [CK08a, JZ09, SZ00a, Zha05b, ASM04, AH06, Bha06, BK01, CT05, GEGMMI08, GZH09, HM07d, LZC05]. Generalization [FMS04, KP02, KP05b]. Generalizations [LR05, RS06a]. Generalized [CB05, FZZ01, FCZ02, MS05b, WX01a, WX01b, ZZ06, BSK05, BHZ06, HT04, Hun04, Jia07, MS05a, NS01, RW07, Sar07, SG09, Sch01, SLM00, XY03, YPM08, vdG08]. Generalizing [Sar08]. generated [MF09]. generating [BW04]. genetic [CKS02]. Geometric [CY04, BB01, DHL09, DS02, GG08, HS02, Kou04]. Geometry [SS00, Kol00]. Gibbs [FGR00, JT08, JH04, PR08a]. glasses [Cha07]. Global [Jan00, AC04]. globally [BD08]. GLR [BLBL02, HT04]. GMM [CHT08]. Goodness [But07, JW07, Jia01, KK09, BRS06, DHV05, FL06, Jan00, Joh04, Jup05, KK04, OGV09, Roh08]. Goodness-of-fit [But07, JW07, Jia01, KK09, DHV05, FL06, Joh04, KK04, OGV09, Roh08]. gradient [Fri01, GPPVW09]. Graph [LMP01, AP06, LPM03, RS02]. Graphical [GHKM01, BC02, GH02, GMS06, Mad00, RMC08]. graphics [FS02b]. graphs [ARS09, LM07b, MW09, MB06, VR09]. Greedy [Fri01, BCDD08, Jon00, LW08]. Grenander [KL05]. group [CR01, GLM03, Li02]. grouped [ZRY09]. groups [DG05, DG06, Zhu03]. growth [Imh01, WH06a]. guide [FWF09]. Gumbel [PSS09].

half [DH02]. half-spaces [DH02]. harmonic [DMP05]. harmonics [MP04]. harmonizable [LR02, LR03]. Hausdorff [SSN09]. haystacks [JS04]. Hazard [CFZZ07, HV05, Ho06, KL03a, Kim06, LNV03]. Hazards [FL02, BPP09, Che04, CZ07, HKST00, KL03b, KLF04, SGM09, Zen05]. heavy [Pen04, PQ06]. heavy-tailed [Pen04]. Herbert [Ano03, Sie03]. Hessian [CL04c]. heterogeneous [DJ04]. heteroscedastic [CW08, DHL09, GK08, Imh01, KL06a, KR09, SM06b]. Hidden [BC03, AR09a, Fuh03, Fuh04, SS00]. Hierarchical [BG05, BST05, JH04, LM07a, MLD09, PR08a, ZRY09]. High [AW09, FF08, KY05, Mar06, MvdGB09, MB06, SBY08, WR09, vdG08, ASJ09a, AGL05, BG09, Büh06, bCH09, Dry05, Gre06, HHM08, Jia07, JT08, JM09, MKB09, MY09, Owe00, PBHT08, PQ06, XH09, ZH08a]. High-dimensional [AW09, FF08, MvdGB09, MB06, WR09, vdG08, BG09, Büh06, bCH09, Dry05, Gre06, HHM08, JT08, MKB09, MY09, PBHT08, XH09, ZH08a]. high-order [AGL05]. High-resolution [Mar06]. Higher [CK08b, DJ04, HS07, HJ08]. Higher-order [HS07]. Hill [DRdH00]. Hochberg [FZ06, Sar08]. Homodyne [BGA07]. homogeneity [HP08, STT03, ZZ06]. homogeneous [Lan04]. Hough [GZ04]. hull [CG06]. hulls [Kol09]. hunting [HMZ04]. Hurst [Coe08, GH07]. Hybrid [Ber04, Yua09]. hypergeometric [BW02]. hyperpriors [BST05]. Hypotheses [DS01, Jia02, BHL03, BHL05, Bar07, BRS06, FR02, KSC07, MR06, Rom05]. Hypothesis [CL09, Kar05, LW02, NS09].

i.i.d [Ren03]. ideal [CP09]. idealism [AH03b]. ideals [GW04b]. Identifiability [AMR09, LS03, SKF09]. Identification [MT05, CR08]. identifying [WW07b]. identity [Jam05b]. if [Men08]. Ignorability [Jae05, LC04]. II [BM04, BP05, PP00]. II. [HOP06]. ill [CD02, CG06]. ill-posed [CD02, CG06]. Image [PS03, DMM03]. images [HM07c]. imaging [KSC07]. implications [CHS09, Jia02]. importance [CDS06, DGMR07, PSS09]. important [Gel05]. Improved [GLX06, OGV09]. improvements [Rob05]. Improving [Lia09, MP01]. imputation [Kim04, RW02]. inadmissibility [CS05a]. incidental [Mor09]. incomplete [Nau06]. increasing [HV05, NBPW04]. increments [RLP03]. independence [FF08, GBS09, GQR07, Lah03, SS00, Son09]. independencies [MW09]. independent [CB06, Kom04, RLP03, SXS05]. independently [MR06]. index [GLdW03, HJS01, SZ05]. Indicator [Ye03]. indirect [But07, GTZ06]. individual [DM03]. induction [BK01]. inequalities [BM06a, BM06c, CGPT02, CLV06, Efr08, Kol06a, Kol06b, SW06, Tsy06, vdG06]. inequality [CP04]. Inference [GR01, HY03a, HWH07, KS07, Rob09, SMT08, ŞM06a, WF09, AH06, Aud09, Ban07, BG03a, Bun04, CK08b, Cho04, DR06a, DW08, GS09, GK08, HH05, HP06, HOP06, Jia05, KLF04, KZW06b, LF08, Lee09, MS04, Osi09, RW02, RMSE08, Rei03, Rob05, RW00, RBK05, SG07, YZ04, Yua09, ZL09]. inferences [ZC07]. infinite [BG03a, KvdV06, Scr06, WLP07]. infinite-dimensional [BG03a, KvdV06, Scr06, WLP07]. Infinitely [GT01]. Influence [ZCY04, CT02, ZILZ07]. Information [NEW04, Hös03, Ing07, Jia05, LPS08, LP10, Lua04, SXS05, Zha05a, Zha06]. ingredients [DHP00]. inhomogeneous [MS04]. Institute [Bic08]. instrumental [HH05]. Integral [KSC07]. integrals [Ols07]. Integrated [GJW01a]. integration [SP05]. intensity [GCLN02]. intensive [MP07a]. interactions [DS02, WW02]. interactive [FS02b]. interdirections [HP02b]. interesting [Li02]. interference [KM00]. interpretable [JWZ09]. interpretation [Sto03]. intersections [Bha06]. interval [Pen04, PT00]. Intervals [BCD02, CL04a, CLL08, HPT02, HL05a, Myk03]. intervention [MKB09]. intrinsic [BP03, BP05]. introduction [KY04]. invariance [CSW06, CSW07, Li02, LY07]. Invariant [Jer05, Mad00, RS06b, SW04]. Invelope [GJW01a]. inverse [AB06, BLBL02, CD02, CGPT02, CG06, HR08, HR09, KM03, vdVvZ09]. inverse-Gaussian [BLBL02]. invertible [SW07a]. invited [Lai03, Sie03, Zha03]. isomorphism [CY04]. isotonic [Zha02]. isotropic [AS08b, AC09]. Itô [MZ06]. item [CY09]. Iterated [HL05a]. Iterative [JLW07, Bha06, CL04c]. IV [CC06, CC09].

jackknife [JLW02, Lah05]. jackknife-after-bootstrap [Lah05]. jackknifing [Mar03]. Jacobi [Joh08]. John [Ano02a, BB02, Bri02a, Bri02b, Dem02, FS02b, Hub02, Spe02]. joint [RBK05, ZC05a]. jump [BKL^{+}09, Ste00]. jump-penalized [BKL^{+}09]. jumps [ASJ09a, ASJ09b, JT09].

Karl [Owe09]. Kernel [FBJ09, GLdW03, HSS08, WM02, EM05, HH01, KSC07, LF08, LB04, LNV03, OGV09, WY07]. Kernel-type [GLdW03, EM05]. kernels [HMZ04, SS07]. Kiefer [CSW06, CSW07, DHP00, MP07c, WW07a]. KL [Zha06]. KL-entropy [Zha06]. Kleinberg [CHC03]. Kou [AL06, CK06, ML06, WZ06]. Kriging [Ste02]. Kullback [GLX06].

L2Boosting [MRY07]. lack [BM06b]. Laplace [BW02]. Large [BP03, BP05, BMW03, Gil00, BJYZ09, BC09, BL08b, GKK^{+}00, HRW07, Kar08a, Kar08b, KM07, LF09a, LW02, MR06, MR07, RMSE08, RW00, XY03]. large-dimensional [BJYZ09, Kar08a]. larger [Bic07, CL07a, CT07a, CT07b, EHT07, FS07, Rit07]. Largest [Joh08, Joh01]. Lasso [MB06, MY09, FK00, Rin09, ZHT07, vdG08, BRT09, MRY07, ZH08a]. Lasso-type [MY09, FK00]. latent [AMR09, CY01, PA05]. lattice [ZFLS05]. law [LY07]. laws [SW04]. layout [Ber04]. Learning [JRT08, AT07, Aud09, BCDD08, Gre06, HSS08, MN06, Tsy04]. Least [BDT01, CD01, EHJT04, SZ01, BKL^{+}09, BC07, FSR08, LF09b, Zha09]. least-squares [FSR08]. Leibler [GLX06]. length [AW05, CKS02]. length-biased [AW05]. lengths [CFRC07]. Level [ST07, CC06, SSN09, XW05, Ye03]. Lévy [ASJ07, Jam05a]. life [Bri02a, KMS00]. lifetime [FLZ06]. Likelihood [BW01, Ban07, Drt09, FZZ01, Fuh10, Fuk03, Jen10, Zen05, AS08a, ACD09a, AGL05, BRW09, BW09, BF00, BS07, BPR09, BM08, Cha09, CZ07, CAMMI08, Dah00, DP06, DGT06, DG00, DMR04, DZM09, EL00, ES09, FCZ02, FP04, FZ04, FLZ06, FWF09, FHS09, Fuh06, GvdV01, Gil00, GZH09, GJ09, HS07, Hen04, HMV09, JZ09, KN04, LF08, LRZ03, LC05, LS03, LC04, MK05, Men08, Mor09, NL06, PKJ02, PP09, Pen04, RW02, Ren08, RZ06, Sch07, SM06b, SG07, Tsa04, Wal01, WZ05, WC09, WZ07, ZC05a, Zha08a, ZL08a, ZL08b]. likelihood-based [RW02, WZ07]. Limit [BW07, BRW09, CWTG08, DM08, MS00b, AGL05, Cho04, Loh08, PR06]. Limiting [GMW08b, Lin07, Rob09, ZH06]. limits [Joh08, MS06]. line [HR00, VNG09]. Linear [GT01, JLW07, KT01c, ALL05, ACD09b, BHL03, BHZ06, BK01, Büh06, CL03, CL04b, CL05c, CH06a, CLL08, CKS09, DHV05, GCL01, HH00, HH07, HLT04, HR08, JWZ09, JM04, Jia05, Jia07, KM03, LT05, MK05, MLD09, MS05b, MSW06, Mül09, RZ07, SW04, SW07a, SLM00, VNG09, WZZ09, WM02, Wu07, XH09, YM09, YMW05, ZH08a, ZW09, ZL09, vdG08]. linearity [Owe00]. link [Hel06, HM04, HM07d]. linkage [Hös03]. LM [BD08]. Local [AGL05, BBM05, BM06a, BM06c, BO00, CZ07, CLV06, DK01, FLZ06, FWF09, HLT04, Hua03, Jon00, Kol06a, Kol06b, Owe08, PS04, SW06, Tsy06, ZW09, vdG06, BS07, GQR07, GM07, HC08, Nai04, PKJ02, PSS09, Rob08, Sha00, SP05, SY00, SY03, Spo09, ZILZ07]. localized [LW04]. Locally [FH08, ST01, VvS08, Coe08, Dah00, DP06, YS09]. Location [FYZ01, AY02, CR01, CAMMI08, DSW09, HP02b, Hen04, SBZ04]. loci [GX00]. log [BRW09, BK01, MK05, MLD09]. log-concave [BRW09]. log-likelihood [MK05]. log-linear [BK01, MLD09]. logistic [AGL05, CY09, FHT00, ST01]. long [BHKS06, CL08, CL10, CHS09, CSW06, CSW07, GR03, GK08, Lah03, NL06]. long-range [BHKS06, CSW06, CSW07, Lah03, NL06]. long-run [CHS09]. Longitudinal [GR01, BSK05, HMW06, LWH^{+}07, MS00a, SZT07, WF09, YMW05]. loss [GLX06, LS03, NWJ09, SKF09]. low [GHR04, HY03a, JM09, WW00]. lower [NS09]. LR [BD08]. LRT [BJYZ09]. Lucien [Ano02b, Yan02, vdV02].

MA [LT05]. machine [Fri01, HSS08]. machines [BBM08, SS07, SA09]. made [BRS06]. magnitude [GG00]. Mahalanobis [HP02b]. main [DS02]. Majorant [CD01]. Majorization [ZFLS05]. make [DRdH00]. Malliavin [PR08b]. Manifest [CY01]. manifolds [BP03, BP05, CR01, HL07, Jer05, Jup05, Jup08]. Many [GT01, AMR09, BM00, GCL01, Zen04]. MAP [GJ09]. mapping [GX00]. Margin [KP02, Lec07, RSD07, TvdG05]. Marginal [BR02, CD01, KM07, AR09b, HM08, Zen04, Zen05]. marked [MP07b]. marker [Hös03, VJ03]. Markov [LPM03, ARS09, AP06, AR09a, CWY09, CM05, CS00, CT06, DR06b, DMR04, EHJL08, Fuh03, Fuh04, HS02, LMP01, Mad00, RS02, RSV02]. marks [SGM09]. Martingale [KK04, Mer07, MS06]. matching [RS06b]. Matérn [Loh05]. Matérn-type [Loh05]. Mathematical [Bic08]. matrices [AZ08, BL08b, DMO08, DG12, Kar08a, Kar08b, LL00, RMSE08, STT03, SMT08]. Matrix [MW09, BJYZ09, Bic08, BW02, Kar08b, KMO03, LF09a, LW02, Nad08, SZT07]. maxima [KT01d, LL09a, NSY08]. Maximal [DMM03]. Maximin [Imh01, BD07]. maximization [FM03]. maximizing [DF01]. Maximum [ACD09a, BZ01, EL00, ES09, Lua04, BRW09, BW09, BMT07, BPR09, CT02, CM01, CM02, DP06, DGT06, DG00, DMR04, DZM09, GvdV01, Gil00, GD04, Hen04, JZ09, LRZ03, Mor09, PP09, RZ06, SM06b, Wal01, ZC05a, ZCY04]. maze [ACCHZ08]. MCMC [BC09]. MCMC-based [BC09]. Mean [DJR04, BHL05, BG03a, BHKS06, BS09, CL02, HH00, HM06, KMS00, LCC03, Lia00, MP01, Mar03, Pen04, STT03, WBCL08, WZ00]. mean-squared [HM06]. meaningful [DMM03]. means [AGP07, BP03, BP05, BG09, JZ09, RLP03, ZCH04]. measure [EPdH01, ES09]. measurement [KS09, LY07, Wan04]. measurements [DHM08, ZC05a]. measures [Bha06, CH02, GKM06, NBPW04, RLP03, ZILZ07]. Measuring [SRB07]. mechanics [Hel06, KZW06b]. median [ACD09b, BCZ08, CT02, YZ04]. medians [Zuo03]. meet [HN07]. memory [CL08, CL10, GR03, GK08, MRT08]. meta [Owe09]. meta-analysis [Owe09]. method [BL07, CG06, CC06, CSX09, FZ06, GR05, Ho06, LB04, MM09, MP06, Mor09]. Methodology [HH07]. methods [AR09a, bCH09, CHR00, Cho04, DM08, HP02a, HM03, HH05, HMW06, HM07c, HSS08, JRGN08, LZ06a, Lia09, LNV03, LV04, RS06b, Ste00, WZ07, Zha03, Zha04, Zha05b]. microarray [KM07, TS06]. microarrays [Efr03]. Minimal [VR09, JW03]. Minimax [BGA07, CL04b, Dre01, ZH05, Asl06, GLX06, GG00, MS05a, NS01, Zha05b]. minimaxity [DJ06]. Minimization [CLV08, BM06a, BM06c, CP09, CLV06, Kol06a, Kol06b, SW06, Tsy06, Zha04, vdG06]. Minimum [KS09, CC09, CY04, CT05, DF01, GD04, IK03, WX01a, WX01b, WW02, Xu06, Zha06]. mining [JT08]. minors [DMO08, DG12]. mirror [JRT08]. Mises [GQR07, KL04, Kim06]. Missing [LC04, DGM06, Mül09, NEW04, NKY09, RW02, SG07, WC09]. Misspecification [KvdV06, Pat01]. Mixed [Rad08, CLL08, GM05, Jia01, Jia05, JRGN08, Lua04, MS08, ZZ06]. mixed-effect [GM05]. Mixed-rates [Rad08]. Mixing [Yan00, BCV01, BDN02, RR05, TMG09]. mixture [BMV06, CL09, DHP00, GW00, HZ03, JMP01, Nob04, Ste00]. mixtures [CJL07, CR08, DJ04, DGMR07, GvdV01, GvdV07b, Hen04, Hof03, HWH07, RL05, RS06b]. MLE [GMW08a, GMW08b, MP01]. MLS [Pat01]. MM [Hun04, HL05b]. modality [DK04]. mode [SZ01]. Model [FL02, Tau01, Weg03, ALL05, BCV01, BHL03, BGH09, BB04, BWB^{+}07, BC02, BP04, BS08, BMV06, BO08, Bun04, CP09, Cha06, Che04, CGT08, CAMMI08, DGM06, FL06, GHKM01, GG00, GZH09, HP04b, HM04, HJS01, Jia01, JRGN08, JH04, KMT07, KL04, Kim06, KS09, KM00, LM00, LP06, Lee09, Lia09, LT05, LSV08, LC04, LF09b, MS00a, McC02, NKY09, Pon03, RBK05, RW01, Sch01, Spo00, Wal01, WW00, WY07, ZFS09, ZY08b]. Modeling [Li02, GZ01, Hel06, HKST00, LL08b, Lok08, MvdGB09]. Models [Fuh10, Jen10, LMP01, LM03, Pat01, dHP06, AH03a, AMR09, And02, AP06, AR09a, ADB07, AHHR09, Ban07, BST05, BR02, BHZ06, BDT01, BMW03, BK01, Büh06, BM08, BS09, CFZZ07, CL07b, CR01, CY09, CLL08, CHT08, CWY09, CL09, CKS02, DM03, DMP04, DMR04, DHP00, FH08, FSR08, Fuh03, Fuh04, Fuh06, Fuk03, GK01, GKK^{+}00, GHKM01, GH02, GMS06, GCL01, Gil00, GM05, GL05, GK08, HY03a, HM06, HM07a, HS07, Hof03, HM07d, HHM08, Hun04, Imh01, IS01, Jia05, Jia07, KK04, KL01, KL03a, KL03b, KP05a, KLF04, KS07, KL06a, KY05, LZ06a, LL00, Lan04, LPM03, Lin00, LT05, Lin07, LL09b, MK05, Mad00, MS08, MP05a, MP06, MLD09, Men08, MS04, MPY09, MS05b, MSW06, NEW04, Osi09, PR08a, RMC08, Ren08, RS02, RZ06, ŞM06a, Sko00, SKF09]. models [Ste00, SZ05, SLM00, SGM09, SG07, WLP07, Wan02, Wan04, WF09, WZZ09, Wu07, XH09, YS09, YPM08, ZW03, Zha08a, ZL09, ZZ06, ZL08a, ZL08b, vdG08]. Moderate [IK03]. modified [HS07]. moment [BW04, JSZ04, KY05, NS01, TK03]. Moments [DMO08, DG12, FPW01, HL08, SS00]. Monomial [GW04b]. Monotone [BW01, BW07, Ban07, EL00, HV05, Ho06, KMS00, MN09]. monotonicity [Dur07, GSvdV00, HH00, HH01, JN02]. Monte [AR09b, BPR09, Cho04, CM05, DM08, FM03, Kün05, SG07]. Motion [CD01, GJW01a]. motivated [ACH04]. moving [Jam05a]. much [Bic07, CL07a, CT07a, CT07b, EHT07, FS07, Rit07]. multi [BZ09, XW05, YZ02]. multi-armed [YZ02]. multi-level [XW05]. multi-sample [BZ09]. Multidimensional [Zuo06, BW09, JT09, LPS08, LP10]. multilayer [Fuk03]. multilinear [KT01d]. Multinomial [Lan04, Car02, LL09b]. Multinomial-Poisson [Lan04]. Multiple [FR02, GX00, Rob08, BY01, BB02, CH09, CS05a, CS05b, CSX09, FS02a, HP04a, Kim04, LRS05, RW07, Sar02, Sar06, Zhu03]. multiplicative [LNV03, OW00]. multipoint [Hös03]. Multiprocess [CM05]. Multiresolution [DK01]. Multiscale [DS01, DW08, KN04, Spo09, Wal01, ACDH06, Shi04]. multistage [Bar07]. multistep [Ing04, Shi04]. multistep-multiscale [Shi04]. multitreatment [HZ04]. multitype [MT05]. Multivariate [AK05, Joh08, Kol03, LL08a, MN09, AY02, AS08a, AFR00, AVYZ09, AHHR09, CFZZ07, Car02, CHS09, GQR07, HZ03, HQ07, HP02b, Kol00, LZ06b, Loh08, MM09, RL05, Rei08, RRY03, STT03, Sri03, TS06, TW08, YM09, ZCH04, ZH06]. multiway [CDS06].

Narrow [MR01]. Narrow-band [MR01]. Near [CP09, KL06b]. Near-ideal [CP09]. Nearest [HI01, bCH09, HPS08]. nearest-neighbor [bCH09, HPS08]. nearly [BC07]. necessarily [AH03b]. necessary [Lah03]. NED [Lin07]. Needles [JS04]. needlets [BKMP09a, BKMP09b]. NEFs [BLBL02]. neighbor [BD04, bCH09, HPS08, HI01]. neighbor-balanced [BD04]. neighborhood [CT06]. neighborhoods [CL05a]. nested [HM06, QAW09]. nested-error [HM06]. net [Loh03, ZZ09b]. nets [Owe08]. network [Jon00]. networks [Fuk03]. neural [CL07b, Fuk03, Jon00]. neutral [Jam06]. no [JSZ04]. noise [AH03b, AH05, BCLZ04, JvdM09, Pen06, SA09, WW00]. Noisy [BGA07, GH07, GZ06, HM07c, KSC07]. Non [BZ01, AH03b, CW04, HMZ04, Jia05, Pen06, Ren03]. Non-elliptical [BZ01]. non-Gaussian [CW04, HMZ04, Jia05, Pen06]. non-i.i.d [Ren03]. non-necessarily [AH03b]. Nonanticipating [LP05]. Nonasymptotic [GZ01]. Nonconcave [FP04, BM08, Men08, Zha08a, ZL08a, ZL08b]. nonconvex [CL04b]. nondiscovery [Sar06]. nonelliptical [TT00]. nonelliptically [LD09]. nonequivalence [Wan02]. noniid [GvdV07a]. Nonlinear [CY09, CHS09, HR08, Tau01, CGLZ09, DMP04, GKLT09, Jon00, KMT07, MSW06, Mül09, SW07b, Sko00, Wan04, WY07, YS09, ZSW22]. nonmanipulable [OS09]. nonnecessarily [AH05]. nonnormal [PA05]. Nonparametric [BR03, BG09, CL05a, DP06, DKM09, DN01, EPdH01, GHR04, HH01, HZ03, HH05, HM06, HQ07, HY07, Hof03, HM04, JLT09, JN09, KT01a, KMT07, KSC07, LWH^{+}07, ND03, PP00, RR05, SSG05, SZ05, BZ09, BCLZ02, BL07, BCZ08, BEM01, BDN02, CL03, CL04a, CW08, CZ09a, Car07, CHR00, CFRC07, DHM07, DM01, DL03, FZ04, GW05, GR06, GN06, GM05, HK01, HK05, HWZ07, HM07d, JN00, JN02, KK09, KL04, LW08, Leu05, LB06, LL08a, LZ06b, MM09, MS00a, Rei08, RvdV06, RW00, Spo00, WBCL08, Weg03, YM09, YZ02, YZ04, ZZ09a, Zha00, ZW08]. nonparametrics [HW09]. Nonquadratic [CL05b]. nonregular [BC03]. nonresponse [FHS09]. nonserial [HVW06]. nonsmooth [Mar03]. nonstationary [GKLT09, HWZ07, LM03, MR01, MRT08, PS04, Rob05, ZW09]. Nonsubjective [SDG06]. norm [Kar08a, MN09]. Normal [BLBL02, AGP07, BG03a, BST05, BS08, BG09, CJL07, Car02, CL09, GvdV01, JZ09, Koz03, Lia00, Mad00, MP01, RL05, STT03]. normality [BW09, DH02, EL06, HVW06, HS07, KL05, ZZ09a]. Normalized [FSR08, NBPW04, RLP03]. Note [IKL00a, Bei00, CL03, Mei06a, Nor09, ZSW22]. notions [SZ00a]. NPMLE [AW05, KL06b]. null [KT01a, MR06]. number [BD07, FP04, FR02, LF08, ML07, MR06, Nob04, Ste00, ZZ09b]. numerical [RGN02].

Object [WM07]. Objective [BS08, BP04]. observables [Kom04]. observation [MZ02, RSV02]. observational [MKB09]. observations [ASM04, But07, GKK^{+}00, GvdV07a, GTZ06, GZ06, RSV04, SA09]. observed [ASJ09b, AMR09, BPR09, JT09, Jia05]. observing [VJ03]. obtain [Bha06]. On-line [VNG09]. One [BM08, Men08, Zha08a, ZL08a, ZL08b, Ber04, CS05b, FH08, LP06]. one-compartment [FH08]. one-sided [CS05b]. One-step [BM08, Men08, Zha08a, ZL08a, ZL08b]. one-way [Ber04]. only [BLBL02]. onto [Bha06]. operating [Fuh04]. Operator [Kar08a, HR08]. operators [SBY09]. Optimal [BM00, BB04, BM06b, CL06b, CFH07, Che00, DM03, DMP04, DMP05, DT09, DS02, FYZ01, GTZ06, GM05, HP02b, HOP06, HP08, HWZ07, HY05b, MS08, MM07, Rom05, Tsy04, Bar07, BD07, BC04, DF00, DHL09, FH08, FDR09, Fuh04, GL05, HP04b, HT04, HM07d, IS01, Ing04, Ing07, KM00, Mei06a, NL04, SKF09, ST01, XW05, YS09, HP06]. Optimality [BB01, BD04, Mou04, AGP07, CD02, HY03b, LRS05, Pen06, RSV02, SBR05]. optimization [Dip03, Gre06, Pro00]. optimum [Pro00]. options [Myk03]. Oracle [CGPT02, BM06a, BM06c, CLV06, Efr08, GS08, Kol06a, Kol06b, SW06, Tsy06, ZY08b, vdG06]. Order [IW05, AH06, AGL05, Cha06, CR08, CK08b, CS00, DGT06, FH00, HPS08, HS07, Ing04, STT03]. ordered [MS08]. ordering [AS00, DHP00]. ordinal [Ber04]. oriented [WM07]. Orthogonal [BG01, Loh08, ST07, VW01, WW07b]. other [CKS02]. outcome [BMW03]. Outlier [HM07c]. outliers [AVYZ09, Leu05]. output [BWB^{+}07].

pair [GX00]. panel [BZ09, WZ00, WZ07]. paper [Lai03, Sie03, Zha03]. papers [BBB^{+}04, KY04]. paradigm [KM07]. parallel [CM05]. Parameter [BBLR05, GH02, SV09, Bel00, CL04b, Dah06, GK01, GH07, HM07b, LRZ03, Mor09, MRT08, PSS09]. Parameters [GT01, AMR09, BH04, CR01, FP04, GK01, GCL01, KP05a, LF08, Mei06b, MS06, YS09, ZZ09b]. Parametric [CLL08, LS00, Tau01, FWF09, GHR01, LL09b]. Pareto [HPT02]. Pareto-type [HPT02]. Part [CHC03]. partial [BB01, CZ07, Chi07, FLZ06, KY05, RSV02]. partial-likelihood [FLZ06]. Partially [Jia05, ALL05, BHZ06, WZZ09, XH09, ZL09]. partition [Jam05a]. partitioning [FS02a]. partly [MK05]. pass [ADB07, BDT01]. path [GPPVW09]. paths [Ho06, RZ07]. pattern [BG05, CHC03]. patterns [Zhu03]. Patterson [SBR05]. PCA [JM09]. Pearson [Owe09, Zha08b]. penalization [Kol09]. Penalized [DGT06, MK05, BKL^{+}09, BM08, FP04, KN04, Men08, XH09, Zha08a, ZL08a, ZL08b]. penalizing [SSG05]. penalties [LW04, ZRY09]. penalty [Bei00, Ber04, GG00, TvdG05]. perfect [Hös03]. performance [BBM08, Chi07, Zha09]. Periodic [JR04, LR06, LR08, LB04, PS09]. periodogram [KP03]. periodograms [LL09a]. permutation [JP03]. persistence [Gre06]. Perturbation [ZILZ07, Nad08]. phase [BMW03, RBK05]. Phenomenon [FZZ01, FCZ02]. phi [JW07]. phi-divergences [JW07]. philosopher [Dem02]. Piecewise [RZ07]. plane [AS08b, AC09]. plans [DS02]. plot [DRdH00]. plug [AT07]. plug-in [AT07]. plugged [BR03]. plugged-in [BR03]. plus [WW00]. Point [Sch07, Fuh04, GTZ06, HQ07, HWZ07, KS07, LBM09, LQS00, MP07b, Mei06b, Pon03, SBY08, Spo09]. point-spread [HQ07]. points [ACDH06, BM07, BD07, Hen04, Lin07, Miz02, YS09]. pointwise [PT00, PS03, SV09]. Poisson [AB06, BCLZ04, BG03b, GCLN02, Jam05a, Jam06, Kom04, Lan04, PS00]. pole [GHR01, Hid05]. polynomial [BO00, DF00, DF01, HM07a, Hua03, HC08, Imh01]. polynomials [Gho01]. Population [Bre04, BG01, HL05a, HRW07, LM07a]. posed [CD02, CG06]. positive [GCLN02, LS00, Sto03]. positivity [JN02]. positivity/monotonicity/convexity [JN02]. possibly [JS04]. post [BO08, Bun04, LP06, Lee09]. post-model-selection [LP06]. post-stratification [BO08]. Posterior [BST05, GR06, GvdV07b, BPP09, CK08a, DH02, EHJL08, GGvdV00, GvdV07a, Hua04, JT08, KL01, Nob04, SW01, WLP07, vdVvZ08]. potential [Li02]. power [ABK00, Efr07b, Jan00]. powerful [FS02a]. pre [Mei06b]. pre-change [Mei06b]. precedence [CFH07]. Preconditioning [PBHT08]. Prediction [CH06a, GT01, AR09a, CLL08, DGMS00, HPT02, HM06, Ing07, JLW02, Kom04, Kom06, MSW06, Myk03]. predictions [IW05]. predictive [Asl06, BB04, BGX08, GLX06, Lee09, SDG06, VNG09]. predictor [Coo04, DJR04, IK06]. predictors [CCL02, DHM07, Ing04, LCC03, LD09]. preferences [Nau06]. presence [HH05, Zen05]. presene [JM04]. present [Mei06b]. preserve [ACD09b]. preserving [HM07c]. pretest [ABK00]. prevalent [AW05]. principal [AW09, BHK09, CHS09, HV06, HMW06, Joh01, Nad08, PP09]. principle [FS02a]. principles [CSW06, CSW07]. Prior [BK01, GR06, MLD09]. priors [BS08, BBS09, CM00b, CV03, GH02, Kom06, MZ02, Pau05, SDG06, vdVvZ08]. probabilities [CL00, DH02, KT01d, LMR^{+}08, Tsa04]. probability [Bha06, CL07b, Efr08, GH02, HO04, HS02, Kol03]. problem [AGP07, BD08, BG03b, Cav00, DL03, FH00, HR09, KK09, Mor09, WW00, WW07a, YZ02]. problems [AB06, BM04, CD02, CGPT02, CG06, CFH07, CS05a, Fuh10, HQ07, HM07a, HL08, HR08, Jen10, KM03, KSC07, Lia09, PBHT08, PS00, Zha00]. procedure [Bha06, CS05a, GBS09, Gol09, Mou04, Sar08]. procedures [ABK00, CGMM09, CH09, CS05a, CS05b, GG00, LRS05, RS06a, Sar02, Sar06, Sar07, Yan07]. Process [GJW01a, Jia04, ASJ09b, GCL01, GW04a, GR06, HT01, Jam05a, LRZ03, LQS00, MP04, MS00b, ND03, RGN02, RSV04, WZ00, vdVvZ08]. processes [ASJ07, ACD09a, BW09, BH04, BPR09, BBLR05, BCLZ04, BC07, CW04, CL08, CL10, CGT08, CM03, Coe08, CSW06, CSW07, Dah00, Dah06, DP06, DR06a, DHV05, EdHL06, GM07, Hid05, Ing04, IW05, JT09, Jam05b, Jam06, KY05, LR02, LR03, LR06, LR08, LL09a, MT05, MR01, MP07b, MPR05, MZ06, NBPW04, Nis09, Pau05, PS00, PR08b, RSV02, SW07a, WM02]. processesby [SW04]. Product [PR06, Lin00]. Product-limit [PR06]. professional [Bri02a]. Profile [LF08, CK08a, CK08b]. Profile-kernel [LF08]. program [Bic08]. programming [JN09]. projected [LWT09]. Projection [AY02, Zuo03, BC03, CC06, Pan09, Zhu04, ZCY04, Zuo06]. Projection-based [Zuo03]. projections [Bha06, CG06]. projective [DS02, MP05b]. prone [ZL09]. Propagation [AVYZ09, SV09]. Properties [BHR02, HMW06, HJ08, LMP01, Rin09, ACH04, BC03, CK08a, CY01, DMR04, DZM09, HZ04, HHM08, JLW07, Kim04, LPM03, LB04, SZ00b, ZHCC07]. property [LM07a]. proportion [BCD02, MR06]. Proportional [SGM09, Che04, CZ07, FL02, HKST00, JM04, KL03a, KL03b, Kim06, KLF04, RS06b, Zen05]. propriety [BST05]. proven [GBS09]. Pseudo [RZ06, AR09b, FHS09, GZH09, HP02b]. pseudo-Mahalanobis [HP02b]. pseudo-marginal [AR09b]. Pseudo-maximum [RZ06]. publications [Ano02a, Ano02b, Ano03]. pursuit [Zhu04]. PX [HM08]. pyramids [HW09].

QTL [GX00]. Quadratic [LMR^{+}08, But07, CL05b, CL06b, LM00]. quadrature [Loh03]. Qualitative [DS01, BHL05]. quantify [HM09]. Quantile [HW09, Kim07, WZZ09, Che05, WF09, ZW09, ZY08b]. quantiles [Coe08, HL08, HL05a, Koz03, Lah05, PQ06, Wu05]. quantitative [CY04, GX00]. quantities [CKS02]. Quantum [BGA07, Hel06, Kar05, Lua04, NS09]. Quarter [PX09]. Quarter-fraction [PX09]. Quasi [SM06b, BW09, DP06, FWF09, LC05]. quasi-likelihood [FWF09, LC05]. quasi-maximum [BW09, DP06]. Quasi-maximum-likelihood [SM06b]. quaternary [PX09]. Quotient [Zha08b].

Rademacher [BBM05, BM06a, BM06c, CLV06, Kol06a, Kol06b, SW06, Tsy06, vdG06]. Random [Bic08, HL02, TW08, AH03a, AS08b, AC09, ACDH06, BC02, BCLZ02, BG03b, BW04, CWTG08, CT06, GG08, JH04, Kar08b, KR09, LL00, LB06, Loh05, LC04, LW04, Mar06, Mül09, NSY08, NBPW04, Pan09, PSS09, RLP03, Rei08, SSSW03, SZT07, Zha05a, vdVvZ09]. randomization [RBD09]. randomizations [BB09]. Randomized [YZ02, Dip03, GKK^{+}00, HZH09, Loh08]. randomized-adaptive [HZH09]. randomly [ASM04, MF09]. range [BHKS06, CSW06, CSW07, Lah03, NL06]. Rank [ADB07, GS09, HP04b, And02, HP06, HOP06, HVW06, HP08, ZH06]. Rank-based [ADB07, GS09, HP04b, HP06, HOP06, HP08]. Ranking [CLV08, HM09]. ranks [HP02b, Roh08]. Rate [HM07d, Sar02, ABDJ06, BY01, BS09, DJ06, FDR07, FDR09, GL05, HV05, HWZ07, Ho06, LR05, MP07a, RS06a, SG09, Sto03]. Rate-optimal [HM07d, GL05]. Rates [GW00, SW01, vdVvZ08, AT07, Aud09, BKL^{+}09, Efr07b, GGvdV00, Gho01, GvdV01, GvdV07a, GvdV07b, GMW08a, HP04a, HH07, HL02, Hua04, Jia07, Joh08, LF09a, Rad08, RW07, Sar06, Scr06, SS07, VW01, WLP07]. Ratio [BW01, FZZ01, BF00, Cha09, Drt09, FCZ02, FZ04, Fuk03, HS07, LS03, Tsa04]. rational [IS01]. realization [CHC03, IW05]. recognition [BG05, CHC03]. reconstructing [GKM06, HM07c]. Recovering [CD02, GZ06]. recovery [Kol09, LF09b, MY09]. recurrence [HS02, SM06b]. recurrent [KT01a]. Recursive [Bel00, Kün05, Mad00, MPR05, TMG09]. Redescending [KT01b]. reduced [And02]. Reducing [CPW07]. Reduction [HJPS01, CCL02, CHR00, CL02, Coo04, FBJ09, HY05a, HR09, LCC03, LZC05, LY07, LD09, LWT09, Xia07, ZH08b]. reference [BBS09, CV03]. refinements [Rin09]. Regazzini [Jam05b]. regime [DMR04]. regions [BEM01, DKM09, DG00, LL08a, LL09b, PQ06, Shi04, Tsa04]. Regression [BZ01, GT01, Ren03, And02, BCV01, Bar04, BHL05, BMT07, BO00, BMW03, BCLZ02, BL07, BCZ08, BDN02, Bun04, BTW07, CH06a, CW08, CZ09a, Car07, CR01, CZ07, Che05, CHR00, CvK03, CL02, CL06c, CKS09, CGLZ09, DKM09, DHM07, DM01, DF00, DF01, DM03, DMP04, DHL09, Efr05, Efr07a, EHJT04, FHT00, FBJ09, GK01, GLT06, GW05, GY02, GSvdV00, GR06, GG00, GN06, GL05, GX00, GK08, HH00, HH01, HM03, HM06, HH07, HLT04, HL02, HM07d, HR09, HKST00, Hua03, HHM08, HC08, JWZ09, Jia02, Jon00, JN00, KK04, KK09, Kim07, KLF04, Kou04, KS09, KR09, LW08, LZ06a, Leu05, LZC05, LY07, LL08b, LZ06b, MS00a, MS05a, MW00, Mül09, NEW04, ND03, Ols07, PBHT08, Pon03, Rei08, SBZ02, SBY08, Sch01, ŞM06a, SZ01, Sko00]. regression [ST01, SSG05, VNG09, WBCL08, WF09, WZZ09, Weg03, XH09, Yan07, YMW05, Zen05, Zha02, ZW03, ZH08a, Zha09, ZW08, ZY08b]. regressions [CCL02, LCC03]. Regressors [BZ01]. regret [GG00, SDG06]. regular [CM01, CM02, HS07, RBD09, Xu06]. Regularity [KT01b]. regularization [BL08a, CG06, DKM09, LB04, LW04, Zha09]. Regularized [BL08b, AZ08, LV04, LF09b, RZ07]. reinforced [MF09]. rejection [FDR09]. rejective [AGL05]. Rejoinder [CT07b, Kol06b, KZW06a, WH06b, ZL08b, FHT00, HL02]. rejoinders [BBB^{+}04]. related [CLL08, ZH06]. relations [CFH07]. relative [DH02, IKL00a, IKL00b, SDG06]. relaxations [AW09]. reliability [GW04b, Reb05]. renewal [PP00]. renewal-type [PP00]. Rényi [LP10, LPS08]. repeated [DHM08, ZC05a]. replicated [TS06]. replication [HY03a]. Reply [Fuh10]. representation [HVW06, IK06, Wu05]. representations [MW09, MY09]. reproducing [LB04]. Resampling [LZ06a, LB06, Shi04]. Residual [CL08, CL10, KMS00]. residuals [GCL01, HT01, KY05]. resolution [CC06, CC09, Mar06]. Resolvable [MR07]. Response [BHR02, Ban07, CY09, HR00, MF09, RW02, ZHCC07]. response-adaptive [MF09, ZHCC07]. responses [Mül09]. restricted [MW00, PP09]. restriction [STT03]. restrictions [AH06, JM04, Reb05, TK03]. resulted [FH08]. Results [Sar02, BSK05, CL07b, CC09, CM00b, CS05b, Mar03, Nad08, RLP03, SZ00b, ZC05a]. reversible [DR06b, Ste00]. revisited [CP04, Mou08, Owe09]. ridge [Can03, HM07b, MS05a]. ridge-parameter [HM07b]. Ridgelets [Can03]. Riemannian [Jup05, Jup08]. right [AW05, Jam06, KL04, KS07, VJ03]. right-censored [VJ03]. right-censoring [KL04]. Risk [CG06, Dre01, MN06, Zha02, BM06a, BM06c, Che04, CLV06, Kol06a, Kol06b, LV04, SW06, Spo09, Tsy06, Zha04, vdG06]. risks [BM04, GMW08a, GMW08b]. RKHS [HR09]. RMT [BJYZ09]. Robbins [Ano03, Efr03, Sie03]. Robust [BZ01, BHZ06, BCZ08, bCH09, CAMMI08, DF01, FYZ01, KLF04, MPY09, RW01, YZ04, CZ09a, GR05, GEGMMI08, GY02, GD04, HM07c, Hub02, SBZ02, SBZ04, SBY08]. Robustness [CH09, AC04]. Rodeo [LW08]. root [MM07, PS04, SW07a, TvdG05]. root- [SW07a]. Rosenblatt [Son09]. Rotation [SSSW03]. rules [FF08, NS01, Pen06]. run [CHS09]. Runs [DK01].

Saddlepoint [BW04, JSZ04, RRY03, Kol03]. SAMC [Lia09]. same [IW05]. same-realization [IW05]. sample [AZ08, BZ09, BD08, BP03, BP05, BO08, BMW03, CL00, CY06, Coe08, DL03, FPW01, Gil00, GM07, JM09, Kim04, Kou04, LW02, Mar03, MS00b, Nad08, Ren08, RW00, SZ00b, Wu05, Zha08b, ZZ09a]. sample-fitted [BO08]. sampled [ASJ07]. sampler [AL06, CK06, CK08b, FGR00, KZW06b, ML06, PR08a, WZ06]. samplers [JH04]. samples [BC09, BP04, DM08, FHS09, HRW07, PA05]. sampling [AGL05, BO00, BMW03, Che04, CDS06, DGMR07, Gil00, HP04a, LM07a, LBM09, Loh08, Nea03, Owe08, PSS09]. SAMSI [Bic08]. satisfying [AS00]. saturated [VW01, WW07b]. SCAD [XH09]. SCAD-penalized [XH09]. scalar [GHR04]. Scale [CM00a, Hen04, PSS09, SBZ04]. scanning [MP07a]. Scarf [GW04b]. scatter [DSW09, HP08, KMO03, ZC05b]. scatterplot [Efr01]. scheduling [FH00]. scheme [AH06]. schemes [DGMR07, HT04]. Sciences [Bic08]. scope [HMV09]. score [ZZ06]. scrambled [Loh03, Owe08]. screening [Ste02]. Searching [ACCHZ08]. second [DGT06, DHP00, FPW01]. second-degree [DHP00]. second-order [DGT06]. Selecting [Ing04, WZ05]. Selection [Efr01, FL02, BCV01, BHL03, BGH09, BB04, BP04, Bun04, CP09, CGMM09, GHKM01, GG00, Gre06, HL05b, HN07, IW05, IR05, Jia07, JRGN08, JT08, JS05, Kou04, LM00, LP06, Lee09, LL08b, Lia09, LZ06b, LF09b, MP05a, MB06, PBHT08, WR09, Weg03, ZH08a, ZRY09, ZILZ07, ZY08b]. selector [Bic07, BRT09, CL07a, CT07a, CT07b, EHT07, FS07, Rit07]. self [Coe08, Dah06]. self-similar [Coe08, Dah06]. Semi [Tau01]. Semi-Parametric [Tau01]. semidefinite [AW09]. semimartingale [GK01]. Semiparametric [BMV06, CH06b, CHT08, Hid05, Nai04, ZY08a, ALL05, BRS06, BMW03, Bun04, CWY09, CK08b, DGT06, GLT06, Gil00, GR03, GZH09, KP05a, LL08b, LSV08, LL09b, NKY09, Nis09, Osi09, Ren08, SKF09, Wal01, WZ07, YM09, ZFS09, ZL09]. Semiparametrically [HP06, HOP06]. Separation [LMP01, LPM03]. sequence [BL07, HT01, RSV04, SY00, SY03]. sequences [CSW06, CSW07, HWZ07, HI01, JS04, Lin07, PSS09, Wu05]. Sequential [CDS06, DG00, HM03, Mei06b, Mou08, PS00, Bei00, CL00, CFH07, CY09, Che00, Cho04, CSW06, CSW07, DM08, GK01, LP05, Sie03]. Serial [HVW06]. series [ADB07, AHHR09, BDT01, Bri02b, CL08, CL10, FPW01, GKLT09, GZ01, Ing07, KT01a, KL06a, Lin07, MRT08, Nis09, Rob05, SW07b, SM06b, YM09, ZW08, ZW09, ZSW22]. set [Che04]. Sets [WM07, BM07, Bha06, CJL07, GKK^{+}00, GW05, RvdV06, SSN09]. setting [Efr07a, PS09]. several [GH02, GM07]. shape [DMP05, Dry05, HO04, HP06, HOP06, MW00, Nau06, Reb05, TW08]. shape-restricted [MW00]. shapes [MP05b]. shared [GZH09]. Sharp [Dal05, KT01c, Zha09]. sharpening [CHR00]. short [CL08, CL10, Lah03, NL06]. short- [Lah03, NL06]. shortcoming [IKL00a, IKL00b]. Shrinkage [Kom06, Ber04, Pen06]. shrinking [CL05a]. sib [GX00]. sib-pair [GX00]. sided [CS05b]. Sieve [FZ04, GW00]. sign [HVW06]. sign-and-rank [HVW06]. signal [WW00]. signed [HS07, Roh08]. Significance [DM01]. significant [ZY08a]. Simes [Sar08]. similar [Coe08, Dah06]. simple [Bar07, MP06, RSV04]. simplex [CKS02, DHP00]. Simultaneous [BRT09, Kom04, Lec07, WW07b]. Single [MZ02, HJS01, Sar06, SZ05]. single-index [HJS01, SZ05]. single-step [Sar06]. Singular [Sri03]. singularities [Drt09]. Size [ABK00, Efr07b, CL00, CY06, JM09, LW02, NL04, Rob09]. sizes [XY03]. slab [IR05]. Slice [Nea03]. sliced [LZ07a]. small [DH02, KM07]. Smallest [BEM01]. Smooth [MT04, YPM08, EL00, GvdV07b, HR00, LL00, MP05a, MP06, RSD07]. smoothed [HL05a]. smoothers [Efr01]. Smoothing [CKS09, GKK^{+}00, CPW07, GM05, Lia09, LZ06b, WY07]. Sobolev [Jup05, Jup08]. solution [Chi07, RZ07]. Solutions [PP00]. Some [CL07b, CC09, CM00b, LW02, Sar02, Zha09, BC03, Fuh10, GK08, Jen10, KL06a, Ren08, Zha00]. sources [SXS05]. Space [Fuh10, Jen10, BS09, CM00a, Fuh06, Lin00, QAW09, SSSW03]. space-filling [QAW09]. spaced [ASM04]. spaces [CL04b, Dry05, DH02]. spacings [LL08a]. Sparse [Kol09, LW08, AW09, BM08, CJL07, DJ04, DJ06, HHM08, JS04, Kar08a, LF09b, MY09, Men08, Zha08a, ZL08a, ZL08b]. Sparsistency [LF09a]. sparsity [ABDJ06, ZH08a]. Spatial [BS07, dHP06, BBLR05, Cha09, GLT06, HLT04, Jam06, LZ06a, LWH^{+}07, LC05, NL04]. Specification [GKLT09, CGT08, GL05]. specified [DS02]. spectra [Hid05, VvS08]. Spectral [LR02, LR03, Bri02b, DK04, EPdH01, ES09, GHR01, SW07b, ZSW22, vLBB08]. spectroscopy [SBY09]. Spectrum [Kar08b]. spheres [Dry05]. spherical [BKMP09b, DMP05, MP04, Mar06]. sphericity [HP06]. Spike [IR05, CL07b, LS00]. spin [Cha07]. Spline [WY07, BW07, GKK^{+}00, Hua03, Kou04]. Spline-backfitted [WY07]. splines [CKS09]. split [BM07]. spread [HQ07]. SPRT [Fuh03]. Square [TvdG05, BS09]. squared [DJR04, HM06, HT01]. squares [BKL^{+}09, BC07, FSR08, LF09b, SZ01, Zha09]. Stability [PR08a]. Stable [MP07b, MS06, ACD09a]. Stahel [ZCH04]. State [Fuh10, Jen10, BS09, Fuh06, Lua04]. state-space [BS09]. stationary [CW04, CM03, Dah00, DP06, Hid05, LRZ03, LL09a, Nor09, Rob05, Rob08, dHP06]. Statist [Fuh10, Jen10]. statistic [GM07, HS07, JSZ04, Lah05, TS06]. Statistical [Bic07, BBM08, CL07a, CT07a, CT07b, Che04, DR06a, Dry05, EHT07, FS07, LB04, Lok08, MS04, OW00, RMSE08, Rit07, TV07, Zha04, ZC07, ZL09, Aud09, CHC03, FHT00, Gre06, Hel06, KSC07, KZW06b, MN06, McC02, MM07, Myk03, SZ00a, SZ00b, Tsy04, vdV02]. Statistics [FZZ01, GS08, IKL00a, BG01, Cha09, CLV08, FCZ02, HVW06, Hub02, JW03, KvdV06, MP07a, SW04, TW08, ZH06, Bic08]. status [GMW08a, GMW08b, MK05, VJ03]. Stein [PR08b]. Step [WW07b, BM08, Cha07, CS05a, GBS09, Men08, Sar06, Zha08a, ZL08a, ZL08b]. step-down [GBS09]. Step-up [WW07b, CS05a]. stepdown [LRS05]. Stepup [RS06a, Sar07, LRS05, Sar08]. Stepwise [Sar02, HN07]. Stiefel [CR01]. Stochastic [Lai03, AS00, CHC03, CH02, Dip03, FH00, GW04a, LZ06a, MP07c, Nis09, Sko00, SM06b, TV07]. stopping [PSS09, ZY05]. strategies [CFH07, FH00, IR05, Yan00]. stratification [BO08]. Stratified [GHKM01, FHS09]. straw [JS04]. strength [ST07]. Strings [DK01]. Strong [CSW06, CSW07, Sko00, BC07, HJ08]. strongly [LRZ03]. Structural [SZ00b, Tau01]. Structure [HJPS01, Zhu03, AMR09, AHHR09, Mad00]. structured [LL00, Ols07]. structures [ACDH06, VJ03]. Student [JSZ04]. Studentized [Lah05]. studies [Che00, SZT07, WF09]. study [BD08, CR01]. subclass [Loh05]. subject [HH01]. subsample [NL04]. subset [Gre06, HN07]. Subspace [AR09a]. subspaces [CH06b, RBD09]. success [HS02]. Sufficient [CCL02, JH04, Coo04, Lah03, VR09]. suitability [HM07a]. sum [KY05, ZH06]. sums [BG03b, Lah07, Zha05a]. Superefficiency [CL05a]. supersaturated [BC04, XW05]. Support [YS09, BBM08, BD07, KL06b, SS07, SA09]. surface [CFRC07, HR00]. surfaces [HM03]. surrogate [LY07, NWJ09]. survey [BO00]. surveys [BO08]. survival [GZH09, KL01, PR06, Sch01, Zen04, ZC05a]. survivor [AW05]. symmetric [CV03, GLM03, HWH07, MN09, NS09, RS06b, SMT08]. symmetry [Hel06]. synthesis [LB06]. system [SA09]. systems [GW04b, Lua04, Rob08].

table [BK01]. tables [BB09, CDS06, Lan04]. Tail [KT01d, EdHL06, Kol03]. tailed [Pen04, PQ06]. Tailor [BRS06]. Tailor-made [BRS06]. tale [MRY07]. tapered [DZM09]. Tensor [Lin00, KSC07]. Terry [Hun04]. test [BF00, Cha09, CGT08, CL09, DL03, HY05b, HC08, Joh04, LRS05, Lia00, OS09, Sar08, TW08, ZZ06]. testability [Cat04]. tested [MR06]. testimation [AGP07]. Testing [ASJ09b, BHL05, Cha06, Coo04, DS01, GSvdV00, HH00, HV05, HM07a, JT09, LT05, Lin07, STT03, Son09, TK03, BY01, BG05, But07, CY01, CH09, CSX09, DM01, EdHL06, FR02, GKLT09, GL05, Kar05, KSC07, NS09, OGV09, PS00, Rom05, RW07, Sar02, Sar06, ST01, SRB07, YM09]. Tests [BW01, ACH04, BZ09, BLBL02, BHL03, Bar07, BD08, BRS06, BM06b, CL00, CY09, DHV05, Drt09, FZ04, FL06, GQR07, HP02b, HP04b, HP06, HP08, JW07, Jan00, JP03, Jia01, JN02, Jup05, Jup08, KK04, LW02, LS03, RRY03, Roh08, Shi04, SKF09, WW07b, ZH06]. Texture [LB06]. their [ACH04, BC03, CvK03, Jia02, KY05, KT01d]. theorem [Cho04, KL04, Kim06, Loh08, LZ07b, WW07a]. theorems [AGL05, DM08, Lin07]. theoretical [CL07b, GX00, HM08, Kün05]. Theory [Dre01, GJW01b, Lia09, RGN02, AC04, ASM04, AS00, BW07, BRW09, BP03, BP05, Bic08, Bre04, BMW03, BCLZ02, BCLZ04, BC07, CL04a, CY09, CT05, CS05b, DGM06, FS02a, GW04b, Gil00, GD04, Hös03, JLW02, Kar08b, MS00b, NKY09, Nis09, PP00, Rei03, SW07b, XC08, Zha03, ZSW22, ZY08b]. third [MM07]. Three [KY04, DMP05, MRY07]. three-dimensional [DMP05]. threshold [KS07, LT05, Pon03, ZW03]. Thresholding [KM03, AH03b, AH05, BL08a, CZ09b, DJ06, ZH05]. thresholds [JS05, MM07]. Tilted [CWTG08, Sch07]. time [ADB07, AHHR09, BDT01, Bri02b, CFZZ07, CL08, CL10, DR06a, DGM06, FPW01, FSR08, GK01, GKLT09, GZ01, Ing07, KT01a, KL06a, Lin07, Lok08, MS04, MPR05, MRT08, Mou04, NKY09, Nis09, Rob05, SW07b, SM06b, TS06, VJ03, YM09, ZC05a, ZSW22, ZW08, ZW09]. time-dependent [DGM06]. time-inhomogeneous [MS04]. time-varying [DR06a, FSR08]. times [PSS09, Sch01]. timeseries [KP03]. tolerance [BEM01, LL08a, LL09b]. Tomography [BGA07, Cav00, Pan09]. tool [FS02a]. topography [RL05]. toric [GMS06]. total [BD04]. Tracking [HR00]. Tracy [Joh08]. trail [ACCHZ08]. train [CL07b]. Training [BP04, HI01, Jon00]. trains [LS00]. trait [GX00]. Transform [Mer07, GZ04, MM09]. transformation [CL04c, KS07, LSV08, MK05]. transforms [KK04, Lah03, MS06, Son09]. transience [HS02]. treatment [HY05b]. treatments [HY05b]. Tree [Ols07]. Tree-structured [Ols07]. trees [BM07, SS00, WM07]. trend [BS09]. Trials [BHR02, HZ04]. Trimming [CAMMI08, GEGMMI08, Zuo06]. truly [ACD09b]. truncated [BW04]. Tukey [Ano02a, BB02, Bri02a, Bri02b, CT02, Dem02, FS02b, Hub02, Spe02]. tuning [SV09]. Tusnády [CP04]. twice [PR06]. Two [WZ00, WZ07, BMV06, BMW03, CC06, DS02, DL03, FGR00, MF09, Ren08, RBK05, ST07, WW02, YS09, Ye03, CHC03]. Two-class [CHC03]. two-color [MF09]. two-component [BMV06]. two-factor [DS02, WW02]. two-Level [ST07, CC06, Ye03]. two-phase [BMW03, RBK05]. two-sample [DL03, Ren08]. type [BM04, EM05, FK00, GLdW03, HPT02, KMT07, MY09, PP00, FR02, HRW07, Loh05]. types [Ren08].

UMPU [BLBL02]. unbiased [MZ02, Shi04]. unconditional [AW05]. unidentifiable [Fuk03]. unified [JLW02, LMR^{+}08, LF09b]. Uniform [EM05, FPW01, PS09, SBZ04, AS00, ACDH06, HY03b]. uniformity [Jup08]. Uniformly [SW07a]. unimodal [EL00, SZ01]. Uniqueness [KT01b, TT00]. unit [PS04]. units [MS08]. univariate [CPW07, KLF04]. Universal [HY03b, SBR05, Gol09]. unknown [ABDJ06, BGH09, Car07, GHR01, Hid05, HM07d, Ing04, Joh09, Mei06b, SBZ04, STT03, SA09, Ste00]. unobservable [Pan09]. unstable [BC07]. urn [MF09]. Using [HM09, BKMP09a, Ber04, CR01, Coe08, FF08, HY05a, HL05b, Kar08b, Lia09, LL09b, LF09b, PR08b, RSD07, Shi04, SV09, SS07, Zen04, vdVvZ09].

Valid [LRZ03]. validation [BWB^{+}07, Leu05, WZ05, Yan07]. validity [DGMS00]. Value [Dre01, EPdH01, EL06, EdHL06, ES09, GS09, GLdW03, Spo09, Sto03, dHL03]. value-at-risk [Spo09]. values [Rob09, SY00, SY03, WC09]. Vanishing [IKL00a, IKL00b]. Variable [FL02, HL05b, LL08b, CGMM09, IR05, Jia07, JT08, MB06, WR09, ZRY09, ZH08b]. Variables [Tau01, AMR09, BF00, BC02, BW04, GM07, HH05, HM07a, Lah07, PA05, SS00, Zha05a]. Variance [BL07, BGH09, CW08, CPW07, Gel05, HY03a, HC08, Lah05, LZ07a, NL04, Nor09, Spe02, WBCL08]. variances [Car07]. various [Ren08]. VARMA [HP04b]. varying [ALL05, CFZZ07, DR06a, FSR08, Kim07, MPR05, ŞM06a, WZZ09, ZL09]. varying-coefficient [ZL09]. vector [BHL05, BBM08, BG09, KSC07, Mar03, SS07, SA09]. vectors [STT03]. Vervaat [CSW06, CSW07]. via [BCV01, BL07, BCZ08, EHJL08, Ho06, Hof03, JW07, Kol09, LW04, PX09, ŞM06a, Son09, SDG06]. view [CM00a, FHT00]. Vines [BC02]. Volatility [ASJ07, MM09, MS04].

W [Ano02a, BB02, Bri02a, Bri02b, Dem02, FS02b, Hub02, Spe02]. Wald [BD08]. Wavelet [AH03b, AH05, HP04a, KR09, ZW03, BCZ08, CZ09b, GW05, JS05, MRT08, Pen06, VvS08, Zha05b, ZH05]. Wavelet-based [HP04a]. way [Ber04]. Weak [BDN02, GCL01, OGV09, dHL03, CC09, Jia02, Lah07]. weakly [SW07a]. Weighted [EdHL06, Ren08, DM08, ZCH04, ZC05b]. weights [WZ05]. where [BLBL02]. which [BR03, BM06b]. white [BCLZ04, Men08]. Whittle [GR03, MRT08, PS04, Rob08, SP05]. Who [Men08]. whose [Hid05]. Wicksell [WW07a]. Widom [Joh08]. Wigner [BGA07]. Wilks [FCZ02, FZZ01]. Wishart [CV03, DMO08, DG12, GLM03, LM07b, RMSE08, Sri03]. within [BF00]. Wolfowitz [MP07c, WW07a]. Wong [AL06, CK06, ML06, WZ06]. work [Bri02b, FS02b, JP03, vdV02]. writings [Ano02a, Ano02b, Ano03].

Yakir [Mei06a].

Zhou [AL06, CK06, ML06, WZ06]. zonotope [KMO03].

References

Antoniadis:2006:PIP

[AB06] Anestis Antoniadis and Jéremie Bigot. Poisson inverse problems. Annals of Statistics, 34(5):2132–2158, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

            http://projecteuclid.org/euclid.aos/1169571792.

Abramovich:2006:AUS

[ABDJ06] Felix Abramovich, Yoav Benjamini, David L. Donoho, and Iain M. Johnstone. Adapting to unknown sparsity by controlling the false discovery rate. Annals of Statistics, 34(2):584–653, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

            8966 (electronic). URL http://projecteuclid.org/euclid.
            aos/1151418235.

Albers:2000:SPP

[ABK00] Willem Albers, Pieta C. Boon, and Wilbert C. M. Kallenberg. Size and power of pretest procedures. Annals of Statistics, 28 (1):195–214, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016120369.

Abraham:2004:AGR

[AC04] Christophe Abraham and Benoît Cadre. Asymptotic global robustness in Bayesian decision theory. Annals of Statistics, 32(4): 1341–1366, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626172.

Anderes:2009:CED

[AC09] Ethan Anderes and Sourav Chatterjee. Consistent estimates of deformed isotropic Gaussian random fields on the plane. Annals of Statistics, 37(5A):2324–2350, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1247663757.

Arias-Castro:2008:STE

[ACCHZ08] Ery Arias-Castro, Emmanuel J. Candès, Hannes Helgason, and Ofer Zeitouni. Searching for a trail of evidence in a maze. Annals of Statistics, 36(4):1726–1757, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237298.

Andrews:2009:MLE

[ACD09a] Beth Andrews, Matthew Calder, and Richard A. Davis. Maximum likelihood estimation for α-stable autoregressive processes. Annals of Statistics, 37(4):1946–1982, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1245332837.

Arias-Castro:2009:DMF

[ACD09b] Ery Arias-Castro and David L. Donoho. Does median filtering truly preserve edges better than linear filtering? Annals of Statistics, 37(3):1172–1206, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1239369019.

Arias-Castro:2006:AMD

[ACDH06] Ery Arias-Castro, David L. Donoho, and Xiaoming Huo. Adaptive multiscale detection of filamentary structures in a background of uniform random points. Annals of Statistics, 34(1): 326–349, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576265.

Aerts:2004:BMT

[ACH04] Marc Aerts, Gerda Claeskens, and Jeffrey D. Hart. Bayesianmotivated tests of function fit and their asymptotic frequentist properties. Annals of Statistics, 32(6):2580–2615, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1107794880.

Andrews:2007:RBE

[ADB07] Beth Andrews, Richard A. Davis, and F. Jay Breidt. Rank-based estimation for all-pass time series models. Annals of Statistics, 35(2):844–869, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1183667296.

Almudevar:2000:DME

[AFR00] Anthony Almudevar, Chris Field, and John Robinson. The density of multivariate M-estimates. Annals of Statistics, 28(1): 275–297, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016120373.

Arratia:2005:LCL

[AGL05] Richard Arratia, Larry Goldstein, and Bryan Langholz. Local central limit theorems, the high-order correlations of rejective sampling and logistic likelihood asymptotics. Annals of Statistics, 33(2):871–914, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114339.

Abramovich:2007:OBT

[AGP07] Felix Abramovich, Vadim Grinshtein, and Marianna Pensky. On optimality of Bayesian testimation in the normal means problem. Annals of Statistics, 35(5):2261–2286, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461730.

Akharif:2003:EDR

[AH03a] Abdelhadi Akharif and Marc Hallin. Efficient detection of random coefficients in autoregressive models. Annals of Statistics, 31(2):675–704, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1051027885.

Averkamp:2003:WTN

[AH03b] R. Averkamp and C. Houdré. Wavelet thresholding for nonnecessarily Gaussian noise: idealism. Annals of Statistics, 31 (1):110–151, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1046294459.

Averkamp:2005:WTN

[AH05] R. Averkamp and C. Houdré. Wavelet thresholding for nonnecessarily Gaussian noise: Functionality. Annals of Statistics, 33(5): 2164–2193, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1132936560.

Anevski:2006:GAS

[AH06] D. Anevski and O. Hössjer. A general asymptotic scheme for inference under order restrictions. Annals of Statistics, 34(4): 1874–1930, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567637.

Aue:2009:BDC

[AHHR09] Alexander Aue, Siegfried Hörmann, Lajos Horváth, and Matthew Reimherr. Break detection in the covariance structure of multivariate time series models. Annals of Statistics, 37(6B): 4046–4087, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303536.

Angers:2005:MBF

[AK05] Jean-François Angers and Peter T. Kim. Multivariate Bayesian function estimation. Annals of Statistics, 33(6):2967–2999, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1140191681.

Atchade:2006:DEE

[AL06] Yves F. Atchadé and Jun S. Liu. Discussion of “Equi-energy sampler” by Kou, Zhou and Wong. Annals of Statistics, 34(4):

1620–1628, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567623.

Ahmad:2005:EES

[ALL05] Ibrahim Ahmad, Sittisak Leelahanon, and Qi Li. Efficient estimation of a semiparametric partially linear varying coefficient model. Annals of Statistics, 33(1):258–283, February 2005. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1112967706.

Allman:2009:IPL

[AMR09] Elizabeth S. Allman, Catherine Matias, and John A. Rhodes. Identifiability of parameters in latent structure models with many observed variables. Annals of Statistics, 37(6A):3099– 3132, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515381.

Anderson:2002:CCA

[And02] T. W. Anderson. Canonical correlation analysis and reduced rank regression in autoregressive models. Annals of Statistics, 30 (4):1134–1154, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031689020.

Anonymous:2002:PWJ

[Ano02a] Anonymous. The publications and writings of John w. Tukey. Annals of Statistics, 30(6):1666–1680, December 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1043351253;
 http://www.jstor.org/stable/1558736.

Anonymous:2002:PWL

[Ano02b] Anonymous. The publications and writings of Lucien Le Cam. Annals of Statistics, 30(3):683–687, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1028674837.

Anonymous:2003:PWH

[Ano03] Anonymous. The publications and writings of Herbert Robbins. Annals of Statistics, 31(2):407–413, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1051027874.

Andersson:2006:CME

[AP06] Steen A. Andersson and Michael D. Perlman. Characterizing Markov equivalence classes for AMP chain graph models. Annals of Statistics, 34(2):939–972, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418247.

Andersson:2009:SEP

[AR09a] Sofia Andersson and Tobias Rydén. Subspace estimation and prediction methods for hidden Markov models. Annals of Statistics, 37(6B):4131–4152, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1256303539.

Andrieu:2009:PMA

[AR09b] Christophe Andrieu and Gareth O. Roberts. The pseudomarginal approach for efficient Monte Carlo computations. Annals of Statistics, 37(2):697–725, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693147.

Ali:2009:MEA

[ARS09] R. Ayesha Ali, Thomas S. Richardson, and Peter Spirtes. Markov equivalence for ancestral graphs. Annals of Statistics, 37(5B): 2808–2837, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247836670.

Arcones:2000:ADT

[AS00] Miguel A. Arcones and Francisco J. Samaniego. On the asymptotic distribution theory of a class of consistent estimators of a distribution satisfying a uniform stochastic ordering constraint. Annals of Statistics, 28(1):116–150, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1016120367.

Ait-Sahalia:2008:CFL

[AS08a] Yacine Aït-Sahalia. Closed-form likelihood expansions for multivariate diffusions. Annals of Statistics, 36(2):906–937, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1205420523.

Anderes:2008:EDI

[AS08b] Ethan B. Anderes and Michael L. Stein. Estimating deformations of isotropic Gaussian random fields on the plane. Annals of Statistics, 36(2):719–741, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1205420517.

Ait-Sahalia:2007:VED

[ASJ07] Yacine Aït-Sahalia and Jean Jacod. Volatility estimators for discretely sampled Lévy processes. Annals of Statistics, 35(1): 355–392, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1181100191.

Ait-Sahalia:2009:EDA

[ASJ09a] Yacine Aït-Sahalia and Jean Jacod. Estimating the degree of activity of jumps in high frequency data. Annals of Statistics, 37(5A):2202–2244, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663753.

Ait-Sahalia:2009:TJD

[ASJ09b] Yacine Aït-Sahalia and Jean Jacod. Testing for jumps in a discretely observed process. Annals of Statistics, 37(1):184–222, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115932.

Aslan:2006:AMB

[Asl06] Mihaela Aslan. Asymptotically minimax Bayes predictive densities. Annals of Statistics, 34(6):2921–2938, December 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1179935070.

Ait-Sahalia:2004:EDR

[ASM04] Yacine Aït-Sahalia and Per A. Mykland. Estimators of diffusions with randomly spaced discrete observations: A general theory. Annals of Statistics, 32(5):2186–2222, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1098883787.

Audibert:2007:FLR

[AT07] Jean-Yves Audibert and Alexandre B. Tsybakov. Fast learning rates for plug-in classifiers. Annals of Statistics, 35(2):608–633, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1183667286.

Audibert:2009:FLR

[Aud09] Jean-Yves Audibert. Fast learning rates in statistical inference through aggregation. Annals of Statistics, 37(4):1591–1646, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1245332827.

Alqallaf:2009:POM

[AVYZ09] Fatemah Alqallaf, Stefan Van Aelst, Victor J. Yohai, and Ruben H. Zamar. Propagation of outliers in multivariate data. Annals of Statistics, 37(1):311–331, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1232115936.

Asgharian:2005:ABU

[AW05] Masoud Asgharian and David B. Wolfson. Asymptotic behavior of the unconditional NPMLE of the length-biased survivor function from right censored prevalent cohort data. Annals of Statistics, 33(5):2109–2131, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1132936558.

Amini:2009:HDA

[AW09] Arash A. Amini and Martin J. Wainwright. High-dimensional analysis of semidefinite relaxations for sparse principal components. Annals of Statistics, 37(5B):2877–2921, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1247836672.

Adrover:2002:PEM

[AY02] Jorge Adrover and Víctor Yohai. Projection estimates of multivariate location. Annals of Statistics, 30(6):1760–1781, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1043351256.

Anderson:2008:CRS

[AZ08] Greg W. Anderson and Ofer Zeitouni. A CLT for regularized sample covariance matrices. Annals of Statistics, 36(6): 2553–2576, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1231165179.

Banerjee:2007:LBI

[Ban07] Moulinath Banerjee. Likelihood based inference for monotone response models. Annals of Statistics, 35(3):931–956, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1185303993.

Baraud:2004:CBG

[Bar04] Yannick Baraud. Confidence balls in Gaussian regression. Annals of Statistics, 32(2):528–551, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1083178937.

Bartroff:2007:AOM

[Bar07] Jay Bartroff. Asymptotically optimal multistage tests of simple hypotheses. Annals of Statistics, 35(5):2075–2105, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1194461723.

Bagchi:2001:OPG

[BB01] Bhaskar Bagchi and Sunanda Bagchi. Optimality of partial geometric designs. Annals of Statistics, 29(2):577–594, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1009210554.

Benjamini:2002:JWT

[BB02] Yoav Benjamini and Henry Braun. John w. Tukey’s contributions to multiple comparisons. Annals of Statistics, 30(6):1576–1594, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1043351247;   http://www.jstor.org/stable/
1558730.

Barbieri:2004:OPM

[BB04] Maria Maddalena Barbieri and James O. Berger. Optimal predictive model selection. Annals of Statistics, 32(3):870–897, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1085408489.

Brien:2009:DTE

[BB09] C. J. Brien and R. A. Bailey. Decomposition tables for experiments I. A chain of randomizations. Annals of Statistics, 37(6B): 4184–4213, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303541.

Bartlett:2004:DBP

[BBB^{+}04] Peter L. Bartlett, Peter J. Bickel, Peter Bühlmann, Yoav Freund, Jerome Friedman, Trevor Hastie, Wenxin Jiang, Michael J. Jordan, Vladimir Koltchinskii, Gábor Lugosi, Jon D. McAuliffe, Ya’acov Ritov, Saharan Rosset, Robert E. Schapire, Robert Tibshirani, Nicolas Vayatis, Bin Yu, Tong Zhang, and Ji Zhu. Discussions of boosting papers, and rejoinders. Annals of Statistics, 32 (1):85–134, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1105988581.

Boissy:2005:PEF

[BBLR05] Y. Boissy, B. B. Bhattacharyya, X. Li, and G. D. Richardson. Parameter estimates for fractional autoregressive spatial processes. Annals of Statistics, 33(6):2553–2567, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1140191666.

Bartlett:2005:LRC

[BBM05] Peter L. Bartlett, Olivier Bousquet, and Shahar Mendelson. Local Rademacher complexities. Annals of Statistics, 33(4): 1497–1537, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1123250221.

Blanchard:2008:SPS

[BBM08] Gilles Blanchard, Olivier Bousquet, and Pascal Massart. Statistical performance of support vector machines. Annals of Statistics, 36(2):489–531, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1205420509.

Berger:2009:FDR

[BBS09] James O. Berger, José M. Bernardo, and Dongchu Sun. The formal definition of reference priors. Annals of Statistics, 37 (2):905–938, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693154.

Bedford:2002:VNG

[BC02] Tim Bedford and Roger M. Cooke. Vines — a new graphical model for dependent random variables. Annals of Statistics, 30 (4):1031–1068, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031689016.

Bulutoglu:2003:HPP

[BC03] Dursun A. Bulutoglu and Ching-Shui Cheng. Hidden projection properties of some nonregular fractional factorial designs and their applications. Annals of Statistics, 31(3):1012–1026, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562472.

Bulutoglu:2004:COS

[BC04] Dursun A. Bulutoglu and Ching-Shui Cheng. Construction of E(s^{2})-optimal supersaturated designs. Annals of Statistics, 32 (4):1662–1678, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626182.

Buchmann:2007:ATL

[BC07] Boris Buchmann and Ngai Hang Chan. Asymptotic theory of least squares estimators for nearly unstable processes under strong dependence. Annals of Statistics, 35(5):2001–2017, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1194461720.

Belloni:2009:CCM

[BC09] Alexandre Belloni and Victor Chernozhukov. On the computational complexity of MCMC-based estimators in large samples. Annals of Statistics, 37(4):2011–2055, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1245332839.

Brown:2002:CIB

[BCD02] Lawrence D. Brown, T. Tony Cai, and Anirban DasGupta. Confidence intervals for a binomial proportion and asymptotic expansions. Annals of Statistics, 30(1):160–201, February 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1015362189.

Barron:2008:ALG

[BCDD08] Andrew R. Barron, Albert Cohen, Wolfgang Dahmen, and Ronald A. DeVore. Approximation and learning by greedy algorithms. Annals of Statistics, 36(1):64–94, February 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1201877294.

Chan:2009:RNN

[bCH09] Yao ban Chan and Peter Hall. Robust nearest-neighbor methods for classifying high-dimensional data. Annals of Statistics, 37(6A):3186–3203, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1250515384.

Brown:2002:AET

[BCLZ02] Lawrence D. Brown, T. Tony Cai, Mark G. Low, and Cun-Hui Zhang. Asymptotic equivalence theory for nonparametric regression with random design. Annals of Statistics, 30(3):688–707, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1028674838.

Brown:2004:ETD

[BCLZ04] Lawrence D. Brown, Andrew V. Carter, Mark G. Low, and Cun-Hui Zhang. Equivalence theory for density estimation, Poisson processes and Gaussian white noise with drift. Annals of Statistics, 32(5):2074–2097, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1098883782.

Baraud:2001:AEA

[BCV01] Y Baraud, F. Comte, and G. Viennet. Adaptive estimation in autoregression or β-mixing regression via model selection. Annals of Statistics, 29(3):839–875, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1009210692.

Brown:2008:RNE

[BCZ08] Lawrence D. Brown, T. Tony Cai, and Harrison H. Zhou. Robust nonparametric estimation via wavelet median regression. Annals of Statistics, 36(5):2055–2084, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908084.

Bailey:2004:ONB

[BD04] R. A. Bailey and P. Druilhet. Optimality of neighbor-balanced designs for total effects. Annals of Statistics, 32(4):1650–1661, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1091626181.

Braess:2007:NSP

[BD07] Dietrich Braess and Holger Dette. On the number of support points of maximin and Bayesian optimal designs. Annals of Statistics, 35(2):772–792, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1183667293.

Belloni:2008:BFP

[BD08] Alexandre Belloni and Gustavo Didier. On the Behrens–Fisher problem: A globally convergent algorithm and a finite-sample study of the Wald, LR and LM tests. Annals of Statistics, 36(5): 2377–2408, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1223908096.

Buhlmann:2002:WDB

[BDN02] Peter Bühlmann, Paul Doukhan, and Patrick Ango Nze. Weak dependence beyond mixing and asymptotics for nonparametric regression. Annals of Statistics, 30(2):397–430, April 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1021379859.

Breidt:2001:LAD

[BDT01] F. Jay Breidt, Richard A. Davis, and A. Alexandre Trindade. Least absolute deviation estimation for all-pass time series models. Annals of Statistics, 29(4):919–946, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1013699987.

Beibel:2000:NSD

[Bei00] M. Beibel. A note on sequential detection with exponential penalty for the delay. Annals of Statistics, 28(6):1696–1701, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957476.

Belitser:2000:RED

[Bel00] Eduard Belitser. Recursive estimation of a drifted autoregressive parameter. Annals of Statistics, 28(3):860–870, June 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1015952001.

Bucchianico:2001:SNT

[BEM01] Alessandro Di Bucchianico, John H. Einmahl, and Nino A. Mushkudiani. Smallest nonparametric tolerance regions. Annals of Statistics, 29(5):1320–1343, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013203456.

Beran:2004:HSE

[Ber04] Rudolf Beran. Hybrid shrinkage estimators using penalty bases for the ordinal one-way layout. Annals of Statistics, 32(6): 2532–2558, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1107794878.

Bartolucci:2000:LRT

[BF00] Francesco Bartolucci and Antonio Forcina. A likelihood ratio test for MTP_{2} within binary variables. Annals of Statistics, 28 (4):1206–1218, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956713.

Bloznelis:2001:ODF

[BG01] M. Bloznelis and F. Götze. Orthogonal decomposition of finite population statistics and its applications to distributional asymptotics. Annals of Statistics, 29(3):899–917, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1009210694.

Belitser:2003:ABI

[BG03a] Eduard Belitser and Subhashis Ghosal. Adaptive Bayesian inference on the mean of an infinite-dimensional normal distribution. Annals of Statistics, 31(2):536–559, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1051027880.

Buchmann:2003:DEP

[BG03b] Boris Buchmann and Rudolf Grübel. Decompounding: an estimation problem for Poisson random sums. Annals of Statistics, 31(4):1054–1074, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1059655905.

Blanchard:2005:HTD

[BG05] Gilles Blanchard and Donald Geman. Hierarchical testing designs for pattern recognition. Annals of Statistics, 33(3):1155–1202, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1120224099.

Brown:2009:NEB

[BG09] Lawrence D. Brown and Eitan Greenshtein. Nonparametric empirical Bayes and compound decision approaches to estimation of a high-dimensional vector of normal means. Annals of Statistics, 37(4):1685–1704, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1245332829.

Butucea:2007:MAE

[BGA07] Cristina Butucea, Mădălin Guţă, and Luis Artiles. Minimax and adaptive estimation of the Wigner function in quantum homodyne tomography with noisy data. Annals of Statistics, 35(2): 465–494, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1183667281;   http://www.jstor.org/stable/
25463565.

Baraud:2009:GMS

[BGH09] Yannick Baraud, Christophe Giraud, and Sylvie Huet. Gaussian model selection with an unknown variance. Annals of Statistics, 37(2):630–672, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693145.

Brown:2008:APD

[BGX08] Lawrence D. Brown, Edward I. George, and Xinyi Xu. Admissible predictive density estimation. Annals of Statistics, 36 (3):1156–1170, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1211819560.

Berkes:2004:EEP

[BH04] István Berkes and Lajos Horváth. The efficiency of the estimators of the parameters in GARCH processes. Annals of Statistics, 32 (2):633–655, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178941.

Bhattacharya:2006:IPG

[Bha06] Bhaskar Bhattacharya. An iterative procedure for general probability measures to obtain I-projections onto intersections of convex sets. Annals of Statistics, 34(2):878–902, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1151418244.

Benko:2009:CFP

[BHK09] Michal Benko, Wolfgang Härdle, and Alois Kneip. Common functional principal components. Annals of Statistics, 37(1):1–34, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115926.

Berkes:2006:DBL

[BHKS06] István Berkes, Lajos Horváth, Piotr Kokoszka, and Qi-Man Shao. On discriminating between long-range dependence and changes in mean. Annals of Statistics, 34(3):1140–1165, June 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1152540745.

Baraud:2003:ATL

[BHL03] Y. Baraud, S. Huet, and B. Laurent. Adaptive tests of linear hypotheses by model selection. Annals of Statistics, 31(1):225–251, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1046294463.

Baraud:2005:TCH

[BHL05] Yannick Baraud, Sylvie Huet, and Béatrice Laurent. Testing convex hypotheses on the mean of a Gaussian vector. Application to testing qualitative hypotheses on a regression function. Annals of Statistics, 33(1):214–257, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1112967705.

Bai:2002:APA

[BHR02] Z. D. Bai, Feifang Hu, and William F. Rosenberger. Asymptotic properties of adaptive designs for clinical trials with delayed response. Annals of Statistics, 30(1):122–139, February 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1015362187.

Boente:2006:REG

[BHZ06] Graciela Boente, Xuming He, and Jianhui Zhou. Robust estimates in generalized partially linear models. Annals of Statistics, 34(6):2856–2878, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935067.

Bickel:2007:DDS

[Bic07] Peter J. Bickel. Discussion: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6): 2352–2357, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201012959.

Bickel:2008:RMT

[Bic08] Peter Bickel. Random matrix theory: A program of the Statistics and Applied Mathematical Sciences Institute (SAMSI). Annals of Statistics, 36(6):2551–2552, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1231165178.

Bai:2009:CLL

[BJYZ09] Zhidong Bai, Dandan Jiang, Jian-Feng Yao, and Shurong Zheng. Corrections to LRT on large-dimensional covariance matrix by RMT. Annals of Statistics, 37(6B):3822–3840, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1256303528.

Brooks:2001:PIL

[BK01] S. P. Brooks and R. King. Prior induction in log-linear models for general contingency table analysis. Annals of Statistics, 29(3):715–747, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210687.

Boysen:2009:CRC

[BKL^{+}09] Leif Boysen, Angela Kempe, Volkmar Liebscher, Axel Munk, and Olaf Wittich. Consistencies and rates of convergence of jumppenalized least squares estimators. Annals of Statistics, 37(1): 157–183, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1232115931.

Baldi:2009:ADE

[BKMP09a] P. Baldi, G. Kerkyacharian, D. Marinucci, and D. Picard. Adaptive density estimation for directional data using needlets. Annals of Statistics, 37(6A):3362–3395, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515390.

Baldi:2009:ASN

[BKMP09b] P. Baldi, G. Kerkyacharian, D. Marinucci, and D. Picard. Asymptotics for spherical needlets. Annals of Statistics, 37(3): 1150–1171, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369018.

Brown:2007:VEN

[BL07] Lawrence D. Brown and M. Levine. Variance estimation in nonparametric regression via the difference sequence method. Annals of Statistics, 35(5):2219–2232, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1194461728.

Bickel:2008:CRT

[BL08a] Peter J. Bickel and Elizaveta Levina. Covariance regularization by thresholding. Annals of Statistics, 36(6):2577–2604, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1231165180.

Bickel:2008:REL

[BL08b] Peter J. Bickel and Elizaveta Levina. Regularized estimation of large covariance matrices. Annals of Statistics, 36(1):199–227, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877299.

Bar-Lev:2002:NGI

[BLBL02] Shaul K. Bar-Lev, Daoud Bshouty, and Gérard Letac. Normal, gamma and inverse-Gaussian are the only NEFs where the bilateral UMPU and GLR tests coincide. Annals of Statistics, 30(5): 1524–1534, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1035844987.

Bailey:2000:ODM

[BM00] R. A. Bailey and J. P. Morgan. Optimal design with many blocking factors. Annals of Statistics, 28(2):553–577, April 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1016218230.

Barmi:2004:CED

[BM04] Hammou El Barmi and Hari Mukerjee. Consistent estimation of distributions with type II bias with applications in competing risks problems. Annals of Statistics, 32(1):245–267, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1079120136.

Bartlett:2006:DLR

[BM06a] Peter L. Bartlett and Shahar Mendelson. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2657–2663, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1179935056.

Bischoff:2006:ODW

[BM06b] Wolfgang Bischoff and Frank Miller. Optimal designs which are efficient for lack of fit tests. Annals of Statistics, 34(4):2015–2025, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1162567641.

Blanchard:2006:DLR

[BM06c] Gilles Blanchard and Pascal Massart. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2664–2671, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1179935057.

Banerjee:2007:CSS

[BM07] Moulinath Banerjee and Ian W. McKeague. Confidence sets for split points in decision trees. Annals of Statistics, 35(2):543–574, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1183667284.

Buhlmann:2008:DOS

[BM08] Peter Bühlmann and Lukas Meier. Discussion: One-step sparse estimates in nonconcave penalized likelihood models. Annals of Statistics, 36(4):1534–1541, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237288.

Berrendero:2007:MBF

[BMT07] José R. Berrendero, Beatriz V. M. Mendes, and David E. Tyler. On the maximum bias functions of MM-estimates and constrained M-estimates of regression. Annals of Statistics, 35 (1):13–40, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1181100179.

Bordes:2006:SET

[BMV06] Laurent Bordes, Stéphane Mottelet, and Pierre Vandekerkhove. Semiparametric estimation of a two-component mixture model. Annals of Statistics, 34(3):1204–1232, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1152540747.

Breslow:2003:LST

[BMW03] Norman Breslow, Brad McNeney, and Jon A. Wellner. Large sample theory for semiparametric regression models with twophase, outcome dependent sampling. Annals of Statistics, 31(4): 1110–1139, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1059655907.

Breidt:2000:LPR

[BO00] F. Jay Breidt and Jean D. Opsomer. Local polynomial regression estimators in survey sampling. Annals of Statistics, 28(4): 1026–1053, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956706.

Breidt:2008:EPS

[BO08] F. Jay Breidt and Jean D. Opsomer. Endogenous poststratification in surveys: Classifying with a sample-fitted model. Annals of Statistics, 36(1):403–427, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1201877307.

Bhattacharya:2003:LST

[BP03] Rabi Bhattacharya and Vic Patrangenaru. Large sample theory of intrinsic and extrinsic sample means on manifolds. Annals of Statistics, 31(1):1–29, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1046294456.

Berger:2004:TSO

[BP04] James O. Berger and Luis R. Pericchi. Training samples in objective Bayesian model selection. Annals of Statistics, 32(3):841– 869, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1085408488.

Bhattacharya:2005:LST

[BP05] Rabi Bhattacharya and Vic Patrangenaru. Large sample theory of intrinsic and extrinsic sample means on manifolds — II. Annals of Statistics, 33(3):1225–1259, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1120224101.

Blasi:2009:APH

[BPP09] Pierpaolo De Blasi, Giovanni Peccati, and Igor Prünster. Asymptotics for posterior hazards. Annals of Statistics, 37(4):1906– 1945, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1245332836.

Beskos:2009:MCM

[BPR09] Alexandros Beskos, Omiros Papaspiliopoulos, and Gareth Roberts. Monte Carlo maximum likelihood estimation for discretely observed diffusion processes. Annals of Statistics, 37(1): 223–245, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1232115933.

Bergsma:2002:MMC

[BR02] Wicher P. Bergsma and Tamás Rudas. Marginal models for categorical data. Annals of Statistics, 30(1):140–159, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015362188.

Bickel:2003:NEW

[BR03] Peter J. Bickel and Ya’acov Ritov. Nonparametric estimators which can be “plugged-in”. Annals of Statistics, 31(4):1033–1053, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1059655904.

Breiman:2004:PTB

[Bre04] Leo Breiman. Population theory for boosting ensembles. Annals of Statistics, 32(1):1–11, February 2004. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1079120126.

Brillinger:2002:JWTa

[Bri02a] David R. Brillinger. John W. Tukey: his life and professional contributions. Annals of Statistics, 30(6):1535–1575, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1043351246;   http://www.jstor.org/stable/1558729;
http://www.stat.berkeley.edu/users/brill/Papers/life.
pdf.

Brillinger:2002:JWTb

[Bri02b] David R. Brillinger. John W. Tukey’s work on time series and spectral analysis. Annals of Statistics, 30(6):1595–1618, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1043351248;   http://www.jstor.org/stable/1558731;
http://www.stat.berkeley.edu/users/brill/Papers/time.
pdf..

Bickel:2006:TMT

[BRS06] Peter J. Bickel, Ya’acov Ritov, and Thomas M. Stoker. Tailormade tests for goodness of fit to semiparametric hypotheses. Annals of Statistics, 34(2):721–741, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418238.

Bickel:2009:SAL

[BRT09] Peter J. Bickel, Ya’acov Ritov, and Alexandre B. Tsybakov. Simultaneous analysis of Lasso and Dantzig selector. Annals of Statistics, 37(4):1705–1732, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1245332830.

Balabdaoui:2009:LDT

[BRW09] Fadoua Balabdaoui, Kaspar Rufibach, and Jon A. Wellner. Limit distribution theory for maximum likelihood estimation of a logconcave density. Annals of Statistics, 37(3):1299–1331, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1239369023.

Belomestny:2007:SAL

[BS07] Denis Belomestny and Vladimir Spokoiny. Spatial aggregation of local likelihood estimates with applications to classification. Annals of Statistics, 35(5):2287–2311, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461731.

Berger:2008:OPB

[BS08] James O. Berger and Dongchu Sun. Objective priors for the bivariate normal model. Annals of Statistics, 36(2):963–982, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1205420525.

Burman:2009:ETS

[BS09] Prabir Burman and Robert H. Shumway. Estimation of trend in state-space models: Asymptotic mean square error and rate of convergence. Annals of Statistics, 37(6B):3715–3742, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1256303525.

Balan:2005:ARG

[BSK05] R. M. Balan and I. Schiopu-Kratina. Asymptotic results with generalized estimating equations for longitudinal data. Annals of Statistics, 33(2):522–541, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114327.

Berger:2005:PPA

[BST05] James O. Berger, William Strawderman, and Dejun Tang. Posterior propriety and admissibility of hyperpriors in normal hierarchical models. Annals of Statistics, 33(2):606–646, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1117114331.

Bunea:2007:AGR

[BTW07] Florentina Bunea, Alexandre B. Tsybakov, and Marten H. Wegkamp. Aggregation for Gaussian regression. Annals of Statistics, 35(4):1674–1697, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405626.

Buhlmann:2006:BHD

[Büh06] Peter Bühlmann. Boosting for high-dimensional linear models. Annals of Statistics, 34(2):559–583, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1151418234.

Bunea:2004:CCS

[Bun04] Florentina Bunea. Consistent covariate selection and post model selection inference in semiparametric regression. Annals of Statistics, 32(3):898–927, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1085408490.

Butucea:2007:GFT

[But07] Cristina Butucea. Goodness-of-fit testing and quadratic functional estimation from indirect observations. Annals of Statistics, 35(5):1907–1930, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1194461716.

Banerjee:2001:LRT

[BW01] Moulinath Banerjee and Jon A. Wellner. Likelihood ratio tests for monotone functions. Annals of Statistics, 29(6):1699– 1731, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015345959.

Butler:2002:LAH

[BW02] Roland W. Butler and Andrew T. A. Wood. Laplace approximations for hypergeometric functions with matrix argument. Annals of Statistics, 30(4):1155–1177, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031689021.

Butler:2004:SAM

[BW04] Ronald W. Butler and Andrew T. A. Wood. Saddlepoint approximation for moment generating functions of truncated random variables. Annals of Statistics, 32(6):2712–2730, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1107794884.

Balabdaoui:2007:EKM

[BW07] Fadoua Balabdaoui and Jon A. Wellner. Estimation of a kmonotone density: Limit distribution theory and the spline connection. Annals of Statistics, 35(6):2536–2564, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1201012971.

Bardet:2009:ANQ

[BW09] Jean-Marc Bardet and Olivier Wintenberger. Asymptotic normality of the quasi-maximum likelihood estimator for multidimensional causal processes. Annals of Statistics, 37(5B): 2730–2759, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247836667.

Bayarri:2007:CMV

[BWB^{+}07] M. J. Bayarri, D. Walsh, J. O. Berger, J. Cafeo, G. Garcia-Donato, F. Liu, J. Palomo, R. J. Parthasarathy, R. Paulo, and J. Sacks. Computer model validation with functional output. Annals of Statistics, 35(5):1874–1906, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461715.

Benjamini:2001:CFD

[BY01] Yoav Benjamini and Daniel Yekutieli. The control of the false discovery rate in multiple testing under dependency. Annals of Statistics, 29(4):1165–1188, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013699998.

Buhlmann:2002:AB

[BY02] Peter Bühlmann and Bin Yu. Analyzing bagging. Annals of Statistics, 30(4):927–961, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031689014.

Berrendero:2001:MBC

[BZ01] José Berrendero and Ruben H. Zamar. Maximum bias curves for robust regression with non-elliptical regressors. Annals of Statistics, 29(1):224–251, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/996986507.

Balakrishnan:2009:NMS

[BZ09] N. Balakrishnan and Xingqiu Zhao. New multi-sample nonparametric tests for panel count data. Annals of Statistics, 37 (3):1112–1149, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369017.

Cuesta-Albertos:2008:TLR

[CAMMI08] Juan A. Cuesta-Albertos, Carlos Matrán, and Agustín Mayo-Iscar. Trimming and likelihood: Robust location and dispersion estimation in the elliptical model. Annals of Statistics, 36(5): 2284–2318, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1223908093.

Candes:2003:RER

[Can03] Emmanuel J. Candès. Ridgelets: estimating with ridge functions. Annals of Statistics, 31(5):1561–1599, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1065705119.

Carter:2002:DDB

[Car02] Andrew V. Carter. Deficiency distance between multinomial and multivariate normal experiments. Annals of Statistics, 30(3):708– 730, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1028674839.

Carter:2007:AAN

[Car07] Andrew V. Carter. Asymptotic approximation of nonparametric regression experiments with unknown variances. Annals of Statistics, 35(4):1644–1673, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405625.

Cator:2004:TCA

[Cat04] Eric A. Cator. On the testability of the CAR assumption. Annals of Statistics, 32(5):1957–1980, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1098883778.

Cavalier:2000:EED

[Cav00] Laurent Cavalier. Efficient estimation of a density in a problem of tomography. Annals of Statistics, 28(2):630–647, April 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1016218233.

Chatterjee:2005:GBE

[CB05] Snigdhansu Chatterjee and Arup Bose. Generalized bootstrap for estimating equations. Annals of Statistics, 33(1):414–436, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1112967711.

Chen:2006:EIC

[CB06] Aiyou Chen and Peter J. Bickel. Efficient independent component analysis. Annals of Statistics, 34(6):2825–2855, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1179935066.

Chen:2006:DPM

[CC06] Hegang H. Chen and Ching-Shui Cheng. Doubling and projection: A method of constructing two-level designs of resolution IV. Annals of Statistics, 34(1):546–558, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1146576274.

Chen:2009:SRD

[CC09] Hegang H. Chen and Ching-Shui Cheng. Some results on 2^{n−m} designs of resolution IV with (weak) minimum aberration. Annals of Statistics, 37(6A):3600–3615, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515398.

Chiaromonte:2002:SDR

[CCL02] Francesca Chiaromonte, R. Dennis Cook, and Bing Li. Sufficient dimensions reduction in regressions with categorical predictors. Annals of Statistics, 30(2):475–497, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1021379862.

Carolan:2001:MDL

[CD01] Chris Carolan and Richard Dykstra. Marginal densities of the least concave majorant of Brownian motion. Annals of Statistics, 29(6):1732–1750, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015345960.

Candes:2002:REI

[CD02] Emmanuel J. Candès and David L. Donoho. Recovering edges in ill-posed inverse problems: optimality of curvelet frames. Annals of Statistics, 30(3):784–842, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028674842.

Chen:2006:SIS

[CDS06] Yuguo Chen, Ian H. Dinwoodie, and Seth Sullivant. Sequential importance sampling for multiway tables. Annals of Statistics, 34 (1):523–545, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576273.

Chan:2007:OSC

[CFH07] Hock Peng Chan, Cheng-Der Fuh, and Inchi Hu. Optimal strategies for a class of sequential control problems with precedence relations. Annals of Statistics, 35(4):1722–1748, August 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1188405628.

Cuevas:2007:NAE

[CFRC07] Antonio Cuevas, Ricardo Fraiman, and Alberto Rodríguez-Casal. A nonparametric approach to the estimation of lengths and surface areas. Annals of Statistics, 35(3):1031–1051, July 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1185303997.

Cai:2007:HMV

[CFZZ07] Jianwen Cai, Jianqing Fan, Haibo Zhou, and Yong Zhou. Hazard models with varying coefficients for multivariate failure time data. Annals of Statistics, 35(1):324–354, February 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1181100190.

Cavalier:2006:RHM

[CG06] L. Cavalier and Yu. Golubev. Risk hull method and regularization by projections of ill-posed inverse problems. Annals of Statistics, 34(4):1653–1677, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1162567628.

Cui:2009:CAN

[CGLZ09] Xia Cui, Wensheng Guo, Lu Lin, and Lixing Zhu. Covariateadjusted nonlinear regression. Annals of Statistics, 37(4): 1839–1870, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1245332834.

Casella:2009:CBP

[CGMM09] George Casella, F. Javier Girón, M. Lina Martínez, and Elías Moreno. Consistency of Bayesian procedures for variable selection. Annals of Statistics, 37(3):1207–1228, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1239369020.

Cavalier:2002:OII

[CGPT02] L. Cavalier, G. K. Golubev, D. Picard, and A. B. Tsybakov. Oracle inequalities for inverse problems. Annals of Statistics, 30(3):843–874, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028674843.

Chen:2008:TMS

[CGT08] Song Xi Chen, Jiti Gao, and Cheng Yong Tang. A test for model specification of diffusion processes. Annals of Statistics, 36(1): 167–198, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201877298.

Claeskens:2002:EDS

[CH02] Gerda Claeskens and Peter Hall. Effect of dependence on stochastic measures of accuracy of density estimations. Annals of Statistics, 30(2):431–454, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1021379860.

Cai:2006:PFL

[CH06a] T. Tony Cai and Peter Hall. Prediction in functional linear regression. Annals of Statistics, 34(5):2159–2179, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1169571793.

Chen:2006:SEF

[CH06b] Willa W. Chen and Clifford M. Hurvich. Semiparametric estimation of fractional cointegrating subspaces. Annals of Statistics, 34(6):2939–2979, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935071.

Clarke:2009:RMT

[CH09] Sandy Clarke and Peter Hall. Robustness of multiple testing procedures against dependence. Annals of Statistics, 37(1):332–358, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115937.

Chambaz:2006:TOM

[Cha06] Antoine Chambaz. Testing the order of a model. Annals of Statistics, 34(3):1166–1203, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1152540746.

Chatterjee:2007:ESG

[Cha07] Sourav Chatterjee. Estimation in spin glasses: A first step. Annals of Statistics, 35(5):1931–1946, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461717.

Chan:2009:DSC

[Cha09] Hock Peng Chan. Detection of spatial clustering with average likelihood ratio test statistics. Annals of Statistics, 37(6B): 3985–4010, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303534.

Chen:2003:CSR

[CHC03] Dechang Chen, Peng Huang, and Xiuzhen Cheng. A concrete statistical realization of Kleinberg’s stochastic discrimination for pattern recognition. Part I. Two-class classification. Annals of Statistics, 31(5):1393–1413, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1065705112.

Chen:2000:OSD

[Che00] Kani Chen. Optimal sequential designs of case-control studies. Annals of Statistics, 28(5):1452–1471, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015957402.

Chen:2004:SEP

[Che04] Kani Chen. Statistical estimation in the proportional hazards model with risk set sampling. Annals of Statistics, 32(4): 1513–1532, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626177.

Chernozhukov:2005:EQR

[Che05] Victor Chernozhukov. Extremal quantile regression. Annals of Statistics, 33(2):806–839, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114337.

Chi:2007:PFC

[Chi07] Zhiyi Chi. On the performance of FDR control: Constraints and a partial solution. Annals of Statistics, 35(4):1409–1431, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1188405616.

Chopin:2004:CLT

[Cho04] Nicolas Chopin. Central limit theorem for sequential Monte Carlo methods and its application to Bayesian inference. Annals of Statistics, 32(6):2385–2411, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794873.

Choi:2000:DSM

[CHR00] Edwin Choi, Peter Hall, and Valentin Rousson. Data sharpening methods for bias reduction in nonparametric regression. Annals of Statistics, 28(5):1339–1355, October 2000. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015957396.

Chen:2009:NPC

[CHS09] Xiaohong Chen, Lars Peter Hansen, and José Scheinkman. Nonlinear principal components and long-run implications of multivariate diffusions. Annals of Statistics, 37(6B):4279–4312, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1256303544.

Chen:2008:SEG

[CHT08] Xiaohong Chen, Han Hong, and Alessandro Tarozzi. Semiparametric efficiency in GMM models with auxiliary data. Annals of Statistics, 36(2):808–843, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1205420520.

Cai:2007:ECS

[CJL07] T. Tony Cai, Jiashun Jin, and Mark G. Low. Estimation and confidence sets for sparse normal mixtures. Annals of Statistics, 35(6):2421–2449, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012967.

Chen:2006:DEE

[CK06] Ming-Hui Chen and Sungduk Kim. Discussion of “Equi-energy sampler” by Kou, Zhou and Wong. Annals of Statistics, 34(4): 1629–1635, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567624.

Cheng:2008:GFP

[CK08a] Guang Cheng and Michael R. Kosorok. General frequentist properties of the posterior profile distribution. Annals of Statistics, 36 (4):1819–1853, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237301.

Cheng:2008:HOS

[CK08b] Guang Cheng and Michael R. Kosorok. Higher order semiparametric frequentist inference with the profile sampler. Annals of Statistics, 36(4):1786–1818, August 2008. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237300.

Cohen:2002:BEG

[CKS02] Arthur Cohen, J. H. B. Kemperman, and Harold Sackrowitz. On the bias in estimating genetic length and other quantities in simplex constrained models. Annals of Statistics, 30(1):202–219, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015362190.

Crambes:2009:SSE

[CKS09] Christophe Crambes, Alois Kneip, and Pascal Sarda. Smoothing splines estimators for functional linear regression. Annals of Statistics, 37(1):35–72, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115927.

Chan:2000:AAE

[CL00] Hock Peng Chan and Tze Leung Lai. Asymptotic approximations for error probabilities of sequential or fixed sample size tests in exponential families. Annals of Statistics, 28(6):1638–1669, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957474.

Cook:2002:DRC

[CL02] R. Dennis Cook and Bing Li. Dimension reduction for conditional mean in regression. Annals of Statistics, 30(2):455–474, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1021379861.

Cai:2003:NNE

[CL03] T. Tony Cai and Mark G. Low. A note on nonparametric estimation of linear functionals. Annals of Statistics, 31(4):1140–1153, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1059655908.

Cai:2004:ATN

[CL04a] T. Tony Cai and Mark G. Low. An adaptation theory for nonparametric confidence intervals. Annals of Statistics, 32(5):

1805–1840, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883773.

Cai:2004:MEL

[CL04b] T. Tony Cai and Mark G. Low. Minimax estimation of linear functionals over nonconvex parameter spaces. Annals of Statistics, 32(2):552–576, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1083178938.

Cook:2004:DDI

[CL04c] R. Dennis Cook and Bing Li. Determining the dimension of iterative Hessian transformation. Annals of Statistics, 32(6): 2501–2531, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1107794877.

Cai:2005:NES

[CL05a] T. Tony Cai and Mark G. Low. Nonparametric estimation over shrinking neighborhoods: Superefficiency and adaptation. Annals of Statistics, 33(1):184–213, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1112967704.

Cai:2005:NEQ

[CL05b] T. Tony Cai and Mark G. Low. Nonquadratic estimators of a quadratic functional. Annals of Statistics, 33(6):2930–2956, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1140191679.

Cai:2005:AEL

[CL05c] T. Tony Cai and Mark G. Low. On adaptive estimation of linear functionals. Annals of Statistics, 33(5):2311–2343, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936565.

Cai:2006:ACB

[CL06a] T. Tony Cai and Mark G. Low. Adaptive confidence balls. Annals of Statistics, 34(1):202–228, February 2006. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1146576261.

Cai:2006:OAE

[CL06b] T. Tony Cai and Mark G. Low. Optimal adaptive estimation of a quadratic functional. Annals of Statistics, 34(5):2298–2325, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1169571798.

Coram:2006:CBE

[CL06c] Marc Coram and Steven P. Lalley. Consistency of Bayes estimators of a binary regression function. Annals of Statistics, 34 (3):1233–1269, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540748.

Cai:2007:DDS

[CL07a] T. Tony Cai and Jinchi Lv. Discussion: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6):2365–2369, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012961.

Chan:2007:STR

[CL07b] Hock Peng Chan and Wei-Liem Loh. Some theoretical results on neural spike train probability models. Annals of Statistics, 35(6): 2691–2722, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201012977.

Chan:2008:REP

[CL08] Ngai Hang Chan and Shiqing Ling. Residual empirical processes for long and short memory time series. Annals of Statistics, 36 (5):2453–2470, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1223908099. See correction [CL10].

Chen:2009:HTN

[CL09] Jiahua Chen and Pengfei Li. Hypothesis test for normal mixture models: The EM approach. Annals of Statistics, 37(5A): 2523–2542, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247663764.

Chan:2010:CRE

[CL10] Ngai Hang Chan and Shiqing Ling. Correction: “Residual empirical processes for long and short memory time series”. Annals of Statistics, 38(6):3839, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1291126976. See [CL08].

Chatterjee:2008:PBA

[CLL08] Snigdhansu Chatterjee, Partha Lahiri, and Huilin Li. Parametric bootstrap approximation to the distribution of EBLUP and related prediction intervals in linear mixed models. Annals of Statistics, 36(3):1221–1245, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1211819562.

Clemencon:2006:DLR

[CLV06] Stéphan Clémençon, Gábor Lugosi, and Nicolas Vayatis. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2672–2676, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1179935058.

Clemencon:2008:REM

[CLV08] Stéphan Clémençon, Gábor Lugosi, and Nicolas Vayatis. Ranking and empirical minimization of U-statistics. Annals of Statistics, 36(2):844–874, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1205420521.

Chaudhuri:2000:SSV

[CM00a] Probal Chaudhuri and J. S. Marron. Scale space view of curve estimation. Annals of Statistics, 28(2):408–428, April 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1016218224.

Cifarelli:2000:SNR

[CM00b] Donato Michele Cifarelli and Eugenio Melilli. Some new results for Dirichlet priors. Annals of Statistics, 28(5):1390–1413, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957399.

Cheng:2001:BRF

[CM01] Ching-Shui Cheng and Rahul Mukerjee. Blocked regular fractional factorial designs with maximum estimation capacity. Annals of Statistics, 29(2):530–548, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1009210551. See correction

[CM02].

Cheng:2002:CBR

[CM02] Ching-Shui Cheng and Rahul Mukerjee. Correction: “Blocked regular fractional factorial designs with maximum estimation capacity”. Annals of Statistics, 30(3):925–926, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1028674846. See [CM01].

Clerc:2003:EDS

[CM03] Maureen Clerc and Stéphane Mallat. Estimating deformations of stationary processes. Annals of Statistics, 31(6):1772–1821, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1074290327.

Craiu:2005:MPA

[CM05] Radu V. Craiu and Xiao-Li Meng. Multiprocess parallel antithetic coupling for backward and forward Markov chain Monte Carlo. Annals of Statistics, 33(2):661–697, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1117114333.

Coeurjolly:2008:HEE

[Coe08] Jean-François Coeurjolly. Hurst exponent estimation of locally self-similar Gaussian processes using sample quantiles. Annals of Statistics, 36(3):1404–1434, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1211819569.

Cook:2004:TPC

[Coo04] R. Dennis Cook. Testing predictor contributions in sufficient dimension reduction. Annals of Statistics, 32(3):1062–1092, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1085408495.

Carter:2004:TIR

[CP04] Andrew Carter and David Pollard. Tusnády’s inequality revisited. Annals of Statistics, 32(6):2731–2741, December 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1107794885.

Candes:2009:NIM

[CP09] Emmanuel J. Candès and Yaniv Plan. Near-ideal model selection by ℓ_{1} minimization. Annals of Statistics, 37(5A):2145–2177, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663751.

Cheng:2007:RVU

[CPW07] Ming-Yen Cheng, Liang Peng, and Jyh-Shyang Wu. Reducing variance in univariate smoothing. Annals of Statistics, 35 (2):522–542, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1183667283.

Chang:2001:ELR

[CR01] Ted Chang and Louis-Paul Rivest. M-estimation for location and regression parameters in group models: a case study using Stiefel manifolds. Annals of Statistics, 29(3):784–814, June 2001. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1009210690.

Carroll:2006:D

[CR06] Raymond J. Carroll and David Ruppert. Discussion. Annals of Statistics, 34(5):2098–2104, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571787.

Chambaz:2008:BBO

[CR08] Antoine Chambaz and Judith Rousseau. Bounds for Bayesian order identification with application to mixtures. Annals of Statistics, 36(2):938–962, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1205420524.

Csiszar:2000:CBM

[CS00] Imre Csiszár and Paul C. Shields. The consistency of the BIC Markov order estimator. Annals of Statistics, 28(6):1601– 1619, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957472.

Cohen:2005:CBP

[CS05a] Arthur Cohen and Harold B. Sackrowitz. Characterization of Bayes procedures for multiple endpoint problems and inadmissibility of the step-up procedure. Annals of Statistics, 33(1): 145–158, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1112967702.

Cohen:2005:DTR

[CS05b] Arthur Cohen and Harold B. Sackrowitz. Decision theory results for one-sided multiple comparison procedures. Annals of Statistics, 33(1):126–144, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1112967701.

Csorgo:2006:SIP

[CSW06] Miklós Csörgő, Barbara Szyszkowicz, and Lihong Wang. Strong invariance principles for sequential Bahadur–Kiefer and Vervaat error processes of long-range dependent sequences. Annals of Statistics, 34(2):1013–1044, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418250.

Csorgo:2007:SIP

[CSW07] Miklós Csörgő, Barbara Szyszkowicz, and Lihong Wang. Strong invariance principles for sequential Bahadur–Kiefer and Vervaat error processes of long-range dependent sequences. Annals of Statistics, 35(6):2815–2817, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012981.

Cohen:2009:NMT

[CSX09] Arthur Cohen, Harold B. Sackrowitz, and Minya Xu. A new multiple testing method in the dependent case. Annals of Statistics, 37(3):1518–1544, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369030.

Chen:2002:IFM

[CT02] Zhiqiang Chen and David E. Tyler. The influence function and maximum bias of Tukey’s median. Annals of Statistics, 30(6): 1737–1759, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1043351255.

Cheng:2005:GTM

[CT05] Ching-Shui Cheng and Boxin Tang. A general theory of minimum aberration and its applications. Annals of Statistics, 33 (2):944–958, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1117114341.

Csiszar:2006:CEB

[CT06] Imre Csiszár and Zsolt Talata. Consistent estimation of the basic neighborhood of Markov random fields. Annals of Statistics, 34 (1):123–145, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576258.

Candes:2007:DSS

[CT07a] Emmanuel Candes and Terence Tao. The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6):2313–2351, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012958.

Candes:2007:RDS

[CT07b] Emmanuel Candès and Terence Tao. Rejoinder: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6):2392–2404, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1201012965.

Consonni:2003:ECR

[CV03] Guido Consonni and Piero Veronese. Enriched conjugate and reference priors for the Wishart family on symmetric cones. Annals of Statistics, 31(5):1491–1516, October 2003. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1065705116.

Claeskens:2003:BCB

[CvK03] Gerda Claeskens and Ingrid van Keilegom. Bootstrap confidence bands for regression curves and their derivatives. Annals of Statistics, 31(6):1852–1884, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290329.

Chan:2004:EFD

[CW04] Grace Chan and Andrew T. A. Wood. Estimation of fractal dimension for a class of non-Gaussian stationary processes and fields. Annals of Statistics, 32(3):1222–1260, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1085408501.

Cai:2008:AVF

[CW08] T. Tony Cai and Lie Wang. Adaptive variance function estimation in heteroscedastic nonparametric regression. Annals of Statistics, 36(5):2025–2054, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908083.

Chamandy:2008:TEC

[CWTG08] N. Chamandy, K. J. Worsley, J. Taylor, and F. Gosselin. Tilted Euler characteristic densities for central limit random fields, with application to “bubbles”. Annals of Statistics, 36(5):2471–2507, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1223908100.

Chen:2009:EEC

[CWY09] Xiaohong Chen, Wei Biao Wu, and Yanping Yi. Efficient estimation of copula-based semiparametric Markov models. Annals of Statistics, 37(6B):4214–4253, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1256303542.

Clarke:2001:MCT

[CY01] B. Clarke and A. Yuan. Manifest characterization and testing for certain latent properties. Annals of Statistics, 29(3):876–898, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1009210693.

Cheng:2004:GIM

[CY04] Shao-Wei Cheng and Kenny Q. Ye. Geometric isomorphism and minimum aberration for factorial designs with quantitative factors. Annals of Statistics, 32(5):2168–2185, October 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1098883786.

Clarke:2006:CFE

[CY06] B. Clarke and Ao Yuan. Closed form expressions for Bayesian sample size. Annals of Statistics, 34(3):1293–1330, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1152540750.

Chang:2009:NSD

[CY09] Hua-Hua Chang and Zhiliang Ying. Nonlinear sequential designs for logistic item response theory models with applications to computerized adaptive tests. Annals of Statistics, 37(3):1466–1488, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1239369028.

Chen:2007:LPL

[CZ07] Songnian Chen and Lingzhi Zhou. Local partial likelihood estimation in proportional hazards regression. Annals of Statistics, 35(2):888–916, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1183667298.

Cai:2009:AEA

[CZ09a] T. Tony Cai and Harrison H. Zhou. Asymptotic equivalence and adaptive estimation for robust nonparametric regression. Annals of Statistics, 37(6A):3204–3235, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515385.

Cai:2009:DDB

[CZ09b] T. Tony Cai and Harrison H. Zhou. A data-driven block thresholding approach to wavelet estimation. Annals of Statistics, 37 (2):569–595, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693142.

Dahlhaus:2000:LAL

[Dah00] Rainer Dahlhaus. A likelihood approximation for locally stationary processes. Annals of Statistics, 28(6):1762–1794, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957480.

Dahlhaus:2006:EPE

[Dah06] Rainer Dahlhaus. Efficient parameter estimation for self-similar processes. Annals of Statistics, 34(2):1045–1047, April 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1151418251.

Dalalyan:2005:SAE

[Dal05] Arnak Dalalyan. Sharp adaptive estimation of the drift function for ergodic diffusions. Annals of Statistics, 33(6):2507–2528, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1140191664.

Dempster:2002:JWT

[Dem02] A. P. Dempster. John w. Tukey as “philosopher”. Annals of Statistics, 30(6):1619–1628, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1043351249;
 http://www.jstor.org/stable/1558732.

Dette:2000:COD

[DF00] Holger Dette and Tobias Franke. Constrained D- and D_{1}-optimal designs for polynomial regression. Annals of Statistics, 28(6): 1702–1727, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957477.

Dette:2001:RDP

[DF01] Holger Dette and Tobias Franke. Robust designs for polynomial regression by maximizing a minimum of D- and D_{1}-efficiencies. Annals of Statistics, 29(4):1024–1049, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1013699990.

Dmitrienko:2000:SCR

[DG00] A. Dmitrienko and Z. Govindarajulu. Sequential confidence regions for maximum likelihood estimates. Annals of Statistics, 28 (5):1472–1501, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957403.

Davies:2005:BG

[DG05] P. Laurie Davies and Ursula Gather. Breakdown and groups. Annals of Statistics, 33(3):977–1035, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1120224093.

Davies:2006:ADB

[DG06] P. Laurie Davies and Ursula Gather. Addendum to the discussion of “breakdown and groups”. Annals of Statistics, 34 (3):1577–1579, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540760.

Drton:2012:CMM

[DG12] Mathias Drton and Aldo Goia. Correction on moments of minors of Wishart matrices. Annals of Statistics, 40(2):1283–1284, April 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1342625469. See [DMO08].

Dupuy:2006:ATC

[DGM06] Jean-François Dupuy, Ion Grama, and Mounir Mesbah. Asymptotic theory for the Cox model with missing time-dependent covariate. Annals of Statistics, 34(2):903–924, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1151418245.

Douc:2007:CAM

[DGMR07] R. Douc, A. Guillin, J.-M. Marin, and C. P. Robert. Convergence of adaptive mixtures of importance sampling schemes. Annals of Statistics, 35(1):420–448, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1181100193.

Datta:2000:BPA

[DGMS00] Gauri Sankar Datta, Malay Ghosh, Rahul Mukerjee, and Trevor J. Sweeting. Bayesian prediction with approximate frequentist validity. Annals of Statistics, 28(5):1414–1426, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957400.

Dalalyan:2006:PML

[DGT06] A. S. Dalalyan, G. K. Golubev, and A. B. Tsybakov. Penalized maximum likelihood and semiparametric second-order efficiency. Annals of Statistics, 34(1):169–201, February 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1146576260.

Dudley:2002:ANS

[DH02] R. M. Dudley and D. Haughton. Asymptotic normality with small relative errors of posterior probabilities of half-spaces. Annals of Statistics, 30(5):1311–1344, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1035844978.

deHaan:2003:WCE

[dHL03] Laurens de Haan and Tao Lin. Weak consistency of extreme value estimators in C[0, 1]. Annals of Statistics, 31(6):1996–2012, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1074290334.

Dette:2009:GCC

[DHL09] Holger Dette and Tim Holland-Letz. A geometric characterization of c-optimal designs for heteroscedastic regression. Annals of Statistics, 37(6B):4088–4103, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1256303537.

Delaigle:2007:ACN

[DHM07] Aurore Delaigle, Peter Hall, and Hans-Georg Müller. Accelerated convergence for nonparametric regression with coarsened predictors. Annals of Statistics, 35(6):2639–2653, December 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1201012975.

Delaigle:2008:DRM

[DHM08] Aurore Delaigle, Peter Hall, and Alexander Meister. On deconvolution with repeated measurements. Annals of Statistics, 36 (2):665–685, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1205420515.

Draper:2000:KOS

[DHP00] Norman R. Draper, Berthold Heiligers, and Friedrich Pukelsheim. Kiefer ordering of simplex designs for second-degree mixture models with four or more ingredients. Annals of Statistics, 28 (2):578–590, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016218231.

deHaan:2006:SEM

[dHP06] Laurens de Haan and Teresa T. Pereira. Spatial extremes: Models for the stationary case. Annals of Statistics, 34(1):146–168, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1146576259.

Delgado:2005:DFG

[DHV05] Miguel A. Delgado, Javier Hidalgo, and Carlos Velasco. Distribution free goodness-of-fit tests for linear processes. Annals of Statistics, 33(6):2568–2609, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191667.

Dippon:2003:ARS

[Dip03] Jürgen Dippon. Accelerated randomized stochastic optimization. Annals of Statistics, 31(4):1260–1281, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1059655913.

Donoho:2004:HCD

[DJ04] David Donoho and Jiashun Jin. Higher criticism for detecting sparse heterogeneous mixtures. Annals of Statistics, 32(3):962– 994, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1085408492.

Donoho:2006:AMF

[DJ06] David Donoho and Jiashun Jin. Asymptotic minimaxity of false discovery rate thresholding for sparse exponential data. Annals of Statistics, 34(6):2980–3018, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935072.

Das:2004:MSE

[DJR04] Kalyan Das, Jiming Jiang, and J. N. K. Rao. Mean squared error of empirical predictor. Annals of Statistics, 32(2):818–840, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1083178948.

Davies:2001:LER

[DK01] P. L. Davies and A. Kovac. Local extremes, runs, strings and multiresolution. Annals of Statistics, 29(1):1–65, February 2001. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/996986501.

Davies:2004:DSD

[DK04] P. Laurie Davies and Arne Kovac. Densities, spectral densities and modality. Annals of Statistics, 32(3):1093–1136, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1085408496.

Davies:2009:NRC

[DKM09] P. L. Davies, A. Kovac, and M. Meise. Nonparametric regression, confidence regions and regularization. Annals of Statistics, 37(5B):2597–2625, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247836662.

Ducharme:2003:EAN

[DL03] Gilles R. Ducharme and Teresa Ledwina. Efficient and adaptive nonparametric test for the two-sample problem. Annals of Statistics, 31(6):2036–2058, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290336.

Delgado:2001:STN

[DM01] Miguel A. Delgado and Wenceslao González Manteiga. Significance testing in nonparametric regression based on the bootstrap. Annals of Statistics, 29(5):1469–1507, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1013203462.

Dette:2003:ODE

[DM03] Holger Dette and Viatcheslav B. Melas. Optimal designs for estimating individual coefficients in Fourier regression models. Annals of Statistics, 31(5):1669–1692, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1065705122.

Douc:2008:LTW

[DM08] Randal Douc and Eric Moulines. Limit theorems for weighted samples with applications to sequential Monte Carlo methods. Annals of Statistics, 36(5):2344–2376, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1223908095.

Desolneux:2003:MME

[DMM03] Agnès Desolneux, Lionel Moisan, and Jean-Michel Morel. Maximal meaningful events and applications to image analysis. Annals of Statistics, 31(6):1822–1851, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1074290328.

Drton:2008:MMW

[DMO08] Mathias Drton, Hélène Massam, and Ingram Olkin. Moments of minors of Wishart matrices. Annals of Statistics, 36(5): 2261–2283, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1223908092. See correction [DG12].

Dette:2004:ODC

[DMP04] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal designs for a class of nonlinear regression models. Annals of Statistics, 32(5):2142–2167, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1098883785.

Dette:2005:ODT

[DMP05] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal designs for three-dimensional shape analysis with spherical harmonic descriptors. Annals of Statistics, 33(6):2758–2788, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1140191672.

Douc:2004:APM

[DMR04] Randal Douc, Éric Moulines, and Tobias Rydén. Asymptotic properties of the maximum likelihood estimator in autoregressive models with Markov regime. Annals of Statistics, 32(5): 2254–2304, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883789.

Dette:2001:NAC

[DN01] Holger Dette and Natalie Neumeyer. Nonparametric analysis of covariance. Annals of Statistics, 29(5):1361–1400, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1013203458.

Dahlhaus:2006:NQM

[DP06] Rainer Dahlhaus and Wolfgang Polonik. Nonparametric quasimaximum likelihood estimation for Gaussian locally stationary processes. Annals of Statistics, 34(6):2790–2824, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1179935065.

Dahlhaus:2006:SIT

[DR06a] Rainer Dahlhaus and Suhasini Subba Rao. Statistical inference for time-varying ARCH processes. Annals of Statistics, 34 (3):1075–1114, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540743.

Diaconis:2006:BAR

[DR06b] Persi Diaconis and Silke W. W. Rolles. Bayesian analysis for reversible Markov chains. Annals of Statistics, 34(3):1270–1292, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1152540749.

Drees:2000:HMH

[DRdH00] Holger Drees, Sidney Resnick, and Laurens de Haan. How to make a Hill plot. Annals of Statistics, 28(1):254–274, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1016120372.

Drees:2001:MRB

[Dre01] Holger Drees. Minimax risk bounds in extreme value theory. Annals of Statistics, 29(1):266–294, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/996986509.

Drton:2009:LRT

[Drt09] Mathias Drton. Likelihood ratio tests and singularities. Annals of Statistics, 37(2):979–1012, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693157.

Dryden:2005:SAH

[Dry05] Ian L. Dryden. Statistical analysis on high-dimensional spheres and shape spaces. Annals of Statistics, 33(4):1643–1665, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1123250225.

Dumbgen:2001:MTQ

[DS01] Lutz Dümbgen and Vladimir G. Spokoiny. Multiscale testing of qualitative hypotheses. Annals of Statistics, 29(1):124–152, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/996986504.

Dey:2002:OFF

[DS02] Aloke Dey and Chung-Yi Suen. Optimal fractional factorial plans for main effects and specified two-factor interactions: a projective geometric approach. Annals of Statistics, 30(5):1512–1523, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844986.

Dudley:2009:DFL

[DSW09] R. M. Dudley, Sergiy Sidenko, and Zuoqin Wang. Differentiability of t-functionals of location and scatter. Annals of Statistics, 37(2):939–960, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693155.

Dette:2009:ODD

[DT09] Holger Dette and Stefanie Titoff. Optimal discrimination designs. Annals of Statistics, 37(4):2056–2082, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1245332840.

Durot:2007:EMC

[Dur07] Cécile Durot. On the L_{p}-error of monotonicity constrained estimators. Annals of Statistics, 35(3):1080–1104, July 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1185303999.

Dumbgen:2008:MIA

[DW08] Lutz Dümbgen and Günther Walther. Multiscale inference about a density. Annals of Statistics, 36(4):1758–1785, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1216237299.

Du:2009:FDA

[DZM09] Juan Du, Hao Zhang, and V. S. Mandrekar. Fixed-domain asymptotic properties of tapered maximum likelihood estimators. Annals of Statistics, 37(6A):3330–3361, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515389.

Einmahl:2006:WAT

[EdHL06] John H. J. Einmahl, Laurens de Haan, and Deyuan Li. Weighted approximations of tail copula processes with application to testing the bivariate extreme value condition. Annals of Statistics, 34 (4):1987–2014, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567640.

Efron:2001:SCS

[Efr01] Bradley Efron. Selection criteria for scatterplot smoothers. Annals of Statistics, 29(2):470–505, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1009210549.

Efron:2003:REB

[Efr03] Bradley Efron. Robbins, empirical Bayes and microarrays. Annals of Statistics, 31(2):366–378, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1051027871.

Efromovich:2004:DEB

[Efr04] Sam Efromovich. Density estimation for biased data. Annals of Statistics, 32(3):1137–1161, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1085408497.

Efromovich:2005:EDR

[Efr05] Sam Efromovich. Estimation of the density of regression errors. Annals of Statistics, 33(5):2194–2227, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1132936561.

Efromovich:2007:CDE

[Efr07a] Sam Efromovich. Conditional density estimation in a regression setting. Annals of Statistics, 35(6):2504–2535, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1201012970.

Efron:2007:SPF

[Efr07b] Bradley Efron. Size, power and false discovery rates. Annals of Statistics, 35(4):1351–1377, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405614.

Efromovich:2008:AEO

[Efr08] Sam Efromovich. Adaptive estimation of and oracle inequalities for probability densities and characteristic functions. Annals of Statistics, 36(3):1127–1155, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1211819559.

Eaton:2008:EFP

[EHJL08] Morris L. Eaton, James P. Hobert, Galin L. Jones, and Wen-Lin Lai. Evaluation of formal posterior distributions via Markov chain arguments. Annals of Statistics, 36(5):2423–2452, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1223908098.

Efron:2004:LAR

[EHJT04] Bradley Efron, Trevor Hastie, Iain Johnstone, and Robert Tibshirani. Least angle regression. Annals of Statistics, 32(2): 407–499, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178935.

Efron:2007:DDS

[EHT07] Bradley Efron, Trevor Hastie, and Robert Tibshirani. Discussion: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6):2358–2364, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201012960.

Eggermont:2000:MLE

[EL00] P. P. B. Eggermont and V. N. LaRiccia. Maximum likelihood estimation of smooth monotone and unimodal densities. Annals of Statistics, 28(3):922–947, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015952005.

Einmahl:2006:ANE

[EL06] John H. J. Einmahl and Tao Lin. Asymptotic normality of extreme value estimators on C[0, 1]. Annals of Statistics, 34(1): 469–492, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576271.

Einmahl:2005:UBC

[EM05] Uwe Einmahl and David M. Mason. Uniform in bandwidth consistency of kernel-type function estimators. Annals of Statistics, 33(3):1380–1403, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1120224106.

Einmahl:2001:NES

[EPdH01] John H. J. Einmahl, Vladimir I. Piterbarg, and Laurens de Haan. Nonparametric estimation of the spectral measure of an extreme value distribution. Annals of Statistics, 29(5):1401–1423, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203459.

Einmahl:2009:MEL

[ES09] John H. J. Einmahl and Johan Segers. Maximum empirical likelihood estimation of the spectral measure of an extremevalue distribution. Annals of Statistics, 37(5B):2953–2989, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247836674.

Fukumizu:2009:KDR

[FBJ09] Kenji Fukumizu, Francis R. Bach, and Michael I. Jordan. Kernel dimension reduction in regression. Annals of Statistics, 37(4): 1871–1905, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1245332835.

Fan:2002:CGL

[FCZ02] J. Fan, Zhang C, and J. Zhang. Correction: “Generalized likelihood ratio statistics and Wilks phenomenon”. Annals of Statistics, 30(6):1811, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1043351258. See [FZZ01].

Finner:2007:DFD

[FDR07] Helmut Finner, Thorsten Dickhaus, and Markus Roters. Dependency and false discovery rate: Asymptotics. Annals of Statistics, 35(4):1432–1455, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405617.

Finner:2009:FDR

[FDR09] Helmut Finner, Thorsten Dickhaus, and Markus Roters. On the false discovery rate and an asymptotically optimal rejection curve. Annals of Statistics, 37(2):596–618, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1236693143.

Fan:2008:HDC

[FF08] Jianqing Fan and Yingying Fan. High-dimensional classification using features annealed independence rules. Annals of Statistics, 36(6):2605–2637, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1231165181.

Frigessi:2000:ACT

[FGR00] Arnoldo Frigessi, Jørund Gåsemyr, and Håvard Rue. Antithetic coupling of two Gibbs sampler chains. Annals of Statistics, 28 (4):1128–1149, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956710.

Fuh:2000:AES

[FH00] Cheng-Der Fuh and Inchi Hu. Asymptotically efficient strategies for a stochastic scheduling problem with order constraints. Annals of Statistics, 28(6):1670–1695, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015957475.

Fang:2008:LDO

[FH08] X. Fang and A. S. Hedayat. Locally D-optimal designs based on a class of composed models resulted from blending Emax and one-compartment models. Annals of Statistics, 36(1):428–444, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877308.

Fang:2009:PEL

[FHS09] Fang Fang, Quan Hong, and Jun Shao. A pseudo empirical likelihood approach for stratified samples with nonresponse. Annals of Statistics, 37(1):371–393, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115939.

Friedman:2000:ALR

[FHT00] Jerome Friedman, Trevor Hastie, and Robert Tibshirani. Additive logistic regression: a statistical view of boosting (with discussion and a rejoinder by the authors). Annals of Statistics, 28(2):337–407, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016218223.

Fu:2000:ALT

[FK00] Wenjiang Fu and Keith Knight. Asymptotics for lasso-type estimators. Annals of Statistics, 28(5):1356–1378, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1015957397.

Fan:2002:VSC

[FL02] Jianqing Fan and Runze Li. Variable selection for Cox’s proportional hazards model and frailty model. Annals of Statistics, 30 (1):74–99, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015362185.

Fromont:2006:AGF

[FL06] Magalie Fromont and Béatrice Laurent. Adaptive goodness-of-fit tests in a density model. Annals of Statistics, 34(2):680–720, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1151418237.

Fan:2006:LPL

[FLZ06] Jianqing Fan, Huazhen Lin, and Yong Zhou. Local partiallikelihood estimation for lifetime data. Annals of Statistics, 34 (1):290–325, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576264.

Fort:2003:CMC

[FM03] Gersende Fort and Eric Moulines. Convergence of the Monte Carlo expectation maximization for curved exponential families. Annals of Statistics, 31(4):1220–1259, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1059655912.

Freund:2004:GBA

[FMS04] Yoav Freund, Yishay Mansour, and Robert E. Schapire. Generalization bounds for averaged classifiers. Annals of Statistics, 32 (4):1698–1722, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626184.

Fan:2004:NPL

[FP04] Jianqing Fan and Heng Peng. Nonconcave penalized likelihood with a diverging number of parameters. Annals of Statistics, 32(3):928–961, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408491.

Findley:2001:UCS

[FPW01] David F. Findley, Benedikt M. Pötscher, and Ching-Zong Wei. Uniform convergence of sample second moments of families of time series arrays. Annals of Statistics, 29(3):815–838, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1009210691.

Finner:2002:MHT

[FR02] H. Finner and M. Roters. Multiple hypotheses testing and expected number of Type I errors. Annals of Statistics, 30(1): 220–238, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015362191.

Friedman:2001:GFA

[Fri01] Jerome H. Friedman. Greedy function approximation: A gradient boosting machine. Annals of Statistics, 29(5):1189–1232, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203451.

Finner:2002:PPP

[FS02a] H. Finner and K. Strassburger. The partitioning principle: a powerful tool in multiple decision theory. Annals of Statistics, 30 (4):1194–1213, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031689023.

Friedman:2002:JWT

[FS02b] Jerome H. Friedman and Werner Stuetzle. John w. Tukey’s work on interactive graphics. Annals of Statistics, 30(6):1629–1639, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1043351250;   http://www.jstor.org/stable/
1558733.

Friedlander:2007:DDS

[FS07] Michael P. Friedlander and Michael A. Saunders. Discussion: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6):2385–2391, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201012964.

Fryzlewicz:2008:NLS

[FSR08] Piotr Fryzlewicz, Theofanis Sapatinas, and Suhasini Subba Rao. Normalized least-squares estimation in time-varying ARCH models. Annals of Statistics, 36(2):742–786, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1205420518.

Fuh:2003:SCH

[Fuh03] Cheng-Der Fuh. SPRT and CUSUM in hidden Markov models. Annals of Statistics, 31(3):942–977, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1056562468.

Fuh:2004:AOC

[Fuh04] Cheng-Der Fuh. Asymptotic operating characteristics of an optimal change point detection in hidden Markov models. Annals of Statistics, 32(5):2305–2339, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1098883790.

Fuh:2006:ELE

[Fuh06] Cheng-Der Fuh. Efficient likelihood estimation in state space models. Annals of Statistics, 34(4):2026–2068, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1162567642. See comments [Jen10] and reply [Fuh10].

Fuh:2010:RSP

[Fuh10] Cheng-Der Fuh. Reply to “On some problems in the article Efficient Likelihood Estimation in State Space Models” by Cheng-Der Fuh [Ann. Statist. 34 (2006) 2026–2068]. Annals of Statistics, 38(2):1282–1285, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1266586631. See [Fuh06,

Jen10].

Fukumizu:2003:LRU

[Fuk03] Kenji Fukumizu. Likelihood ratio of unidentifiable models and multilayer neural networks. Annals of Statistics, 31(3):833–851, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562464.

Fan:2009:LQL

[FWF09] Jianqing Fan, Yichao Wu, and Yang Feng. Local quasi-likelihood with a parametric guide. Annals of Statistics, 37(6B):4153– 4183, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303540.

Fraiman:2001:ORM

[FYZ01] Ricardo Fraiman, Víctor J. Yohai, and Ruben H. Zamar. Optimal robust M-estimates of location. Annals of Statistics, 29 (1):194–223, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/996986506.

Fan:2004:SEL

[FZ04] Jianqing Fan and Jian Zhang. Sieve empirical likelihood ratio tests for nonparametric functions. Annals of Statistics, 32(5): 1858–1907, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883775.

Ferreira:2006:BHM

[FZ06] J. A. Ferreira and A. H. Zwinderman. On the Benjamini– Hochberg method. Annals of Statistics, 34(4):1827–1849, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1162567635.

Fan:2001:GLR

[FZZ01] Jianqing Fan, Chunming Zhang, and Jian Zhang. Generalized likelihood ratio statistics and Wilks phenomenon. Annals of Statistics, 29(1):153–193, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/996986505. See correction

[FCZ02].

Gavrilov:2009:ASP

[GBS09] Yulia Gavrilov, Yoav Benjamini, and Sanat K. Sarkar. An adaptive step-down procedure with proven FDR control under independence. Annals of Statistics, 37(2):619–629, April 2009. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1236693144.

Gemai:2001:WCE

[GCL01] Gemai, Chen, and Richard A. Lockhart. Weak convergence of the empirical process of residuals in linear models with many parameters. Annals of Statistics, 29(3):748–762, June 2001. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1009210688.

Genon-Catalot:2002:AEE

[GCLN02] Valentine Genon-Catalot, Catherine Laredo, and Michael Nussbaum. Asymptotic equivalence of estimating a Poisson intensity and a positive diffusion drift. Annals of Statistics, 30(3):731–753, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1028674840.

Grunwald:2004:GTM

[GD04] Peter D. Grünwald and A. Philip Dawid. Game theory, maximum entropy, minimum discrepancy and robust Bayesian decision theory. Annals of Statistics, 32(4):1367–1433, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1091626173.

Garcia-Escudero:2008:GTA

[GEGMMI08] Luis A. García-Escudero, Alfonso Gordaliza, Carlos Matrán, and Agustin Mayo-Iscar. A general trimming approach to robust cluster analysis. Annals of Statistics, 36(3):1324–1345, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1211819566.

Gelman:2005:AVW

[Gel05] Andrew Gelman. Analysis of variance — why it is more important than ever. Annals of Statistics, 33(1):1–53, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1112967698.

Goldenshluger:2000:AMR

[GG00] Alexander Goldenshluger and Eitan Greenshtein. Asymptotically minimax regret procedures in regression model selection and the magnitude of the dimension penalty. Annals of Statistics, 28(6): 1620–1637, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957473.

Gill:2008:AGC

[GG08] Richard D. Gill and Peter D. Grünwald. An algorithmic and a geometric characterization of coarsening at random. Annals of Statistics, 36(5):2409–2422, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908097.

Ghosal:2000:CRP

[GGvdV00] Subhashis Ghosal, Jayanta K. Ghosh, and Aad W. van der Vaart. Convergence rates of posterior distributions. Annals of Statistics, 28(2):500–531, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016218228.

Geiger:2002:PPD

[GH02] Dan Geiger and David Heckerman. Parameter priors for directed acyclic graphical models and the characterization of several probability distributions. Annals of Statistics, 30(5):1412–1440, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844981.

Gloter:2007:EHP

[GH07] Arnaud Gloter and Marc Hoffmann. Estimation of the Hurst parameter from discrete noisy data. Annals of Statistics, 35(5): 1947–1974, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1194461718.

Geiger:2001:SEF

[GHKM01] Dan Geiger, David Heckerman, Henry King, and Christopher Meek. Stratified exponential families: Graphical models and model selection. Annals of Statistics, 29(2):505–529, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1009210550.

Ghosal:2001:CRD

[Gho01] Subhashis Ghosal. Convergence rates for density estimation with Bernstein polynomials. Annals of Statistics, 29(5):1264–1280, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203453.

Giraitis:2001:GEP

[GHR01] L. Giraitis, J. Hidalgo, and P. M. Robinson. Gaussian estimation of parametric spectral density with unknown pole. Annals of Statistics, 29(4):987–1023, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013699989.

Gobet:2004:NES

[GHR04] Emmanuel Gobet, Marc Hoffmann, and Markus Reiß. Nonparametric estimation of scalar diffusions based on low frequency data. Annals of Statistics, 32(5):2223–2253, October 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1098883788.

Gilbert:2000:LST

[Gil00] Peter B. Gilbert. Large sample theory of maximum likelihood estimates in semiparametric biased sampling models. Annals of Statistics, 28(1):151–194, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1016120368.

Grendar:2009:AEE

[GJ09] Marian Grendár and George Judge. Asymptotic equivalence of empirical likelihood and Bayesian MAP. Annals of Statistics, 37(5A):2445–2457, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663761.

Groeneboom:2001:CPE

[GJW01a] Piet Groeneboom, Geurt Jongbloed, and Jon A. Wellner. A canonical process for estimation of convex functions: The “invelope” of integrated Brownian motion +t^{4}. Annals of Statistics, 29(6):1620–1652, December 2001. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015345957.

Groeneboom:2001:ECF

[GJW01b] Piet Groeneboom, Geurt Jongbloed, and Jon A. Wellner. Estimation of a convex function: Characterizations and asymptotic theory. Annals of Statistics, 29(6):1653–1698, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1015345958.

Galtchouk:2001:SEP

[GK01] L. Galtchouk and V. Konev. On sequential estimation of parameters in semimartingale regression models with continuous time parameter. Annals of Statistics, 29(5):1508–1536, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203463.

Guo:2008:AIS

[GK08] Hongwen Guo and Hira L. Koul. Asymptotic inference in some heteroscedastic regression models with long memory design and errors. Annals of Statistics, 36(1):458–487, February 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1201877310.

Gao:2000:SSA

[GKK^{+}00] Fangyu Gao, Ronald Klein, Barbara Klein, Xiwu Lin, Grace Wahba, and Dong Xiang. Smoothing spline ANOVA models for large data sets with Bernoulli observations and the randomized GACV. Annals of Statistics, 28(6):1570–1600, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957471.

Gao:2009:STN

[GKLT09] Jiti Gao, Maxwell King, Zudi Lu, and Dag Tjøstheim. Specification testing in nonlinear and nonstationary time series autoregression. Annals of Statistics, 37(6B):3893–3928, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1256303531.

Gardner:2006:CAR

[GKM06] Richard J. Gardner, Markus Kiderlen, and Peyman Milanfar. Convergence of algorithms for reconstructing convex bodies and directional measures. Annals of Statistics, 34(3):1331–1374, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1152540751.

Guerre:2005:DDR

[GL05] Emmanuel Guerre and Pascal Lavergne. Data-driven rateoptimal specification testing in regression models. Annals of Statistics, 33(2):840–870, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114338.

Groeneboom:2003:KTE

[GLdW03] P. Groeneboom, H. P. Lopuhaä, and P. P. de Wolf. Kerneltype estimators for the extreme value index. Annals of Statistics, 31(6):1956–1995, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290333.

Graczyk:2003:CWD

[GLM03] Piotr Graczyk, Gérard Letac, and Hélène Massam. The complex Wishart distribution and the symmetric group. Annals of Statistics, 31(1):287–309, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1046294466.

Gao:2006:ESS

[GLT06] Jiti Gao, Zudi Lu, and Dag Tjøstheim. Estimation in semiparametric spatial regression. Annals of Statistics, 34(3):1395–1435, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1152540753.

George:2006:IMP

[GLX06] Edward I. George, Feng Liang, and Xinyi Xu. Improved minimax predictive densities under Kullback–Leibler loss. Annals of Statistics, 34(1):78–91, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1146576256.

Gu:2005:OSN

[GM05] Chong Gu and Ping Ma. Optimal smoothing in nonparametric mixed-effect models. Annals of Statistics, 33(3):1357–1379, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1120224105.

Gine:2007:LUS

[GM07] Evarist Giné and David M. Mason. On local U-statistic processes and the estimation of densities of functions of several sample variables. Annals of Statistics, 35(3):1105–1145, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1185304000.

Geiger:2006:TAG

[GMS06] Dan Geiger, Christopher Meek, and Bernd Sturmfels. On the toric algebra of graphical models. Annals of Statistics, 34(3): 1463–1492, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540755.

Groeneboom:2008:CSDa

[GMW08a] Piet Groeneboom, Marloes H. Maathuis, and Jon A. Wellner. Current status data with competing risks: Consistency and rates of convergence of the MLE. Annals of Statistics, 36(3): 1031–1063, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1211819555.

Groeneboom:2008:CSDb

[GMW08b] Piet Groeneboom, Marloes H. Maathuis, and Jon A. Wellner. Current status data with competing risks: Limiting distribution of the MLE. Annals of Statistics, 36(3):1064–1089, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1211819556.

Grama:2006:AEN

[GN06] Ion G. Grama and Michael H. Neumann. Asymptotic equivalence of nonparametric autoregression and nonparametric regression. Annals of Statistics, 34(4):1701–1732, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1162567630.

Goldenshluger:2009:UPA

[Gol09] Alexander Goldenshluger. A universal procedure for aggregating estimators. Annals of Statistics, 37(1):542–568, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1232115945.

Genovese:2009:PDG

[GPPVW09] Christopher R. Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. On the path density of a gradient field. Annals of Statistics, 37(6A):3236–3271, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1250515386.

Genest:2007:ALE

[GQR07] Christian Genest, Jean-François Quessy, and Bruno Rémillard. Asymptotic local efficiency of Cramér-von Mises tests for multivariate independence. Annals of Statistics, 35(1):166–191, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1181100185.

Gill:2001:CIC

[GR01] Richard D. Gill and James M. Robins. Causal inference for complex longitudinal data: The continuous case. Annals of Statistics, 29(6):1785–1811, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015345962.

Giraitis:2003:EES

[GR03] L. Giraitis and P. M. Robinson. Edgeworth expansions for semiparametric Whittle estimation of long memory. Annals of Statistics, 31(4):1325–1375, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1059655915.

Gallegos:2005:RMC

[GR05] María Teresa Gallegos and Gunter Ritter. A robust method for cluster analysis. Annals of Statistics, 33(1):347–380, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1112967709.

Ghosal:2006:PCG

[GR06] Subhashis Ghosal and Anindya Roy. Posterior consistency of Gaussian process prior for nonparametric binary regression. Annals of Statistics, 34(5):2413–2429, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1169571802.

Greenshtein:2006:BSS

[Gre06] Eitan Greenshtein. Best subset selection, persistence in highdimensional statistical learning and optimization under l_{1} constraint. Annals of Statistics, 34(5):2367–2386, October 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1169571800.

Grama:2008:SEO

[GS08] Ion Grama and Vladimir Spokoiny. Statistics of extremes by oracle estimation. Annals of Statistics, 36(4):1619–1648, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1216237294.

Genest:2009:RBI

[GS09] Christian Genest and Johan Segers. Rank-based inference for bivariate extreme-value copulas. Annals of Statistics, 37(5B): 2990–3022, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247836675.

Ghosal:2000:TMR

[GSvdV00] Subhashis Ghosal, Arusharka Sen, and Aad W. van der Vaart. Testing monotonicity of regression. Annals of Statistics, 28(4): 1054–1082, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956707.

Goldenshluger:2001:APE

[GT01] A. Goldenshluger and A. Tsybakov. Adaptive prediction and estimation in linear regression with infinitely many parameters. Annals of Statistics, 29(6):1601–1619, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015345956.

Goldenshluger:2006:OCP

[GTZ06] A. Goldenshluger, A. Tsybakov, and A. Zeevi. Optimal change-point estimation from indirect observations. Annals of Statistics, 34(1):350–372, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1146576266.

Ghosal:2001:ERC

[GvdV01] Subhashis Ghosal and Aad W. van der Vaart. Entropies and rates of convergence for maximum likelihood and Bayes estimation for mixtures of normal densities. Annals of Statistics, 29(5): 1233–1263, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1013203452.

Ghosal:2007:CRP

[GvdV07a] Subhashis Ghosal and Aad van der Vaart. Convergence rates of posterior distributions for noniid observations. Annals of Statistics, 35(1):192–223, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1181100186.

Ghosal:2007:PCR

[GvdV07b] Subhashis Ghosal and Aad van der Vaart. Posterior convergence rates of Dirichlet mixtures at smooth densities. Annals of Statistics, 35(2):697–723, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1183667289.

Genovese:2000:RCG

[GW00] Christopher R. Genovese and Larry Wasserman. Rates of convergence for the Gaussian mixture sieve. Annals of Statistics, 28 (4):1105–1127, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956709.

Genovese:2004:SPA

[GW04a] Christopher Genovese and Larry Wasserman. A stochastic process approach to false discovery control. Annals of Statistics, 32 (3):1035–1061, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408494.

Giglio:2004:MIS

[GW04b] Beatrice Giglio and Henry P. Wynn. Monomial ideals and the scarf complex for coherent systems in reliability theory. Annals of Statistics, 32(3):1289–1311, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1085408503.

Genovese:2005:CSN

[GW05] Christopher R. Genovese and Larry Wasserman. Confidence sets for nonparametric wavelet regression. Annals of Statistics, 33 (2):698–729, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1117114334.

Genovese:2008:ACB

[GW08] Christopher Genovese and Larry Wasserman. Adaptive confidence bands. Annals of Statistics, 36(2):875–905, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1205420522.

Guo:2000:MRA

[GX00] Sunwei Guo and Momiao Xiong. Multiple regression approach to mapping of quantitative trait loci (QTL) based on sib-pair data: a theoretical analysis. Annals of Statistics, 28(5):1245–1278, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957393.

Gervini:2002:CRF

[GY02] Daniel Gervini and Victor J. Yohai. A class of robust and fully efficient regression estimators. Annals of Statistics, 30(2):583–616, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1021379866.

Goldenshluger:2001:NBA

[GZ01] Alexander Goldenshluger and Assaf Zeevi. Nonasymptotic bounds for autoregressive time series modeling. Annals of Statistics, 29(2):417–444, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1009210547.

Goldenshluger:2004:HTE

[GZ04] Alexander Goldenshluger and Assaf Zeevi. The Hough transform estimator. Annals of Statistics, 32(5):1908–1932, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1098883776.

Goldenshluger:2006:RCB

[GZ06] Alexander Goldenshluger and Assaf Zeevi. Recovering convex boundaries from blurred and noisy observations. Annals of Statistics, 34(3):1375–1394, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1152540752.

Gorfine:2009:CCS

[GZH09] Malka Gorfine, David M. Zucker, and Li Hsu. Case-control survival analysis with a general semiparametric shared frailty model: A pseudo full likelihood approach. Annals of Statistics, 37(3):1489–1517, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369029.

Huang:2008:AVC

[HC08] Li-Shan Huang and Jianwei Chen. Analysis of variance, coefficient of determination and F -test for local polynomial regression. Annals of Statistics, 36(5):2085–2109, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1223908085.

Helland:2006:ESM

[Hel06] Inge S. Helland. Extended statistical modeling under symmetry; the link toward quantum mechanics. Annals of Statistics, 34 (1):42–77, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576255.

Hennig:2004:BPM

[Hen04] Christian Hennig. Breakdown points for maximum likelihood estimators of location–scale mixtures. Annals of Statistics, 32 (4):1313–1340, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626171.

Hall:2000:TMR

[HH00] Peter Hall and Nancy E. Heckman. Testing for monotonicity of a regression mean by calibrating for linear functions. Annals of Statistics, 28(1):20–39, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1016120363.

Hall:2001:NKR

[HH01] Peter Hall and Li-Shan Huang. Nonparametric kernel regression subject to monotonicity constraints. Annals of Statistics, 29(3):624–647, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210683.

Hall:2005:NMI

[HH05] Peter Hall and Joel L. Horowitz. Nonparametric methods for inference in the presence of instrumental variables. Annals of Statistics, 33(6):2904–2929, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191678.

Hall:2007:MCR

[HH07] Peter Hall and Joel L. Horowitz. Methodology and convergence rates for functional linear regression. Annals of Statistics, 35 (1):70–91, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1181100181.

Huang:2008:APB

[HHM08] Jian Huang, Joel L. Horowitz, and Shuangge Ma. Asymptotic properties of bridge estimators in sparse high-dimensional regression models. Annals of Statistics, 36(2):587–613, April 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1205420512.

Holst:2001:NNC

[HI01] M. Holst and A. Irle. Nearest neighbor classification with dependent training sequences. Annals of Statistics, 29(5):1424–1442, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203460.

Hidalgo:2005:SES

[Hid05] Javier Hidalgo. Semiparametric estimation for stationary processes whose spectra have an unknown pole. Annals of Statistics, 33(4):1843–1889, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250231.

Hall:2008:PHC

[HJ08] Peter Hall and Jiashun Jin. Properties of higher criticism under strong dependence. Annals of Statistics, 36(1):381–402, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877306.

Hristache:2001:SAA

[HJPS01] Marian Hristache, Anatoli Juditsky, Jörg Polzehl, and Vladimir Spokoiny. Structure adaptive approach for dimension reduction. Annals of Statistics, 29(6):1537–1566, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015345954.

Hristache:2001:DEI

[HJS01] Marian Hristache, Anatoli Juditsky, and Vladimir Spokoiny. Direct estimation of the index coefficient in a single-index model. Annals of Statistics, 29(3):593–623, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1009210682.

Hall:2001:BND

[HK01] Peter Hall and Kee-Hoon Kang. Bootstrapping nonparametric density estimators with empirically chosen bandwidths. Annals of Statistics, 29(5):1443–1468, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013203461.

Hall:2005:BCN

[HK05] Peter Hall and Kee-Hoon Kang. Bandwidth choice for nonparametric classification. Annals of Statistics, 33(1):284–306, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1112967707.

Huang:2000:FAM

[HKST00] Jianhua Z. Huang, Charles Kooperberg, Charles J. Stone, and Young K. Truong. Functional ANOVA modeling for proportional hazards regression. Annals of Statistics, 28(4):961–999, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015956704.

Hoffman:2002:RRA

[HL02] M. Hoffman and O. Lepski. Random rates in anisotropic regression (with a discussion and a rejoinder by the authors). Annals of Statistics, 30(2):325–396, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1021379858.

Ho:2005:ISB

[HL05a] Yvonne H. S. Ho and Stephen M. S. Lee. Iterated smoothed bootstrap confidence intervals for population quantiles. Annals of Statistics, 33(1):437–462, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1112967712.

Hunter:2005:VSU

[HL05b] David R. Hunter and Runze Li. Variable selection using MM algorithms. Annals of Statistics, 33(4):1617–1642, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1123250224.

Hendriks:2007:ADA

[HL07] Harrie Hendriks and Zinoviy Landsman. Asymptotic data analysis on manifolds. Annals of Statistics, 35(1):109–131, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1181100183.

Hall:2008:EDM

[HL08] Peter Hall and Soumendra N. Lahiri. Estimation of distributions, moments and quantiles in deconvolution problems. Annals of Statistics, 36(5):2110–2134, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908086.

Hallin:2004:LLS

[HLT04] Marc Hallin, Zudi Lu, and Lanh T. Tran. Local linear spatial regression. Annals of Statistics, 32(6):2469–2500, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1107794876.

Hall:2003:SMD

[HM03] Peter Hall and Ilya Molchanov. Sequential methods for designadaptive estimation of discontinuities in regression curves and surfaces. Annals of Statistics, 31(3):921–941, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1056562467.

Horowitz:2004:NEA

[HM04] Joel L. Horowitz and Enno Mammen. Nonparametric estimation of an additive model with a link function. Annals of Statistics, 32(6):2412–2443, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794874.

Hall:2006:NEM

[HM06] Peter Hall and Tapabrata Maiti. Nonparametric estimation of mean-squared prediction error in nested-error regression models. Annals of Statistics, 34(4):1733–1750, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1162567631.

Hall:2007:TSP

[HM07a] Peter Hall and Yanyuan Ma. Testing the suitability of polynomial models in errors-in-variables problems. Annals of Statistics, 35(6):2620–2638, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012974.

Hall:2007:RPA

[HM07b] Peter Hall and Alexander Meister. A ridge-parameter approach to deconvolution. Annals of Statistics, 35(4):1535–1558, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1188405621.

Hillebrand:2007:ORC

[HM07c] Martin Hillebrand and Christine H. Müller. Outlier robust corner-preserving methods for reconstructing noisy images. Annals of Statistics, 35(1):132–165, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1181100184.

Horowitz:2007:ROE

[HM07d] Joel L. Horowitz and Enno Mammen. Rate-optimal estimation for a general class of nonparametric regression models with unknown link functions. Annals of Statistics, 35(6):2589–2619, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201012973.

Hobert:2008:TCD

[HM08] James P. Hobert and Dobrin Marchev. A theoretical comparison of the data augmentation, marginal augmentation and PX–DA algorithms. Annals of Statistics, 36(2):532–554, April 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1205420510.

Hall:2009:UBQ

[HM09] Peter Hall and Hugh Miller. Using the bootstrap to quantify the authority of an empirical ranking. Annals of Statistics, 37(6B): 3929–3959, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303532.

Hjort:2009:ESE

[HMV09] Nils Lid Hjort, Ian W. McKeague, and Ingrid Van Keilegom. Extending the scope of empirical likelihood. Annals of Statistics, 37(3):1079–1111, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369016.

Hall:2006:PPC

[HMW06] Peter Hall, Hans-Georg Müller, and Jane-Ling Wang. Properties of principal component methods for functional and longitudinal data analysis. Annals of Statistics, 34(3):1493–1517, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1152540756.

Hall:2009:EFD

[HMY09] Peter Hall, Hans-Georg Müller, and Fang Yao. Estimation of functional derivatives. Annals of Statistics, 37(6A):3307– 3329, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515388.

Hall:2004:BHN

[HMZ04] Peter Hall, Michael C. Minnotte, and Chunming Zhang. Bump hunting with non-Gaussian kernels. Annals of Statistics, 32(5): 2124–2141, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883784.

Huo:2007:WDS

[HN07] Xiaoming Huo and Xuelei (Sherry) Ni. When do stepwise algorithms meet subset selection criteria? Annals of Statistics, 35(2):870–887, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1183667297.

Hall:2004:APS

[HO04] Peter Hall and Hong Ooi. Attributing a probability to the shape of a probability density. Annals of Statistics, 32(5):2098–2123, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1098883783.

Ho:2006:BMM

[Ho06] Man-Wai Ho. A Bayes method for a monotone hazard rate via Spaths. Annals of Statistics, 34(2):820–836, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1151418242.

Hoff:2003:NEC

[Hof03] Peter D. Hoff. Nonparametric estimation of convex models via mixtures. Annals of Statistics, 31(1):174–200, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1046294461.

Hallin:2006:SERb

[HOP06] Marc Hallin, Hannu Oja, and Davy Paindaveine. Semiparametrically efficient rank-based inference for shape. II. optimal Restimation of shape. Annals of Statistics, 34(6):2757–2789, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1179935064.

Hossjer:2003:AET

[Hös03] Ola Hössjer. Asymptotic estimation theory of multipoint linkage analysis under perfect marker information. Annals of Statistics, 31(4):1075–1109, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1059655906.

Hall:2002:NMB

[HP02a] Peter Hall and Byeong U. Park. New methods for bias correction at endpoints and boundaries. Annals of Statistics, 30(5): 1460–1479, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1035844983.

Hallin:2002:OTM

[HP02b] Marc Hallin and Davy Paindaveine. Optimal tests for multivariate location based on interdirections and pseudo-Mahalanobis ranks. Annals of Statistics, 30(4):1103–1133, August 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031689019.

Hall:2004:WBE

[HP04a] Peter Hall and Spiridon Penev. Wavelet-based estimation with multiple sampling rates. Annals of Statistics, 32(5):1933–1956, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1098883777.

Hallin:2004:RBO

[HP04b] Marc Hallin and Davy Paindaveine. Rank-based optimal tests of the adequacy of an elliptic VARMA model. Annals of Statistics, 32(6):2642–2678, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794882.

Hallin:2006:SERa

[HP06] Marc Hallin and Davy Paindaveine. Semiparametrically efficient rank-based inference for shape. I. Optimal rank-based tests for sphericity. Annals of Statistics, 34(6):2707–2756, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1179935063.

Hallin:2008:ORB

[HP08] Marc Hallin and Davy Paindaveine. Optimal rank-based tests for homogeneity of scatter. Annals of Statistics, 36(3):1261–1298, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1211819564.

Hall:2008:CNO

[HPS08] Peter Hall, Byeong U. Park, and Richard J. Samworth. Choice of neighbor order in nearest-neighbor classification. Annals of Statistics, 36(5):2135–2152, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908087.

Hall:2002:EEC

[HPT02] Peter Hall, Liang Peng, and Nader Tajvidi. Effect of extrapolation on coverage accuracy of prediction intervals computed from Pareto-type data. Annals of Statistics, 30(3):875–895, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1028674844.

Hall:2007:NEP

[HQ07] Peter Hall and Peihua Qiu. Nonparametric estimation of a pointspread function in multivariate problems. Annals of Statistics, 35 (4):1512–1534, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1188405620.

Hall:2000:TSF

[HR00] Peter Hall and Christian Rau. Tracking a smooth fault line in a response surface. Annals of Statistics, 28(3):713–733, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1015951995.

Hoffmann:2008:NEL

[HR08] Marc Hoffmann and Markus Reiss. Nonlinear estimation for linear inverse problems with error in the operator. Annals of Statistics, 36(1):310–336, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201877303.

Hsing:2009:RFI

[HR09] Tailen Hsing and Haobo Ren. An RKHS formulation of the inverse regression dimension-reduction problem. Annals of Statistics, 37(2):726–755, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693148.

Hu:2007:CTL

[HRW07] Zhishui Hu, John Robinson, and Qiying Wang. Cramér-type large deviations for samples from a finite population. Annals of Statistics, 35(2):673–696, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1183667288.

Hobert:2002:CRT

[HS02] James P. Hobert and Jason Schweinsberg. Conditions for recurrence and transience of a Markov chain on Z^{+} and estimation of a geometric success probability. Annals of Statistics, 30(4): 1214–1223, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031689024.

He:2007:HOA

[HS07] Heping He and Thomas A. Severini. Higher-order asymptotic normality of approximations to the modified signed likelihood ratio statistic for regular models. Annals of Statistics, 35(5): 2054–2074, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1194461722.

Hofmann:2008:KMM

[HSS08] Thomas Hofmann, Bernhard Schölkopf, and Alexander J. Smola. Kernel methods in machine learning. Annals of Statistics, 36 (3):1171–1220, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1211819561.

Horvath:2001:EPS

[HT01] Lajos Horváth and Gilles Teyssière. Empirical process of the squared residuals of an arch sequence. Annals of Statistics, 29 (2):445–469, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210548.

Han:2004:GEC

[HT04] Dong Han and Fugee Tsung. A generalized EWMA control chart and its comparison with the optimal EWMA, CUSUM and GLR schemes. Annals of Statistics, 32(1):316–339, February 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1079120139.

Huang:2003:LAP

[Hua03] Jianhua Z. Huang. Local asymptotics for polynomial spline regression. Annals of Statistics, 31(5):1600–1635, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1065705120.

Huang:2004:CRP

[Hua04] Tzee-Ming Huang. Convergence rates for posterior distributions and adaptive estimation. Annals of Statistics, 32(4):1556–1593, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1091626179.

Huber:2002:JWT

[Hub02] Peter J. Huber. John w. Tukey’s contributions to robust statistics. Annals of Statistics, 30(6):1640–1648, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1043351251;   http://www.jstor.org/stable/1558734.

Hunter:2004:MAG

[Hun04] David R. Hunter. MM algorithms for generalized Bradley–Terry models. Annals of Statistics, 32(1):384–406, February 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1079120141.

Hall:2005:TMI

[HV05] Peter Hall and Ingrid Van Keilegom. Testing for monotone increasing hazard rate. Annals of Statistics, 33(3):1109–1137, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1120224097.

Hall:2006:AEE

[HV06] Peter Hall and Céline Vial. Assessing extrema of empirical principal component functions. Annals of Statistics, 34(3):1518–1544, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1152540757.

Hallin:2006:SNS

[HVW06] Marc Hallin, Catherine Vermandele, and Bas Werker. Serial and nonserial sign-and-rank statistics: Asymptotic representation and asymptotic normality. Annals of Statistics, 34(1): 254–289, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576263.

Hjort:2009:QPB

[HW09] Nils Lid Hjort and Stephen G. Walker. Quantile pyramids for Bayesian nonparametrics. Annals of Statistics, 37(1):105–131, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115929.

Hunter:2007:IMS

[HWH07] David R. Hunter, Shaoli Wang, and Thomas P. Hettmansperger. Inference for mixtures of symmetric distributions. Annals of Statistics, 35(1):224–251, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1181100187.

Hariz:2007:ORC

[HWZ07] Samir Ben Hariz, Jonathan J. Wylie, and Qiang Zhang. Optimal rate of convergence for nonparametric change-point estimators for nonstationary sequences. Annals of Statistics, 35(4): 1802–1826, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1188405631.

Hall:2003:ICV

[HY03a] Peter Hall and Qiwei Yao. Inference in components of variance models with low replication. Annals of Statistics, 31(2):414–441, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1051027875.

Hedayat:2003:UOB

[HY03b] A. S. Hedayat and Min Yang. Universal optimality of balanced uniform crossover designs. Annals of Statistics, 31(3):978–983, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562469.

Hall:2005:ACD

[HY05a] Peter Hall and Qiwei Yao. Approximating conditional distribution functions using dimension reduction. Annals of Statistics, 33(3):1404–1421, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1120224107.

Hedayat:2005:OEC

[HY05b] A. S. Hedayat and Min Yang. Optimal and efficient crossover designs for comparing test treatments with a control treatment. Annals of Statistics, 33(2):915–943, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114340.

Hall:2007:NEW

[HY07] Peter Hall and Adonis Yatchew. Nonparametric estimation when data on derivatives are available. Annals of Statistics, 35(1): 300–323, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1181100189.

Hall:2003:NEC

[HZ03] Peter Hall and Xiao-Hua Zhou. Nonparametric estimation of component distributions in a multivariate mixture. Annals of Statistics, 31(1):201–224, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1046294462.

Hu:2004:APD

[HZ04] Feifang Hu and Li-Xin Zhang. Asymptotic properties of doubly adaptive biased coin designs for multitreatment clinical trials. Annals of Statistics, 32(1):268–301, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1079120137.

Hu:2009:ERA

[HZH09] Feifang Hu, Li-Xin Zhang, and Xuming He. Efficient randomizedadaptive designs. Annals of Statistics, 37(5A):2543–2560, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663765.

Inglot:2003:MDM

[IK03] Tadeusz Inglot and Wilbert C. M. Kallenberg. Moderate deviations of minimum contrast estimators under contamination. Annals of Statistics, 31(3):852–879, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1056562465.

Inoue:2006:ERF

[IK06] Akihiko Inoue and Yukio Kasahara. Explicit representation of finite predictor coefficients and its applications. Annals of Statistics, 34(2):973–993, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418248.

Inglot:2000:CNV

[IKL00a] Tadeusz Inglot, Wilbert C. M. Kallenberg, and Teresa Ledwina. Correction note: “Vanishing shortcoming and asymptotic relative efficiency” (Annals of Statistics 28 (2000), 215–238). Annals of Statistics, 28(6):1795, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015957481. See [IKL00b].

Inglot:2000:VSA

[IKL00b] Tadeusz Inglot, Wilbert C. M. Kallenberg, and Teresa Ledwina. Vanishing shortcoming and asymptotic relative efficiency. Annals of Statistics, 28(1):215–238, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1016120370. See correction

[IKL00a].

Imhof:2001:MDE

[Imh01] Lorens A. Imhof. Maximin designs for exponential growth models and heteroscedastic polynomial models. Annals of Statistics, 29 (2):561–576, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210553.

Ing:2004:SOM

[Ing04] Ching-Kang Ing. Selecting optimal multistep predictors for autoregressive processes of unknown order. Annals of Statistics, 32(2):693–722, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178943.

Ing:2007:APE

[Ing07] Ching-Kang Ing. Accumulated prediction errors, information criteria and optimal forecasting for autoregressive time series. Annals of Statistics, 35(3):1238–1277, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1185304005.

Ishwaran:2005:SSV

[IR05] Hemant Ishwaran and J. Sunil Rao. Spike and slab variable selection: Frequentist and Bayesian strategies. Annals of Statistics, 33(2):730–773, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1117114335.

Imhof:2001:ODR

[IS01] Lorens Imhof and William J. Studden. E-optimal designs for rational models. Annals of Statistics, 29(3):763–783, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1009210689.

Ing:2005:OSS

[IW05] Ching-Kang Ing and Ching-Zong Wei. Order selection for samerealization predictions in autoregressive processes. Annals of Statistics, 33(5):2423–2474, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1132936568.

Jaeger:2005:ICD

[Jae05] Manfred Jaeger. Ignorability for categorical data. Annals of Statistics, 33(4):1964–1981, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250234.

James:2005:BPP

[Jam05a] Lancelot F. James. Bayesian Poisson process partition calculus with an application to Bayesian Lévy moving averages. Annals of Statistics, 33(4):1771–1799, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250229.

James:2005:FDP

[Jam05b] Lancelot F. James. Functionals of Dirichlet processes, the Cifarelli–Regazzini identity and beta–gamma processes. Annals of Statistics, 33(2):647–660, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114332.

James:2006:PCS

[Jam06] Lancelot F. James. Poisson calculus for spatial neutral to the right processes. Annals of Statistics, 34(1):416–440, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1146576269.

Janssen:2000:GPF

[Jan00] Arnold Janssen. Global power functions of goodness of fit tests. Annals of Statistics, 28(1):239–253, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1016120371.

Jensen:2010:SPA

[Jen10] Jens Ledet Jensen. On some problems in the article “Efficient Likelihood Estimation in State Space Models” by Cheng–Der Fuh [Ann. Statist. 34 (2006) 2026–2068]. Annals of Statistics, 38 (2):1279–1281, April 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1266586630. See [Fuh06].

Jermyn:2005:IBE

[Jer05] Ian H. Jermyn. Invariant Bayesian estimation on manifolds. Annals of Statistics, 33(2):583–605, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114330.

Jones:2004:SBG

[JH04] Galin L. Jones and James P. Hobert. Sufficient burn-in for Gibbs samplers for a hierarchical random effects model. Annals of Statistics, 32(2):784–817, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1083178947.

Jiang:2001:GFT

[Jia01] Jiming Jiang. Goodness-of-fit tests for mixed model diagnostics. Annals of Statistics, 29(4):1137–1164, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1013699997.

Jiang:2002:WBH

[Jia02] Wenxin Jiang. On weak base hypotheses and their implications for boosting regression and classification. Annals of Statistics, 30(1):51–73, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015362184.

Jiang:2004:PCA

[Jia04] Wenxin Jiang. Process consistency for AdaBoost. Annals of Statistics, 32(1):13–29, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1079120128.

Jiang:2005:POI

[Jia05] Jiming Jiang. Partially observed information and inference about non-Gaussian mixed linear models. Annals of Statistics, 33(6): 2695–2731, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1140191670.

Jiang:2007:BVS

[Jia07] Wenxin Jiang. Bayesian variable selection for high dimensional generalized linear models: Convergence rates of the fitted densities. Annals of Statistics, 35(4):1487–1511, August 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1188405619.

Juditsky:2009:NECa

[JLT09] Anatoli B. Juditsky, Oleg V. Lepski, and Alexandre B. Tsybakov. Nonparametric estimation of composite functions. Annals of Statistics, 37(3):1360–1404, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1239369025.

Jiang:2002:UJT

[JLW02] Jiming Jiang, P. Lahiri, and Shu-Mei Wan. A unified jackknife theory for empirical best prediction with M-estimation. Annals of Statistics, 30(6):1782–1810, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1043351257.

Jiang:2007:IEE

[JLW07] Jiming Jiang, Yihui Luan, and You-Gan Wang. Iterative estimating equations: Linear convergence and asymptotic properties. Annals of Statistics, 35(5):2233–2260, October 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1194461729.

Jensen:2004:EPC

[JM04] Soøren Tolver Jensen and Jesper Madsen. Estimation of proportional covariances in the presene of certain linear restrictions. Annals of Statistics, 32(1):219–232, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1079120134.

Jung:2009:PCH

[JM09] Sungkyu Jung and J. S. Marron. PCA consistency in high dimension, low sample size context. Annals of Statistics, 37(6B): 4104–4130, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303538.

James:2001:CEM

[JMP01] Lancelot F. James, David J. Marchette, and Carey E. Priebe. Consistent estimation of mixture complexity. Annals of Statistics, 29(5):1281–1296, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013203454.

Juditsky:2000:FAN

[JN00] Anatoli Juditsky and Arkadii Nemirovski. Functional aggregation for nonparametric regression. Annals of Statistics, 28 (3):681–712, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015951994.

Juditsky:2002:NTP

[JN02] Anatoli Juditsky and Arkadi Nemirovski. On nonparametric tests of positivity/monotonicity/convexity. Annals of Statistics, 30(2):498–527, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1021379863.

Juditsky:2009:NECb

[JN09] Anatoli B. Juditsky and Arkadi S. Nemirovski. Nonparametric estimation by convex programming. Annals of Statistics, 37(5A): 2278–2300, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247663755.

Johnstone:2001:DLE

[Joh01] Iain M. Johnstone. On the distribution of the largest eigenvalue in principal components analysis. Annals of Statistics, 29 (2):295–327, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210544.

Johnson:2004:BTG

[Joh04] Valen E. Johnson. A Bayesian χ ^{2} test for goodness-of-fit. Annals of Statistics, 32(6):2361–2384, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1107794872.

Johnstone:2008:MAJ

[Joh08] Iain M. Johnstone. Multivariate analysis and Jacobi ensembles: Largest eigenvalue, Tracy–Widom limits and rates of convergence. Annals of Statistics, 36(6):2638–2716, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1231165182.

Johannes:2009:DUE

[Joh09] Jan Johannes. Deconvolution with unknown error distribution. Annals of Statistics, 37(5A):2301–2323, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1247663756.

Jones:2000:LGA

[Jon00] L. K. Jones. Local greedy approximation for nonlinear regression and neural network training. Annals of Statistics, 28(5): 1379–1389, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957398.

Janssen:2003:HDB

[JP03] Arnold Janssen and Thorsten Pauls. How do bootstrap and permutation tests work? Annals of Statistics, 31(3):768–806, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562462.

Johnstone:2004:PBD

[JR04] Iain M. Johnstone and Marc Raimondo. Periodic boxcar deconvolution and Diophantine approximation. Annals of Statistics, 32 (5):1781–1804, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883772.

Jiang:2008:FMM

[JRGN08] Jiming Jiang, J. Sunil Rao, Zhonghua Gu, and Thuan Nguyen. Fence methods for mixed model selection. Annals of Statistics, 36 (4):1669–1692, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237296.

Juditsky:2008:LMA

[JRT08] A. Juditsky, P. Rigollet, and A. B. Tsybakov. Learning by mirror averaging. Annals of Statistics, 36(5):2183–2206, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1223908089.

Johnstone:2004:NSH

[JS04] Iain M. Johnstone and Bernard W. Silverman. Needles and straw in haystacks: Empirical Bayes estimates of possibly sparse sequences. Annals of Statistics, 32(4):1594–1649, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1091626180.

Johnstone:2005:EBS

[JS05] Iain M. Johnstone and Bernard W. Silverman. Empirical Bayes selection of wavelet thresholds. Annals of Statistics, 33(4): 1700–1752, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1123250227.

Jing:2004:SAS

[JSZ04] Bing-Yi Jing, Qi-Man Shao, and Wang Zhou. Saddlepoint approximation for Student’s t-statistic with no moment conditions. Annals of Statistics, 32(6):2679–2711, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1107794883.

Jiang:2008:GPV

[JT08] Wenxin Jiang and Martin A. Tanner. Gibbs posterior for variable selection in high-dimensional classification and data mining. Annals of Statistics, 36(5):2207–2231, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1223908090.

Jacod:2009:TCA

[JT09] Jean Jacod and Viktor Todorov. Testing for common arrivals of jumps for discretely observed multidimensional processes. Annals of Statistics, 37(4):1792–1838, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1245332833.

Jupp:2005:STG

[Jup05] P. E. Jupp. Sobolev tests of goodness of fit of distributions on compact Riemannian manifolds. Annals of Statistics, 33(6): 2957–2966, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1140191680.

Jupp:2008:DDS

[Jup08] P. E. Jupp. Data-driven Sobolev tests of uniformity on compact Riemannian manifolds. Annals of Statistics, 36(3):1246–1260, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1211819563.

Jongbloed:2009:ECD

[JvdM09] Geurt Jongbloed and Frank H. van der Meulen. Estimating a concave distribution function from data corrupted with additive noise. Annals of Statistics, 37(2):782–815, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1236693150.

Jing:2003:EEU

[JW03] Bing-Yi Jing and Qiying Wang. Edgeworth expansion for Ustatistics under minimal conditions. Annals of Statistics, 31(4): 1376–1391, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1059655916.

Jager:2007:GFT

[JW07] Leah Jager and Jon A. Wellner. Goodness-of-fit tests via phidivergences. Annals of Statistics, 35(5):2018–2053, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1194461721.

James:2009:FLR

[JWZ09] Gareth M. James, Jing Wang, and Ji Zhu. Functional linear regression that’s interpretable. Annals of Statistics, 37(5A): 2083–2108, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247663749.

Jiang:2009:GML

[JZ09] Wenhua Jiang and Cun-Hui Zhang. General maximum likelihood empirical Bayes estimation of normal means. Annals of Statistics, 37(4):1647–1684, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1245332828.

Kargin:2005:CBE

[Kar05] Vladislav Kargin. On the Chernoff bound for efficiency of quantum hypothesis testing. Annals of Statistics, 33(2):959–976, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1117114342.

Karoui:2008:ONC

[Kar08a] Noureddine El Karoui. Operator norm consistent estimation of large-dimensional sparse covariance matrices. Annals of Statistics, 36(6):2717–2756, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1231165183.

Karoui:2008:SEL

[Kar08b] Noureddine El Karoui. Spectrum estimation for large dimensional covariance matrices using random matrix theory. Annals of Statistics, 36(6):2757–2790, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1231165184.

Kim:2004:FSP

[Kim04] Jae Kwang Kim. Finite sample properties of multiple imputation estimators. Annals of Statistics, 32(2):766–783, April 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1083178946.

Kim:2006:BMT

[Kim06] Yongdai Kim. The Bernstein–von Mises theorem for the proportional hazard model. Annals of Statistics, 34(4):1678–1700, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1162567629.

Kim:2007:QRV

[Kim07] Mi-Ok Kim. Quantile regression with varying coefficients. Annals of Statistics, 35(1):92–108, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1181100182.

Khmaladze:2004:MTG

[KK04] Estate V. Khmaladze and Hira L. Koul. Martingale transforms goodness-of-fit tests in regression models. Annals of Statistics, 32(3):995–1034, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408493.

Khmaladze:2009:GFP

[KK09] Estate V. Khmaladze and Hira L. Koul. Goodness-of-fit problem for errors in nonparametric regression: Distribution free approach. Annals of Statistics, 37(6A):3165–3185, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1250515383.

Kim:2001:PCS

[KL01] Yongdai Kim and Jaeyong Lee. On posterior consistency of survival models. Annals of Statistics, 29(3):666–686, June 2001. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1009210685.

Kim:2003:BAP

[KL03a] Yongdai Kim and Jaeyong Lee. Bayesian analysis of proportional hazard models. Annals of Statistics, 31(2):493–511, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1051027878.

Kim:2003:BBP

[KL03b] Yongdai Kim and Jaeyong Lee. Bayesian bootstrap for proportional hazards models. Annals of Statistics, 31(6):1905–1922, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1074290331.

Kim:2004:BMT

[KL04] Yongdai Kim and Jaeyong Lee. A Bernstein–von Mises theorem in the nonparametric right-censoring model. Annals of Statistics, 32(4):1492–1512, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1091626176.

Kulikov:2005:ANE

[KL05] Vladimir N. Kulikov and Hendrik P. Lopuhaä. Asymptotic normality of the L_{k}-error of the Grenander estimator. Annals of Statistics, 33(5):2228–2255, October 2005. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1132936562.

Koul:2006:FED

[KL06a] Hira L. Koul and Shiqing Ling. Fitting an error distribution in some heteroscedastic time series models. Annals of Statistics, 34 (2):994–1012, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1151418249.

Kulikov:2006:BND

[KL06b] Vladimir N. Kulikov and Hendrik P. Lopuhaä. The behavior of the NPMLE of a decreasing density near the boundaries of the support. Annals of Statistics, 34(2):742–768, April 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1151418239.

Kosorok:2004:RIU

[KLF04] Michael R. Kosorok, Bee Leng Lee, and Jason P. Fine. Robust inference for univariate proportional hazards frailty regression models. Annals of Statistics, 32(4):1448–1491, August 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1091626175.

Kunert:2000:DOD

[KM00] J. Kunert and R. J. Martin. On the determination of optimal designs for an interference model. Annals of Statistics, 28(6): 1728–1742, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957478.

Kalifa:2003:TEL

[KM03] Jérôme Kalifa and Stéphane Mallat. Thresholding estimators for linear inverse problems and deconvolutions. Annals of Statistics, 31(1):58–109, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1046294458.

Kosorok:2007:MAL

[KM07] Michael R. Kosorok and Shuangge Ma. Marginal asymptotics for the “large p, small n” paradigm: With applications to microarray data. Annals of Statistics, 35(4):1456–1486, August 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1188405618.

Koshevoy:2003:SME

[KMO03] Gleb A. Koshevoy, Jyrki Möttönen, and Hannu Oja. A scatter matrix estimate based on the zonotope. Annals of Statistics, 31 (5):1439–1459, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1065705114.

Kochar:2000:EMM

[KMS00] Subhash C. Kochar, Hari Mukerjee, and Francisco J. Samaniego. Estimation of a monotone mean residual life. Annals of Statistics, 28(3):905–921, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015952004.

Karlsen:2007:NEN

[KMT07] Hans Arnfinn Karlsen, Terje Myklebust, and Dag Tjøstheim. Nonparametric estimation in a nonlinear cointegration type model. Annals of Statistics, 35(1):252–299, February 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1181100188.

Kolaczyk:2004:MLA

[KN04] Eric D. Kolaczyk and Robert D. Nowak. Multiscale likelihood analysis and complexity penalized estimation. Annals of Statistics, 32(2):500–527, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1083178936.

Koltchinskii:2000:EGM

[Kol00] V. I. Koltchinskii. Empirical geometry of multivariate data: a deconvolution approach. Annals of Statistics, 28(2):591–629, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1016218232.

Kolassa:2003:MST

[Kol03] John E. Kolassa. Multivariate saddlepoint tail probability approximations. Annals of Statistics, 31(1):274–286, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1046294465.

Koltchinskii:2006:LRC

[Kol06a] Vladimir Koltchinskii. Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6): 2593–2656, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1179935055.

Koltchinskii:2006:RLR

[Kol06b] Vladimir Koltchinskii. Rejoinder: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2697–2706, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935062.

Koltchinskii:2009:SRC

[Kol09] Vladimir Koltchinskii. Sparse recovery in convex hulls via entropy penalization. Annals of Statistics, 37(3):1332–1359, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1239369024.

Komaki:2004:SPI

[Kom04] Fumiyasu Komaki. Simultaneous prediction of independent Poisson observables. Annals of Statistics, 32(4):1744–1769, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1091626186.

Komaki:2006:SPB

[Kom06] Fumiyasu Komaki. Shrinkage priors for Bayesian prediction. Annals of Statistics, 34(2):808–819, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418241.

Kou:2004:FSA

[Kou04] S. C. Kou. From finite sample to asymptotics: A geometric bridge for selection criteria in spline regression. Annals of Statistics, 32(6):2444–2468, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794875.

Kozek:2003:MEN

[Koz03] Andrzej S. Kozek. On M-estimators and normal quantiles. Annals of Statistics, 31(4):1170–1185, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1059655910.

Koltchinskii:2002:EMD

[KP02] V. Koltchinskii and D. Panchenko. Empirical margin distributions and bounding the generalization error of combined classifiers. Annals of Statistics, 30(1):1–50, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015362183.

Kreiss:2003:AAP

[KP03] Jens-Peter Kreiss and Efstathios Paparoditis. Autoregressiveaided periodogram bootstrap for timeseries. Annals of Statistics, 31(6):1923–1955, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290332.

Klaassen:2005:EEB

[KP05a] Chris A. J. Klaassen and Hein Putter. Efficient estimation of Banach parameters in semiparametric models. Annals of Statistics, 33(1):307–346, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1112967708.

Koltchinskii:2005:CCC

[KP05b] Vladimir Koltchinskii and Dmitry Panchenko. Complexities of convex combinations and bounding the generalization error in classification. Annals of Statistics, 33(4):1455–1496, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1123250220.

Kulik:2009:WRR

[KR09] Rafał Kulik and Marc Raimondo. Wavelet regression in random design with heteroscedastic dependent errors. Annals of Statistics, 37(6A):3396–3430, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1250515391.

Kosorok:2007:IUR

[KS07] Michael R. Kosorok and Rui Song. Inference under right censoring for transformation models with a change-point based on a covariate threshold. Annals of Statistics, 35(3):957–989, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1185303994.

Koul:2009:MDR

[KS09] Hira L. Koul and Weixing Song. Minimum distance regression model checking with Berkson measurement errors. Annals of Statistics, 37(1):132–156, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115930.

Koltchinskii:2007:ICN

[KSC07] Vladimir Koltchinskii, Lyudmila Sakhanenko, and Songhe Cai. Integral curves of noisy vector fields and statistical problems in diffusion tensor imaging: Nonparametric kernel estimation and hypotheses testing. Annals of Statistics, 35(4):1576–1607, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1188405623.

Karlsen:2001:NEN

[KT01a] Hans Arnfinn Karlsen and Dag Tjøstheim. Nonparametric estimation in null recurrent time series. Annals of Statistics, 29 (2):372–416, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210546.

Kent:2001:RUC

[KT01b] John T. Kent and David E. Tyler. Regularity and uniqueness for constrained M-estimates and redescending M-estimates. Annals of Statistics, 29(1):252–265, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/996986508.

Klemela:2001:SAE

[KT01c] Jussi Klemelä and Alexandre B. Tsybakov. Sharp adaptive estimation of linear functionals. Annals of Statistics, 29(6):1567– 1600, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015345955.

Kuriki:2001:TPM

[KT01d] Satoshi Kuriki and Akimichi Takemura. Tail probabilities of the maxima of multilinear forms and their applications. Annals of Statistics, 29(2):328–371, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1009210545.

Kunsch:2005:RMC

[Kün05] Hans R. Künsch. Recursive Monte Carlo filters: Algorithms and theoretical analysis. Annals of Statistics, 33(5):1983–2021, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936554.

Kleijn:2006:MID

[KvdV06] B. J. K. Kleijn and A. W. van der Vaart. Misspecification in infinite-dimensional Bayesian statistics. Annals of Statistics, 34 (2):837–877, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1151418243.

Koltchinskii:2004:TPB

[KY04] Vladimir Koltchinskii and Bin Yu. Three papers on boosting: an introduction. Annals of Statistics, 32(1):12, February 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1079120127.

Kulperger:2005:HMP

[KY05] Reg Kulperger and Hao Yu. High moment partial sum processes of residuals in GARCH models and their applications. Annals of Statistics, 33(5):2395–2422, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1132936567.

Kou:2006:R

[KZW06a] S. C. Kou, Qing Zhou, and Wing H. Wong. Rejoinder. Annals of Statistics, 34(4):1646–1652, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1162567627.

Kou:2006:EES

[KZW06b] S. C. Kou, Qing Zhou, and Wing Hung Wong. Equi-energy sampler with applications in statistical inference and statistical mechanics. Annals of Statistics, 34(4):1581–1619, August 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1162567622.

Lahiri:2003:NSC

[Lah03] S. N. Lahiri. A necessary and sufficient condition for asymptotic independence of discrete Fourier transforms under shortand long-range dependence. Annals of Statistics, 31(2):613–641, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1051027883.

Lahiri:2005:CJA

[Lah05] S. N. Lahiri. Consistency of the jackknife-after-bootstrap variance estimator for the bootstrap quantiles of a studentized statistic. Annals of Statistics, 33(5):2475–2506, October 2005. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1132936569.

Lahiri:2007:AES

[Lah07] S. N. Lahiri. Asymptotic expansions for sums of block-variables under weak dependence. Annals of Statistics, 35(3):1324–1350, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1185458986.

Lai:2003:SAI

[Lai03] Tze Leung Lai. Stochastic approximation: invited paper. Annals of Statistics, 31(2):391–406, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1051027873.

Lang:2004:MPH

[Lan04] Joseph B. Lang. Multinomial-Poisson homogeneous models for contingency tables. Annals of Statistics, 32(1):340–383, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1079120140.

Lin:2004:SPM

[LB04] Yi Lin and Lawrence D. Brown. Statistical properties of the method of regularization with periodic Gaussian reproducing kernel. Annals of Statistics, 32(4):1723–1743, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1091626185.

Levina:2006:TSN

[LB06] Elizaveta Levina and Peter J. Bickel. Texture synthesis and nonparametric resampling of random fields. Annals of Statistics, 34 (4):1751–1773, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567632.

Lan:2009:CPE

[LBM09] Yan Lan, Moulinath Banerjee, and George Michailidis. Changepoint estimation under adaptive sampling. Annals of Statistics, 37(4):1752–1791, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1245332832.

Lu:2004:MRL

[LC04] Guobing Lu and John B. Copas. Missing at random, likelihood ignorability and model completeness. Annals of Statistics, 32 (2):754–765, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178945.

Lin:2005:ABS

[LC05] Pei-Sheng Lin and Murray K. Clayton. Analysis of binary spatial data by quasi-likelihood estimating equations. Annals of Statistics, 33(2):542–555, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114328.

Li:2003:DRC

[LCC03] Bing Li, R. Dennis Cook, and Francesca Chiaromonte. Dimension reduction for the conditional mean in regressions with categorical predictors. Annals of Statistics, 31(5):1636–1668, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1065705121.

Li:2009:DRN

[LD09] Bing Li and Yuexiao Dong. Dimension reduction for nonelliptically distributed predictors. Annals of Statistics, 37(3): 1272–1298, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369022.

Lecue:2007:SAM

[Lec07] Guillaume Lecué. Simultaneous adaptation to the margin and to complexity in classification. Annals of Statistics, 35(4): 1698–1721, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1188405627.

Leeb:2009:CPI

[Lee09] Hannes Leeb. Conditional predictive inference post model selection. Annals of Statistics, 37(5B):2838–2876, October 2009. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1247836671.

Leung:2005:CVN

[Leu05] Denis Heng-Yan Leung. Cross-validation in nonparametric regression with outliers. Annals of Statistics, 33(5):2291–2310, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936564.

Lam:2008:PKL

[LF08] Clifford Lam and Jianqing Fan. Profile-kernel likelihood inference with diverging number of parameters. Annals of Statistics, 36(5): 2232–2260, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1223908091.

Lam:2009:SRC

[LF09a] Clifford Lam and Jianqing Fan. Sparsistency and rates of convergence in large covariance matrix estimation. Annals of Statistics, 37(6B):4254–4278, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1256303543.

Lv:2009:UAM

[LF09b] Jinchi Lv and Yingying Fan. A unified approach to model selection and sparse recovery using regularized least squares. Annals of Statistics, 37(6A):3498–3528, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515394.

Li:2002:MTG

[Li02] Heng Li. Modeling through group invariance: an interesting example with potential applications. Annals of Statistics, 30(4): 1069–1080, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031689017.

Liang:2000:EBT

[Lia00] TaChen Liang. On an empirical Bayes test for a normal mean. Annals of Statistics, 28(2):648–655, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1016218234.

Liang:2009:ISU

[Lia09] Faming Liang. Improving SAMC using smoothing methods: Theory and applications to Bayesian model selection problems. Annals of Statistics, 37(5B):2626–2654, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1247836663.

Lin:2000:TPS

[Lin00] Yi Lin. Tensor product space ANOVA models. Annals of Statistics, 28(3):734–755, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015951996.

Ling:2007:TCP

[Lin07] Shiqing Ling. Testing for change points in time series models and limiting theorems for NED sequences. Annals of Statistics, 35(3):1213–1237, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1185304004.

Lam:2000:ESC

[LL00] Tao-Kai Lam and Wei-Liem Loh. Estimating structured correlation matrices in smooth Gaussian random field models. Annals of Statistics, 28(3):880–904, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015952003.

Li:2008:MSB

[LL08a] Jun Li and Regina Y. Liu. Multivariate spacings based on data depth: I. Construction of nonparametric multivariate tolerance regions. Annals of Statistics, 36(3):1299–1323, June 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1211819565.

Li:2008:VSS

[LL08b] Runze Li and Hua Liang. Variable selection in semiparametric regression modeling. Annals of Statistics, 36(1):261–286, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877301.

Lin:2009:MPS

[LL09a] Zhengyan Lin and Weidong Liu. On maxima of periodograms of stationary processes. Annals of Statistics, 37(5B):2676–2695, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247836665.

Liu:2009:BUS

[LL09b] Jiawei Liu and Bruce G. Lindsay. Building and using semiparametric tolerance regions for parametric multinomial models. Annals of Statistics, 37(6A):3644–3659, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515400.

Laurent:2000:AEQ

[LM00] B. Laurent and P. Massart. Adaptive estimation of a quadratic functional by model selection. Annals of Statistics, 28(5): 1302–1338, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957395.

Ling:2003:AEN

[LM03] Shiqing Ling and Michael McAleer. On adaptive estimation in nonstationary ARMA models with GARCH errors. Annals of Statistics, 31(2):642–674, April 2003. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1051027884.

Lahiri:2007:DCP

[LM07a] P. Lahiri and Kanchan Mukherjee. On the design-consistency property of hierarchical Bayes estimators in finite population sampling. Annals of Statistics, 35(2):724–737, April 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1183667290.

Letac:2007:WDD

[LM07b] Gérard Letac and Hélène Massam. Wishart distributions for decomposable graphs. Annals of Statistics, 35(3):1278–1323, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1185304006.

Levitz:2001:SCP

[LMP01] Michael Levitz, David Madigan, and Michael D. Perlman. Separation and completeness properties for AMP chain graph Markov models. Annals of Statistics, 29(6):1751–1784, December 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1015345961. See correction [LPM03].

Lindsay:2008:QDP

[LMR^{+}08] Bruce G. Lindsay, Marianthi Markatou, Surajit Ray, Ke Yang, and Shu-Chuan Chen. Quadratic distances on probabilities: A unified foundation. Annals of Statistics, 36(2):983–1006, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1205420526.

Linton:2003:EMA

[LNV03] Oliver B. Linton, Jens Perch Nielsen, and Sara Van de Geer. Estimating multiplicative and additive hazard functions by kernel methods. Annals of Statistics, 31(2):464–492, April 2003. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1051027877.

Loh:2003:ADS

[Loh03] Wei-Liem Loh. On the asymptotic distribution of scrambled net quadrature. Annals of Statistics, 31(4):1282–1324, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1059655914.

Loh:2005:FDA

[Loh05] Wei-Liem Loh. Fixed-domain asymptotics for a subclass of Matérn-type Gaussian random fields. Annals of Statistics, 33(5): 2344–2394, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1132936566.

Loh:2008:MCL

[Loh08] Wei-Liem Loh. A multivariate central limit theorem for randomized orthogonal array sampling designs in computer experiments. Annals of Statistics, 36(4):1983–2023, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1216237306.

Lok:2008:SMC

[Lok08] Judith J. Lok. Statistical modeling of causal effects in continuous time. Annals of Statistics, 36(3):1464–1507, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1211819571.

Lorden:2005:NEA

[LP05] Gary Lorden and Moshe Pollak. Nonanticipating estimation applied to sequential analysis and changepoint detection. Annals of Statistics, 33(3):1422–1454, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1120224108.

Leeb:2006:COE

[LP06] Hannes Leeb and Benedikt M. Pötscher. Can one estimate the conditional distribution of post-model-selection estimators? Annals of Statistics, 34(5):2554–2591, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1169571807.

Leonenko:2010:CCR

[LP10] Nikolai Leonenko and Luc Pronzato. Correction: “A class of Rényi information estimators for multidimensional densities”. Annals of Statistics, 38(6):3837–3838, December 2010. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1291126975.  See

[LPS08].

Levitz:2003:CSC

[LPM03] Michael Levitz, Michael D. Perlman, and David Madigan. Correction: “Separation and completeness properties for AMP chain graph Markov models”. Annals of Statistics, 31(1):348, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1046294468. See [LMP01].

Leonenko:2008:CRI

[LPS08] Nikolai Leonenko, Luc Pronzato, and Vippal Savani. A class of Rényi information estimators for multidimensional densities. Annals of Statistics, 36(5):2153–2182, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1223908088. See correction [LP10].

Loh:2000:EFP

[LQS00] Ji Meng Loh, Jean M. Quashnock, and Michael L. Stein. Estimating the K function of a point process with an application to cosmology. Annals of Statistics, 28(6):1503–1532, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957468.

Lii:2002:SAH

[LR02] Keh-Shin Lii and Murray Rosenblatt. Spectral analysis for harmonizable processes. Annals of Statistics, 30(1):258–297, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1015362193. See correction [LR03].

Lii:2003:CSA

[LR03] Keh-Shin Lii and Murray Rosenblatt. Correction: “Spectral analysis for harmonizable processes”. Annals of Statistics, 31 (5):1693, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1065705123. See [LR02].

Lehmann:2005:GFE

[LR05] E. L. Lehmann and Joseph P. Romano. Generalizations of the familywise error rate. Annals of Statistics, 33(3):1138–1154, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1120224098.

Lii:2006:EAP

[LR06] Keh-Shin Lii and Murray Rosenblatt. Estimation for almost periodic processes. Annals of Statistics, 34(3):1115–1139, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1152540744.

Lii:2008:EAP

[LR08] Keh-Shin Lii and Murray Rosenblatt. Estimation for almost periodic processes. Annals of Statistics, 36(3):1508, June 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1211819572.

Lehmann:2005:OSS

[LRS05] E. L. Lehmann, Joseph P. Romano, and Juliet Popper Shaffer. On optimality of stepdown and stepup multiple test procedures. Annals of Statistics, 33(3):1084–1108, June 2005. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1120224096.

Lieberman:2003:VAE

[LRZ03] Offer Lieberman, Judith Rousseau, and David M. Zucker. Valid asymptotic expansions for the maximum likelihood estimator of the parameter of a stationary, Gaussian, strongly dependent process. Annals of Statistics, 31(2):586–612, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1051027882.

Li:2000:PDP

[LS00] Lei Li and Terence P. Speed. Parametric deconvolution of positive spike trains. Annals of Statistics, 28(5):1279–1301, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957394.

Liu:2003:ALR

[LS03] Xin Liu and Yongzhao Shao. Asymptotics for likelihood ratio tests under loss of identifiability. Annals of Statistics, 31(3):807– 832, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562463.

Linton:2008:EST

[LSV08] Oliver Linton, Stefan Sperlich, and Ingrid Van Keilegom. Estimation of a semiparametric transformation model. Annals of Statistics, 36(2):686–718, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1205420516.

Ling:2005:TLM

[LT05] Shiqing Ling and Howell Tong. Testing for a linear MA model against threshold MA models. Annals of Statistics, 33(6):2529– 2552, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1140191665.

Luati:2004:MFI

[Lua04] Alessandra Luati. Maximum Fisher information in mixed state quantum systems. Annals of Statistics, 32(4):1770–1779, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1091626187.

Lugosi:2004:BRC

[LV04] Gábor Lugosi and Nicolas Vayatis. On the Bayes-risk consistency of regularized boosting methods. Annals of Statistics, 32 (1):30–55, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1079120129.

Ledoit:2002:SHT

[LW02] Olivier Ledoit and Michael Wolf. Some hypothesis tests for the covariance matrix when the dimension is large compared to the sample size. Annals of Statistics, 30(4):1081–1102, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1031689018.

Lugosi:2004:CRL

[LW04] Gábor Lugosi and Marten Wegkamp. Complexity regularization via localized random penalties. Annals of Statistics, 32(4): 1679–1697, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1091626183.

Lafferty:2008:RSG

[LW08] John Lafferty and Larry Wasserman. Rodeo: Sparse, greedy nonparametric regression. Annals of Statistics, 36(1):28–63, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877293.

Li:2007:NEC

[LWH^{+}07] Yehua Li, Naisyin Wang, Meeyoung Hong, Nancy D. Turner, Joanne R. Lupton, and Raymond J. Carroll. Nonparametric estimation of correlation functions in longitudinal and spatial data, with application to colon carcinogenesis experiments. Annals of Statistics, 35(4):1608–1643, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405624.

Luo:2009:CPD

[LWT09] Ronghua Luo, Hansheng Wang, and Chih-Ling Tsai. Contour projected dimension reduction. Annals of Statistics, 37(6B): 3743–3778, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303526.

Li:2007:SDR

[LY07] Bing Li and Xiangrong Yin. On surrogate dimension reduction for measurement error regression: An invariance law. Annals of Statistics, 35(5):2143–2172, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1194461725.

Lahiri:2006:RMS

[LZ06a] S. N. Lahiri and Jun Zhu. Resampling methods for spatial regression models under a class of stochastic designs. Annals of Statistics, 34(4):1774–1813, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1162567633.

Lin:2006:CSS

[LZ06b] Yi Lin and Hao Helen Zhang. Component selection and smoothing in multivariate nonparametric regression. Annals of Statistics, 34(5):2272–2297, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571797.

Li:2007:ASA

[LZ07a] Yingxing Li and Li-Xing Zhu. Asymptotics for sliced average variance estimation. Annals of Statistics, 35(1):41–69, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1181100180.

Low:2007:CCT

[LZ07b] Mark G. Low and Harrison H. Zhou. A complement to le Cam’s theorem. Annals of Statistics, 35(3):1146–1165, July 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1185304001.

Li:2005:CRG

[LZC05] Bing Li, Hongyuan Zha, and Francesca Chiaromonte. Contour regression: A general approach to dimension reduction. Annals of Statistics, 33(4):1580–1616, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250223.

Madsen:2000:INM

[Mad00] Jesper Madsen. Invariant normal models with recursive graphical Markov structure. Annals of Statistics, 28(4):1150–1178, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015956711.

Marcheselli:2003:ARJ

[Mar03] Marzia Marcheselli. Asymptotic results in jackknifing nonsmooth functions of the sample mean vector. Annals of Statistics, 31(6): 1885–1904, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1074290330.

Marinucci:2006:HRA

[Mar06] Domenico Marinucci. High-resolution asymptotics for the angular bispectrum of spherical random fields. Annals of Statistics, 34(1):1–41, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576254.

Meinshausen:2006:HDG

[MB06] Nicolai Meinshausen and Peter Bühlmann. High-dimensional graphs and variable selection with the lasso. Annals of Statistics, 34(3):1436–1462, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540754.

McCullagh:2002:WSM

[McC02] Peter McCullagh. What is a statistical model? Annals of Statistics, 30(5):1225–1310, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1035844977.

Mei:2006:CNO

[Mei06a] Yajun Mei. Comments on “A note on optimal detection of a change in distribution,” by Benjamin Yakir. Annals of Statistics, 34(3):1570–1576, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540759.

Mei:2006:SCP

[Mei06b] Yajun Mei. Sequential change-point detection when unknown parameters are present in the pre-change distribution. Annals of Statistics, 34(1):92–122, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1146576257.

Meng:2008:DOS

[Men08] Xiao-Li Meng. Discussion: One-step sparse estimates in nonconcave penalized likelihood models: Who cares if it is a white cat or a black cat? Annals of Statistics, 36(4):1542–1552, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1216237289.

Merkouris:2007:TME

[Mer07] T. Merkouris. Transform martingale estimating functions. Annals of Statistics, 35(5):1975–2000, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461719.

May:2009:ARA

[MF09] Caterina May and Nancy Flournoy. Asymptotics in responseadaptive designs generated by a two-color, randomly reinforced urn. Annals of Statistics, 37(2):1058–1078, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1236693160.

Mizera:2002:DDP

[Miz02] Ivan Mizera. On depth and deep points: a calculus. Annals of Statistics, 30(6):1681–1736, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1043351254.

Ma:2005:PLL

[MK05] Shuangge Ma and Michael R. Kosorok. Penalized log-likelihood estimation for partly linear transformation models with current status data. Annals of Statistics, 33(5):2256–2290, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936563.

Maathuis:2009:EHD

[MKB09] Marloes H. Maathuis, Markus Kalisch, and Peter Bühlmann. Estimating high-dimensional intervention effects from observational data. Annals of Statistics, 37(6A):3133–3164, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1250515382.

Minary:2006:DEE

[ML06] Peter Minary and Michael Levitt. Discussion of “Equi-energy sampler” by Kou, Zhou and Wong. Annals of Statistics, 34(4): 1636–1641, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567625.

Mao:2007:ENC

[ML07] Chang Xuan Mao and Bruce G. Lindsay. Estimating the number of classes. Annals of Statistics, 35(2):917–930, April 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1183667299.

Massam:2009:CPD

[MLD09] Hélène Massam, Jinnan Liu, and Adrian Dobra. A conjugate prior for discrete hierarchical log-linear models. Annals of Statistics, 37(6A):3431–3467, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1250515392.

Merkl:2007:OTR

[MM07] Franz Merkl and Leila Mohammadi. Optimal third root asymptotic bounds in the statistical estimation of thresholds. Annals of Statistics, 35(5):2193–2218, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1194461727.

Malliavin:2009:FTM

[MM09] Paul Malliavin and Maria Elvira Mancino. A Fourier transform method for nonparametric estimation of multivariate volatility. Annals of Statistics, 37(4):1983–2010, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1245332838.

Massart:2006:RBS

[MN06] Pascal Massart and Élodie Nédélec. Risk bounds for statistical learning. Annals of Statistics, 34(5):2326–2366, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1169571799.

McNeil:2009:MAC

[MN09] Alexander J. McNeil and Johanna Nešlehová. Multivariate Archimedean copulas, d-monotone functions and ℓ_{1}-norm symmetric distributions. Annals of Statistics, 37(5B):3059–3097, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247836677.

Moreira:2009:MLM

[Mor09] Marcelo J. Moreira. A maximum likelihood method for the incidental parameter problem. Annals of Statistics, 37(6A):3660– 3696, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515401.

Moustakides:2004:OCP

[Mou04] George V. Moustakides. Optimality of the CUSUM procedure in continuous time. Annals of Statistics, 32(1):302–315, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1079120138.

Moustakides:2008:SCD

[Mou08] George V. Moustakides. Sequential change detection revisited. Annals of Statistics, 36(2):787–807, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1205420519.

Marchand:2001:IMB

[MP01] Éric Marchand and François Perron. Improving on the MLE of a bounded normal mean. Annals of Statistics, 29(4):1078–1093, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013699994.

Marinucci:2004:EPG

[MP04] Domenico Marinucci and Mauro Piccioni. The empirical process on Gaussian spherical harmonics. Annals of Statistics, 32 (3):1261–1288, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408502.

Mammen:2005:BSS

[MP05a] Enno Mammen and Byeong U. Park. Bandwidth selection for smooth backfitting in additive models. Annals of Statistics, 33 (3):1260–1294, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1120224102.

Mardia:2005:DPS

[MP05b] Kanti V. Mardia and Vic Patrangenaru. Directions and projective shapes. Annals of Statistics, 33(4):1666–1699, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1123250226.

Mammen:2006:SSB

[MP06] Enno Mammen and Byeong U. Park. A simple smooth backfitting method for additive models. Annals of Statistics, 34(5): 2252–2271, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1169571796.

McElroy:2007:CIR

[MP07a] Tucker McElroy and Dimitris N. Politis. Computer-intensive rate estimation, diverging statistics and scanning. Annals of Statistics, 35(4):1827–1848, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1188405632.

McElroy:2007:SMP

[MP07b] Tucker McElroy and Dimitris N. Politis. Stable marked point processes. Annals of Statistics, 35(1):393–419, February 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1181100192.

Mokkadem:2007:CKW

[MP07c] Abdelkader Mokkadem and Mariane Pelletier. A companion for the Kiefer–Wolfowitz–Blum stochastic approximation algorithm. Annals of Statistics, 35(4):1749–1772, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1188405629.

Moulines:2005:RET

[MPR05] Eric Moulines, Pierre Priouret, and François Roueff. On recursive estimation for time varying autoregressive processes. Annals of Statistics, 33(6):2610–2654, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191668.

Muler:2009:REA

[MPY09] Nora Muler, Daniel Peña, and Víctor J. Yohai. Robust estimation for ARMA models. Annals of Statistics, 37(2):816–840, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1236693151.

Marinucci:2001:NBA

[MR01] D. Marinucci and P. M. Robinson. Narrow-band analysis of nonstationary processes. Annals of Statistics, 29(4):947–986, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013699988.

Meinshausen:2006:EPF

[MR06] Nicolai Meinshausen and John Rice. Estimating the proportion of false null hypotheses among a large number of independently tested hypotheses. Annals of Statistics, 34(1):373–393, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1146576267.

Morgan:2007:RDL

[MR07] J. P. Morgan and Brian H. Reck. Resolvable designs with large blocks. Annals of Statistics, 35(2):747–771, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1183667292.

Moulines:2008:WWE

[MRT08] E. Moulines, F. Roueff, and M. S. Taqqu. A wavelet Whittle estimator of the memory parameter of a nonstationary Gaussian time series. Annals of Statistics, 36(4):1925–1956, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1216237304.

Meinshausen:2007:DTT

[MRY07] N. Meinshausen, G. Rocha, and B. Yu. Discussion: A tale of three cousins: Lasso, L2Boosting and Dantzig. Annals of Statistics, 35(6):2373–2384, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012963.

Martinussen:2000:NDA

[MS00a] Torben Martinussen and Thomas H. Scheike. A nonparametric dynamic additive regression model for longitudinal data. Annals of Statistics, 28(4):1000–1025, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015956705.

Mikosch:2000:LTS

[MS00b] Thomas Mikosch and Cătălin Stărică. Limit theory for the sample autocorrelations and extremes of a GARCH (1,1) process. Annals of Statistics, 28(5):1427–1451, October 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015957401.

Mercurio:2004:SIT

[MS04] Danilo Mercurio and Vladimir Spokoiny. Statistical inference for time-inhomogeneous volatility models. Annals of Statistics, 32(2):577–602, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178939.

Maruyama:2005:NCG

[MS05a] Yuzo Maruyama and William E. Strawderman. A new class of generalized Bayes minimax ridge regression estimators. Annals of Statistics, 33(4):1753–1770, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250228.

Muller:2005:GFL

[MS05b] Hans-Georg Müller and Ulrich Stadtmüller. Generalized functional linear models. Annals of Statistics, 33(2):774–805, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1117114336.

Mikosch:2006:SLM

[MS06] Thomas Mikosch and Daniel Straumann. Stable limits of martingale transforms with application to the estimation of GARCH parameters. Annals of Statistics, 34(1):493–522, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1146576272.

Majumdar:2008:ODM

[MS08] Dibyen Majumdar and John Stufken. Optimal designs for mixed models in experiments based on ordered units. Annals of Statistics, 36(3):1090–1107, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1211819557.

Muller:2006:EPL

[MSW06] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Efficient prediction for linear and nonlinear autoregressive models. Annals of Statistics, 34(5):2496–2533, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1169571805.

Mammen:2004:SDA

[MT04] Enno Mammen and Alexandre B. Tsybakov. Smooth discrimination analysis. Annals of Statistics, 32(5):2340–2341, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1098883791.

Maaouia:2005:IMB

[MT05] F. Maaouia and A. Touati. Identification of multitype branching processes. Annals of Statistics, 33(6):2655–2694, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1140191669.

Muller:2009:ELF

[Mül09] Ursula U. Müller. Estimating linear functionals in nonlinear regression with responses missing at random. Annals of Statistics, 37(5A):2245–2277, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663754.

Meier:2009:HDA

[MvdGB09] Lukas Meier, Sara van de Geer, and Peter Bühlmann. Highdimensional additive modeling. Annals of Statistics, 37(6B): 3779–3821, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1256303527.

Meyer:2000:DFS

[MW00] Mary Meyer and Michael Woodroofe. On the degrees of freedom in shape-restricted regression. Annals of Statistics, 28(4): 1083–1104, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015956708.

Marchetti:2009:MRI

[MW09] Giovanni M. Marchetti and Nanny Wermuth. Matrix representations and independencies in directed acyclic graphs. Annals of Statistics, 37(2):961–978, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693156.

Meinshausen:2009:LTR

[MY09] Nicolai Meinshausen and Bin Yu. Lasso-type recovery of sparse representations for high-dimensional data. Annals of Statistics, 37(1):246–270, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115934.

Mykland:2003:FOS

[Myk03] Per Aslak Mykland. Financial options and statistical prediction intervals. Annals of Statistics, 31(5):1413–1438, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1065705113.

Meng:2002:SOU

[MZ02] Xiao-Li Meng and Alan M. Zaslavsky. Single observation unbiased priors. Annals of Statistics, 30(5):1345–1375, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1035844979.

Mykland:2006:ADI

[MZ06] Per Aslak Mykland and Lan Zhang. ANOVA for diffusions and Itô processes. Annals of Statistics, 34(4):1931–1963, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1162567638.

Nadler:2008:FSA

[Nad08] Boaz Nadler. Finite sample approximation results for principal component analysis: A matrix perturbation approach. Annals of Statistics, 36(6):2791–2817, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1231165185.

Naito:2004:SDE

[Nai04] Kanta Naito. Semiparametric density estimation by local L_{2}- fitting. Annals of Statistics, 32(3):1162–1191, June 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1085408498.

Nau:2006:SIP

[Nau06] Robert Nau. The shape of incomplete preferences. Annals of Statistics, 34(5):2430–2448, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571803.

Nieto-Barajas:2004:NRM

[NBPW04] Luis E. Nieto-Barajas, Igor Prünster, and Stephen G. Walker. Normalized random measures driven by increasing additive processes. Annals of Statistics, 32(6):2343–2360, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1107794871.

Neumeyer:2003:NCR

[ND03] Natalie Neumeyer and Holger Dette. Nonparametric comparison of regression curves: an empirical process approach. Annals of Statistics, 31(3):880–920, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1056562466.

Neal:2003:SS

[Nea03] Radford M. Neal. Slice sampling. Annals of Statistics, 31(3):705– 767, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1056562461.

Nan:2004:IBC

[NEW04] Bin Nan, Mary J. Emond, and Jon A. Wellner. Information bounds for Cox regression models with missing data. Annals of Statistics, 32(2):723–753, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1083178944.

Nishiyama:2009:ATS

[Nis09] Yoichi Nishiyama. Asymptotic theory of semiparametric Zestimators for stochastic processes with applications to ergodic diffusions and time series. Annals of Statistics, 37(6A):3555– 3579, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515396.

Nan:2009:ATS

[NKY09] Bin Nan, John D. Kalbfleisch, and Menggang Yu. Asymptotic theory for the semiparametric accelerated failure time model with missing data. Annals of Statistics, 37(5A):2351–2376, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663758.

Nordman:2004:OSS

[NL04] Daniel J. Nordman and Soumendra N. Lahiri. On optimal spatial subsample size for variance estimation. Annals of Statistics, 32 (5):1981–2027, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883779.

Nordman:2006:FDE

[NL06] Daniel J. Nordman and Soumendra N. Lahiri. A frequency domain empirical likelihood for short- and long-range dependence. Annals of Statistics, 34(6):3019–3050, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1179935073.

Nobile:2004:PDN

[Nob04] Agostino Nobile. On the posterior distribution of the number of components in a finite mixture. Annals of Statistics, 32(5): 2044–2073, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1098883781.

Nordman:2009:NSB

[Nor09] Daniel J. Nordman. A note on the stationary bootstrap’s variance. Annals of Statistics, 37(1):359–370, February 2009. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1232115938.

Noubiap:2001:ACG

[NS01] Roger Fandom Noubiap and Wilfried Seidel. An algorithm for calculating Γ-minimax decision rules under generalized moment conditions. Annals of Statistics, 29(4):1094–1116, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013699995.

Nussbaum:2009:CLB

[NS09] Michael Nussbaum and Arleta Szkoła. The Chernoff lower bound for symmetric quantum hypothesis testing. Annals of Statistics, 37(2):1040–1057, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693159.

Nardi:2008:DMA

[NSY08] Yuval Nardi, David O. Siegmund, and Benjamin Yakir. The distribution of maxima of approximately Gaussian random fields. Annals of Statistics, 36(3):1375–1403, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1211819568.

Nguyen:2009:SLF

[NWJ09] XuanLong Nguyen, Martin J. Wainwright, and Michael I. Jordan. On surrogate loss functions and f-divergences. Annals of Statistics, 37(2):876–904, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693153.

Omelka:2009:IKE

[OGV09] Marek Omelka, Irène Gijbels, and Noël Veraverbeke. Improved kernel estimation of copulas: Weak convergence and goodnessof-fit testing. Annals of Statistics, 37(5B):3023–3058, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1247836676.

Olshen:2007:TSR

[Ols07] Richard A. Olshen. Tree-structured regression and the differentiation of integrals. Annals of Statistics, 35(1):1–12, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1181100178.

Olszewski:2009:NT

[OS09] Wojciech Olszewski and Alvaro Sandroni. A nonmanipulable test. Annals of Statistics, 37(2):1013–1039, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1236693158.

Osius:2009:AIS

[Osi09] Gerhard Osius. Asymptotic inference for semiparametric association models. Annals of Statistics, 37(1):459–489, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115942.

Ossiander:2000:SEM

[OW00] Mina Ossiander and Edward C. Waymire. Statistical estimation for multiplicative cascades. Annals of Statistics, 28(6): 1533–1560, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957469.

Owen:2000:ALH

[Owe00] Art B. Owen. Assessing linearity in high dimensions. Annals of Statistics, 28(1):1–19, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1016120362.

Owen:2008:LAS

[Owe08] Art B. Owen. Local antithetic sampling with scrambled nets. Annals of Statistics, 36(5):2319–2343, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1223908094.

Owen:2009:KPM

[Owe09] Art B. Owen. Karl Pearson’s meta-analysis revisited. Annals of Statistics, 37(6B):3867–3892, December 2009. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1256303530.

Papadopoulos:2005:CSF

[PA05] Savas Papadopoulos and Yasuo Amemiya. Correlated samples with fixed and nonnormal latent variables. Annals of Statistics, 33(6):2732–2757, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191671.

Panaretos:2009:RTU

[Pan09] Victor M. Panaretos. On random tomography with unobservable projection angles. Annals of Statistics, 37(6A):3272–3306, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1250515387.

Patilea:2001:CMM

[Pat01] Valentin Patilea. Convex models, MLS and misspecification. Annals of Statistics, 29(1):94–123, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/996986503.

Paulo:2005:DPG

[Pau05] Rui Paulo. Default priors for Gaussian processes. Annals of Statistics, 33(2):556–582, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114329.

Paul:2008:PFS

[PBHT08] Debashis Paul, Eric Bair, Trevor Hastie, and Robert Tibshirani. “preconditioning” for feature selection and regression in highdimensional problems. Annals of Statistics, 36(4):1595–1618, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1216237293.

Peng:2004:ELB

[Pen04] Liang Peng. Empirical-likelihood-based confidence interval for the mean with a heavy-tailed distribution. Annals of Statistics, 32(3):1192–1214, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408499.

Pensky:2006:FOB

[Pen06] Marianna Pensky. Frequentist optimality of Bayesian wavelet shrinkage rules for Gaussian and non-Gaussian noise. Annals of Statistics, 34(2):769–807, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418240.

Pere:2006:D

[Per06] Anneli Pere. Discussion. Annals of Statistics, 34(5):2105–2112, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1169571788.

Park:2002:LLD

[PKJ02] B. U. Park, W. C. Kim, and M. C. Jones. On local likelihood density estimation. Annals of Statistics, 30(5):1480–1495, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844984.

Pons:2003:ECR

[Pon03] Odile Pons. Estimation in a Cox regression model with a changepoint according to a threshold in a covariate. Annals of Statistics, 31(2):442–463, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1051027876.

Pitts:2000:NER

[PP00] Susan M. Pitts and Konstadinos Politis. Nonparametric estimation in renewal theory. II: Solutions of renewal-type equations. Annals of Statistics, 28(1):88–115, February 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1016120366.

Paul:2009:CRM

[PP09] Debashis Paul and Jie Peng. Consistency of restricted maximum likelihood estimators of principal components. Annals of Statistics, 37(3):1229–1271, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1239369021.

Peng:2006:CRH

[PQ06] Liang Peng and Yongcheng Qi. Confidence regions for high quantiles of a heavy tailed distribution. Annals of Statistics, 34(4): 1964–1986, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567639.

Patilea:2006:PLE

[PR06] Valentin Patilea and Jean-Marie Rolin. Product-limit estimators of the survival function with twice censored data. Annals of Statistics, 34(2):925–938, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418246.

Papaspiliopoulos:2008:SGS

[PR08a] Omiros Papaspiliopoulos and Gareth Roberts. Stability of the Gibbs sampler for Bayesian hierarchical models. Annals of Statistics, 36(1):95–117, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201877295.

Privault:2008:SED

[PR08b] Nicolas Privault and Anthony Réveillac. Stein estimation for the drift of Gaussian processes using the Malliavin calculus. Annals of Statistics, 36(5):2531–2550, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1223908102.

Pronzato:2000:AOO

[Pro00] Luc Pronzato. Adaptive optimization and D-optimum experimental design. Annals of Statistics, 28(6):1743–1761, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015957479.

Peskir:2000:STP

[PS00] G. Peskir and A. N. Shiryaev. Sequential testing problems for Poisson processes. Annals of Statistics, 28(3):837–859, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1015952000.

Polzehl:2003:IDP

[PS03] Jörg Polzehl and Vladimir Spokoiny. Image denoising: pointwise adaptive approach. Annals of Statistics, 31(1):30–57, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1046294457.

Phillips:2004:LWE

[PS04] Peter C. B. Phillips and Katsumi Shimotsu. Local Whittle estimation in nonstationary and unit root cases. Annals of Statistics, 32(2):656–692, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178942.

Pensky:2009:FDP

[PS09] Marianna Pensky and Theofanis Sapatinas. Functional deconvolution in a periodic setting: Uniform case. Annals of Statistics, 37 (1):73–104, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1232115928.

Park:2009:EGS

[PSS09] Yonil Park, Sergey Sheetlin, and John L. Spouge. Estimating the Gumbel scale parameter for local alignment of random sequences by importance sampling with stopping times. Annals of Statistics, 37(6A):3697–3714, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1250515402.

Picard:2000:ACI

[PT00] Dominique Picard and Karine Tribouley. Adaptive confidence interval for pointwise curve estimation. Annals of Statistics, 28 (1):298–335, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1016120374.

Phoa:2009:QFF

[PX09] Frederick K. H. Phoa and Hongquan Xu. Quarter-fraction factorial designs constructed via quaternary codes. Annals of Statistics, 37(5A):2561–2581, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663766.

Qian:2009:CNS

[QAW09] Peter Z. G. Qian, Mingyao Ai, and C. F. Jeff Wu. Construction of nested space-filling designs. Annals of Statistics, 37(6A): 3616–3643, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515399.

Radchenko:2008:MRA

[Rad08] Peter Radchenko. Mixed-rates asymptotics. Annals of Statistics, 36(1):287–309, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201877302.

Ranjan:2009:ECR

[RBD09] Pritam Ranjan, Derek R. Bingham, and Angela M. Dean. Existence and construction of randomization defining contrast subspaces for regular factorial designs. Annals of Statistics, 37(6A): 3580–3599, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1250515397.

Rubin-Bleuer:2005:TPF

[RBK05] Susana Rubin-Bleuer and Ioana Schiopu Kratina. On the twophase framework for joint model and design-based inference. Annals of Statistics, 33(6):2789–2810, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1140191673.

Reboul:2005:EFU

[Reb05] L. Reboul. Estimation of a function under shape restrictions. Applications to reliability. Annals of Statistics, 33(3):1330–1356, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1120224104.

Reid:2003:ATI

[Rei03] N. Reid. Asymptotics and the theory of inference. Annals of Statistics, 31(6):1695–2095, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290325.

Reiss:2008:AEN

[Rei08] Markus Reiß. Asymptotic equivalence for nonparametric regression with multivariate and random design. Annals of Statistics, 36(4):1957–1982, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237305.

Ren:2003:RME

[Ren03] Jian-Jian Ren. Regression M-estimators with non-i.i.d. doubly censored data. Annals of Statistics, 31(4):1186–1219, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1059655911.

Ren:2008:WEL

[Ren08] Jian-Jian Ren. Weighted empirical likelihood in some twosample semiparametric models with various types of censored data. Annals of Statistics, 36(1):147–166, February 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1201877297.

Regazzini:2002:TNA

[RGN02] Eugenio Regazzini, Alessandra Guglielmi, and Giulia Di Nunno. Theory and numerical analysis for exact distributions of functionals of a Dirichlet process. Annals of Statistics, 30(5):1376–1411, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844980.

Rinaldo:2009:PRF

[Rin09] Alessandro Rinaldo. Properties and refinements of the fused lasso. Annals of Statistics, 37(5B):2922–2952, October 2009. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1247836673.

Ritov:2007:DDS

[Rit07] Ya’acov Ritov. Discussion: The Dantzig selector: Statistical estimation when p is much larger than n. Annals of Statistics, 35(6): 2370–2372, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201012962.

Ray:2005:TMN

[RL05] Surajit Ray and Bruce G. Lindsay. The topography of multivariate normal mixtures. Annals of Statistics, 33(5):2042–2065, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936556.

Regazzini:2003:DRM

[RLP03] Eugenio Regazzini, Antonio Lijoi, and Igor Prünster. Distributional results for means of normalized random measures with independent increments. Annals of Statistics, 31(2):560–585, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1051027881.

Rajaratnam:2008:FCE

[RMC08] Bala Rajaratnam, Hélène Massam, and Carlos M. Carvalho. Flexible covariance estimation in graphical Gaussian models. Annals of Statistics, 36(6):2818–2849, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1231165186.

Rao:2008:SEI

[RMSE08] N. Raj Rao, James A. Mingo, Roland Speicher, and Alan Edelman. Statistical eigen-inference from large Wishart matrices. Annals of Statistics, 36(6):2850–2885, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1231165187.

Robinson:2005:EII

[Rob05] P. M. Robinson. Efficiency improvements in inference on stationary and nonstationary fractional time series. Annals of Statistics, 33(4):1800–1842, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1123250230.

Robinson:2008:MLW

[Rob08] P. M. Robinson. Multiple local Whittle estimation in stationary systems. Annals of Statistics, 36(5):2508–2530, October 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1223908101.

Robert:2009:ILC

[Rob09] Christian Y. Robert. Inference for the limiting cluster size distribution of extreme values. Annals of Statistics, 37(1):271–310, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115935.

Rohde:2008:AGF

[Roh08] Angelika Rohde. Adaptive goodness-of-fit tests based on signed ranks. Annals of Statistics, 36(3):1346–1374, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1211819567.

Romano:2005:OTE

[Rom05] Joseph P. Romano. Optimal testing of equivalence hypotheses. Annals of Statistics, 33(3):1036–1047, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1120224094.

Roueff:2005:NEM

[RR05] François Roueff and Tobias Rydén. Nonparametric estimation of mixing densities for discrete distributions. Annals of Statistics, 33(5):2066–2108, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1132936557.

Robinson:2003:SAT

[RRY03] J. Robinson, E. Ronchetti, and G. A. Young. Saddlepoint approximations and tests based on multivariate M-estimates. Annals of Statistics, 31(4):1154–1169, August 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1059655909.

Richardson:2002:AGM

[RS02] Thomas Richardson and Peter Spirtes. Ancestral graph Markov models. Annals of Statistics, 30(4):962–1030, August 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031689015.

Romano:2006:SPC

[RS06a] Joseph P. Romano and Azeem M. Shaikh. Stepup procedures for control of generalizations of the familywise error rate. Annals of Statistics, 34(4):1850–1873, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1162567636.

Rubin:2006:AIM

[RS06b] Donald B. Rubin and Elizabeth A. Stuart. Affinely invariant matching methods with discriminant mixtures of proportional ellipsoidally symmetric distributions. Annals of Statistics, 34(4): 1814–1826, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567634.

Rudin:2007:ABA

[RSD07] Cynthia Rudin, Robert E. Schapire, and Ingrid Daubechies. Analysis of boosting algorithms using the smooth margin function. Annals of Statistics, 35(6):2723–2768, December 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1201012978.

Rosenberg:2002:BOM

[RSV02] Dinah Rosenberg, Eilon Solan, and Nicolas Vieille. Blackwell optimality in Markov decision processes with partial observation. Annals of Statistics, 30(4):1178–1193, August 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1031689022.

Rosenberg:2004:ASO

[RSV04] Dinah Rosenberg, Eilon Solan, and Nicolas Vieille. Approximating a sequence of observations by a simple process. Annals of Statistics, 32(6):2742–2775, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794886.

Robins:2006:ANC

[RvdV06] James Robins and Aad van der Vaart. Adaptive nonparametric confidence sets. Annals of Statistics, 34(1):229–253, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1146576262.

Romano:2000:FSN

[RW00] Joseph P. Romano and Michael Wolf. Finite sample nonparametric inference and large sample efficiency. Annals of Statistics, 28(3):756–778, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015951997.

Ruckstuhl:2001:RFB

[RW01] A. F. Ruckstuhl and A. H. Welsh. Robust fitting of the binomial model. Annals of Statistics, 29(4):1117–1136, August 2001. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1013699996.

Rao:2002:ELB

[RW02] J. N. K. Rao and Qihua Wang. Empirical likelihood-based inference under imputation for missing response data. Annals of Statistics, 30(3):896–924, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028674845.

Romano:2007:CGE

[RW07] Joseph P. Romano and Michael Wolf. Control of generalized error rates in multiple testing. Annals of Statistics, 35(4):1378–1408, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1188405615.

Robinson:2006:PML

[RZ06] Peter M. Robinson and Paolo Zaffaroni. Pseudo-maximum likelihood estimation of ARCH(∞) models. Annals of Statistics, 34 (3):1049–1074, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540742.

Rosset:2007:PLR

[RZ07] Saharon Rosset and Ji Zhu. Piecewise linear regularized solution paths. Annals of Statistics, 35(3):1012–1030, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1185303996.

Steinwart:2009:CSV

[SA09] Ingo Steinwart and Marian Anghel. Consistency of support vector machines for forecasting the evolution of an unknown ergodic dynamical system from observations with unknown noise. Annals of Statistics, 37(2):841–875, April 2009. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693152.

Sarkar:2002:SRF

[Sar02] Sanat K. Sarkar. Some results on false discovery rate in stepwise multiple testing procedures. Annals of Statistics, 30(1):239–257, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1015362192.

Sarkar:2006:FDF

[Sar06] Sanat K. Sarkar. False discovery and false nondiscovery rates in single-step multiple testing procedures. Annals of Statistics, 34 (1):394–415, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1146576268.

Sarkar:2007:SPC

[Sar07] Sanat K. Sarkar. Stepup procedures controlling generalized FWER and generalized FDR. Annals of Statistics, 35(6):2405– 2420, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201012966.

Sarkar:2008:GST

[Sar08] Sanat K. Sarkar. Generalizing Simes’ test and Hochberg’s stepup procedure. Annals of Statistics, 36(1):337–363, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877304.

Shah:2005:UOP

[SBR05] Kirti R. Shah, Mausumi Bose, and Damaraju Raghavarao. Universal optimality of Patterson’s crossover designs. Annals of Statistics, 33(6):2854–2872, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191676.

Salibian-Barrera:2008:HBP

[SBY08] Matías Salibian-Barrera and Víctor J. Yohai. High breakdown point robust regression with censored data. Annals of Statistics, 36(1):118–146, February 2008. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201877296.

Shi:2009:DSE

[SBY09] Tao Shi, Mikhail Belkin, and Bin Yu. Data spectroscopy: Eigenspaces of convolution operators and clustering. Annals of Statistics, 37(6B):3960–3984, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1256303533.

Salibian-Barrera:2002:BRE

[SBZ02] Matias Salibian-Barrera and Ruben H. Zamar. Bootstrapping robust estimates of regression. Annals of Statistics, 30(2):556–582, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1021379865.

Salibian-Barrera:2004:UAR

[SBZ04] Matias Salibian-Barrera and Ruben H. Zamar. Uniform asymptotics for robust location estimates when the scale is unknown. Annals of Statistics, 32(4):1434–1447, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1091626174.

Salibian-Barrera:2006:D

[SBZ06] Matias Salibian-Barrera and Ruben H. Zamar. Discussion. Annals of Statistics, 34(5):2113–2118, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1169571789.

Scheike:2001:GAR

[Sch01] Thomas H. Scheike. A generalized additive regression model for survival times. Annals of Statistics, 29(5):1344–1360, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1013203457.

Schennach:2007:PEE

[Sch07] Susanne M. Schennach. Point estimation with exponentially tilted empirical likelihood. Annals of Statistics, 35(2):634–672, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1183667287.

Scricciolo:2006:CRB

[Scr06] Catia Scricciolo. Convergence rates for Bayesian density estimation of infinite-dimensional exponential families. Annals of Statistics, 34(6):2897–2920, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935069.

Sweeting:2006:NPP

[SDG06] Trevor J. Sweeting, Gauri S. Datta, and Malay Ghosh. Nonsubjective priors via predictive relative entropy regret. Annals of Statistics, 34(1):441–468, February 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1146576270.

Sung:2007:MCL

[SG07] Yun Ju Sung and Charles J. Geyer. Monte Carlo likelihood inference for missing data models. Annals of Statistics, 35 (3):990–1011, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1185303995.

Sarkar:2009:GFD

[SG09] Sanat K. Sarkar and Wenge Guo. On a generalized false discovery rate. Annals of Statistics, 37(3):1545–1565, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1239369031.

Sun:2009:PHM

[SGM09] Yanqing Sun, Peter B. Gilbert, and Ian W. McKeague. Proportional hazards models with continuous marks. Annals of Statistics, 37(1):394–426, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115940.

Shapiro:2000:ACL

[Sha00] Alexander Shapiro. On the asymptotics of constrained local Mestimators. Annals of Statistics, 28(3):948–960, June 2000. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1015952006.

Shimodaira:2004:AUT

[Shi04] Hidetoshi Shimodaira. Approximately unbiased tests of regions using multistep-multiscale bootstrap resampling. Annals of Statistics, 32(6):2616–2641, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1107794881.

Siegmund:2003:HRS

[Sie03] David Siegmund. Herbert Robbins and sequential analysis: invited paper. Annals of Statistics, 31(2):349–365, April 2003. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1051027870.

Song:2009:AOT

[SKF09] Rui Song, Michael R. Kosorok, and Jason P. Fine. On asymptotically optimal tests under loss of identifiability in semiparametric models. Annals of Statistics, 37(5A):2409–2444, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663760.

Skouras:2000:SCN

[Sko00] K. Skouras. Strong consistency in nonlinear stochastic regression models. Annals of Statistics, 28(3):871–879, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015952002.

Sun:2000:CBG

[SLM00] Jiayang Sun, Catherine Loader, and William P. McCormick. Confidence bands in generalized linear models. Annals of Statistics, 28(2):429–460, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1016218225.

Senturk:2006:ICA

[ŞM06a] Damla Şentürk and Hans-Georg Müller. Inference for covariate adjusted regression via varying coefficient models. Annals of Statistics, 34(2):654–679, April 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1151418236.

Straumann:2006:QML

[SM06b] Daniel Straumann and Thomas Mikosch. Quasi-maximum-likelihood estimation in conditionally heteroscedastic time series: A stochastic recurrence equations approach. Annals of Statistics, 34(5):2449–2495, October 2006. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571804.

Schwartzman:2008:IEE

[SMT08] Armin Schwartzman, Walter F. Mascarenhas, and Jonathan E. Taylor. Inference for eigenvalues and eigenvectors of Gaussian symmetric matrices. Annals of Statistics, 36(6):2886–2919, December 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1231165188.

Song:2009:TCI

[Son09] Kyungchul Song. Testing conditional independence via Rosenblatt transforms. Annals of Statistics, 37(6B):4011–4045, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1256303535.

Shimotsu:2005:ELW

[SP05] Katsumi Shimotsu and Peter C. B. Phillips. Exact local Whittle estimation of fractional integration. Annals of Statistics, 33(4): 1890–1933, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1123250232.

Speed:2002:JWT

[Spe02] T. P. Speed. John w. Tukey’s contributions to analysis of variance. Annals of Statistics, 30(6):1649–1665, December 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1043351252;   http://www.jstor.org/stable/1558735.

Spokoiny:2000:ADE

[Spo00] Vladimir G. Spokoiny. Adaptive drift estimation for nonparametric diffusion model. Annals of Statistics, 28(3):815–836, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1015951999.

Spokoiny:2009:MLC

[Spo09] Vladimir Spokoiny. Multiscale local change point detection with applications to value-at-risk. Annals of Statistics, 37(3): 1405–1436, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369026.

Szekely:2007:MTD

[SRB07] Gábor J. Székely, Maria L. Rizzo, and Nail K. Bakirov. Measuring and testing dependence by correlation of distances. Annals of Statistics, 35(6):2769–2794, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012979.

Srivastava:2003:SWM

[Sri03] M. S. Srivastava. Singular Wishart and multivariate beta distributions. Annals of Statistics, 31(5):1537–1560, October 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1065705118.

Settimi:2000:GMC

[SS00] Raffaella Settimi and Jim Q. Smith. Geometry, moments and conditional independence trees with hidden variables. Annals of Statistics, 28(4):1179–1205, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015956712.

Steinwart:2007:FRS

[SS07] Ingo Steinwart and Clint Scovel. Fast rates for support vector machines using Gaussian kernels. Annals of Statistics, 35 (2):575–607, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1183667285.

Studer:2005:NRP

[SSG05] M. Studer, B. Seifert, and T. Gasser. Nonparametric regression penalizing deviations from additivity. Annals of Statistics, 33 (3):1295–1329, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1120224103.

Singh:2009:AHE

[SSN09] Aarti Singh, Clayton Scott, and Robert Nowak. Adaptive Hausdorff estimation of density level sets. Annals of Statistics, 37(5B): 2760–2782, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247836668.

Shafie:2003:RSR

[SSSW03] K. Shafie, D. Siegmund, B. Sigal, and K. J. Worsley. Rotation space random fields with an application to fMRI data. Annals of Statistics, 31(6):1732–1771, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1074290326.

Spruill:2001:LAO

[ST01] M. C. Spruill and Renjin Tu. Locally asymptotically optimal designs for testing in logistic regression. Annals of Statistics, 29 (4):1050–1057, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1013699991.

Stufken:2007:CET

[ST07] John Stufken and Boxin Tang. Complete enumeration of twolevel orthogonal arrays of strength d with d + 2 constraints. Annals of Statistics, 35(2):793–814, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1183667294.

Stephens:2000:BAM

[Ste00] Matthew Stephens. Bayesian analysis of mixture models with an unknown number of components — an alternative to reversible jump methods. Annals of Statistics, 28(1):40–74, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1016120364.

Stein:2002:SEK

[Ste02] Michael L. Stein. The screening effect in kriging. Annals of Statistics, 30(1):298–323, February 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1015362194.

Storey:2003:PFD

[Sto03] John D. Storey. The positive false discovery rate: a Bayesian interpretation and the q-value. Annals of Statistics, 31(6):2013– 2035, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1074290335.

Sasabushi:2003:THM

[STT03] Shoichi Sasabushi, Koji Tanaka, and Takeshi Tsukamoto. Testing homogeneity of multivariate normal mean vectors under an order restriction when the covariance matrices are common but unknown. Annals of Statistics, 31(5):1517–1536, October 2003. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1065705117.

Spokoiny:2009:PTP

[SV09] Vladimir Spokoiny and Céline Vial. Parameter tuning in pointwise adaptation using a propagation approach. Annals of Statistics, 37(5B):2783–2807, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247836669.

Shen:2001:RCP

[SW01] Xiaotong Shen and Larry Wasserman. Rates of convergence of posterior distributions. Annals of Statistics, 29(3):687–714, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1009210686.

Schick:2004:EIL

[SW04] Anton Schick and Wolfgang Wefelmeyer. Estimating invariant laws of linear processesby U-statistics. Annals of Statistics, 32 (2):603–632, April 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1083178940.

Shen:2006:DLR

[SW06] Xiaotong Shen and Lifeng Wang. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2677–2680, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1179935059.

Schick:2007:URC

[SW07a] Anton Schick and Wolfgang Wefelmeyer. Uniformly root-n consistent density estimators for weakly dependent invertible linear processes. Annals of Statistics, 35(2):815–843, April 2007. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1183667295.

Shao:2007:AST

[SW07b] Xiaofeng Shao and Wei Biao Wu. Asymptotic spectral theory for nonlinear time series. Annals of Statistics, 35(4):1773–1801, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1188405630. See correction [ZSW22].

Singh:2005:CII

[SXS05] Kesar Singh, Minge Xie, and William E. Strawderman. Combining information from independent sources through confidence distributions. Annals of Statistics, 33(1):159–183, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1112967703.

Siegmund:2000:AVL

[SY00] David Siegmund and Benjamin Yakir. Approximate p-values for local sequence alignments. Annals of Statistics, 28(3):657–680, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1015951993. See correction [SY03].

Siegmund:2003:CAV

[SY03] David Siegmund and Benjamin Yakir. Correction: “Approximate p-values for local sequence alignments”. Annals of Statistics, 31 (3):1027–1031, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1056562473. See [SY00].

Serfling:2000:GNS

[SZ00a] Robert Serfling and Yijun Zuo. General notions of statistical depth function. Annals of Statistics, 28(2):461–482, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1016218226.

Serfling:2000:SPC

[SZ00b] Robert Serfling and Yijun Zuo. Structural properties and convergence results for contours of sample statistical depth functions. Annals of Statistics, 28(2):483–499, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1016218227.

Shoung:2001:LSE

[SZ01] Jyh-Ming Shoung and Cun-Hui Zhang. Least squares estimators of the mode of a unimodal regression function. Annals of Statistics, 29(3):648–665, June 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1009210684.

Stute:2005:NCS

[SZ05] Winfried Stute and Li-Xing Zhu. Nonparametric checks for single-index models. Annals of Statistics, 33(3):1048–1083, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1120224095.

Sun:2007:ECM

[SZT07] Yan Sun, Wenyang Zhang, and Howell Tong. Estimation of the covariance matrix of random effects in longitudinal studies. Annals of Statistics, 35(6):2795–2814, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1201012980.

Taupin:2001:SPE

[Tau01] Marie-Luce Taupin. Semi-parametric estimation in the nonlinear structural errors-in-variables model. Annals of Statistics, 29 (1):66–93, February 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/996986502.

Thompson:2006:D

[Tho06] Mary Lou Thompson. Discussion. Annals of Statistics, 34(5): 2119–2125, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1169571790.

Tripathi:2003:TCM

[TK03] Gautam Tripathi and Yuichi Kitamura. Testing conditional moment restrictions. Annals of Statistics, 31(6):2059–2095, December 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1074290337.

Tokdar:2009:CRE

[TMG09] Surya T. Tokdar, Ryan Martin, and Jayanta K. Ghosh. Consistency of a recursive estimate of mixing distributions. Annals of Statistics, 37(5A):2502–2522, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663763.

Tai:2006:MEB

[TS06] Yu Chuan Tai and Terence P. Speed. A multivariate empirical Bayes statistic for replicated microarray time course data. Annals of Statistics, 34(5):2387–2412, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571801.

Tsao:2004:BCP

[Tsa04] Min Tsao. Bounds on coverage probabilities of the empirical likelihood ratio confidence regions. Annals of Statistics, 32(3): 1215–1221, June 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1085408500.

Tsybakov:2004:OAC

[Tsy04] Alexander B. Tsybakov. Optimal aggregation of classifiers in statistical learning. Annals of Statistics, 32(1):135–166, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1079120131.

Tsybakov:2006:DLR

[Tsy06] A. B. Tsybakov. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2681–2687, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935060.

Tatsuoka:2000:UFM

[TT00] Kay S. Tatsuoka and David E. Tyler. On the uniqueness of Sfunctionals and M-functionals under nonelliptical distributions. Annals of Statistics, 28(4):1219–1243, August 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1015956714.

Tudor:2007:SAF

[TV07] Ciprian A. Tudor and Frederi G. Viens. Statistical aspects of the fractional stochastic calculus. Annals of Statistics, 35 (3):1183–1212, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1185304003.

Tsybakov:2005:SRP

[TvdG05] A. B. Tsybakov and S. A. van de Geer. Square root penalty: Adaptation to the margin in classification and in edge estimation. Annals of Statistics, 33(3):1203–1224, June 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1120224100.

Taylor:2008:RFM

[TW08] J. E. Taylor and K. J. Worsley. Random fields of multivariate test statistics, with applications to shape analysis. Annals of Statistics, 36(1):1–27, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201877292.

vandeGeer:2006:DLR

[vdG06] Sara van de Geer. Discussion: Local Rademacher complexities and oracle inequalities in risk minimization. Annals of Statistics, 34(6):2688–2696, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1179935061.

vandeGeer:2008:HDG

[vdG08] Sara A. van de Geer. High-dimensional generalized linear models and the lasso. Annals of Statistics, 36(2):614–645, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1205420513.

vanderVaart:2002:SWL

[vdV02] Aad van der Vaart. The statistical work of Lucien Le Cam. Annals of Statistics, 30(3):631–682, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028674836.

vanderVaart:2008:RCP

[vdVvZ08] A. W. van der Vaart and J. H. van Zanten. Rates of contraction of posterior distributions based on Gaussian process priors. Annals of Statistics, 36(3):1435–1463, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1211819570.

vanderVaart:2009:ABE

[vdVvZ09] A. W. van der Vaart and J. H. van Zanten. Adaptive Bayesian estimation using a Gaussian random field with inverse gamma bandwidth. Annals of Statistics, 37(5B):2655–2675, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1247836664.

VanderLaan:2003:CSR

[VJ03] Mark J. Van der Laan and Nicholas P. Jewell. Current status and right-censored data structures when observing a marker at the censoring time. Annals of Statistics, 31(2):512–535, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1051027879.

vonLuxburg:2008:CSC

[vLBB08] Ulrike von Luxburg, Mikhail Belkin, and Olivier Bousquet. Consistency of spectral clustering. Annals of Statistics, 36 (2):555–586, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1205420511.

Vovk:2009:LPL

[VNG09] Vladimir Vovk, Ilia Nouretdinov, and Alex Gammerman. Online predictive linear regression. Annals of Statistics, 37(3): 1566–1590, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369032.

VanderWeele:2009:MSC

[VR09] Tyler J. VanderWeele and James M. Robins. Minimal sufficient causation and directed acyclic graphs. Annals of Statistics, 37 (3):1437–1465, June 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1239369027.

VanBellegem:2008:LAE

[VvS08] Sébastien Van Bellegem and Rainer von Sachs. Locally adaptive estimation of evolutionary wavelet spectra. Annals of Statistics, 36(4):1879–1924, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237303.

Voss:2001:CER

[VW01] Daniel T. Voss and Weizhen Wang. Control of error rates in adaptive analysis of orthogonal saturated designs. Annals of Statistics, 29(4):1058–1065, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013699992.

Walther:2001:MML

[Wal01] Guenther Walther. Multiscale maximum likelihood analysis of a semiparametric model, with applications. Annals of Statistics, 29 (5):1297–1319, October 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1013203455.

Walker:2004:NAB

[Wal04] Stephen Walker. New approaches to Bayesian consistency. Annals of Statistics, 32(5):2028–2043, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1098883780.

Wang:2002:ANG

[Wan02] Yazhen Wang. Asymptotic nonequivalence of GARCH models and diffusions. Annals of Statistics, 30(3):754–783, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1028674841.

Wang:2004:ENM

[Wan04] Liqun Wang. Estimation of nonlinear models with Berkson measurement errors. Annals of Statistics, 32(6):2559–2579, December 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1107794879.

Wang:2008:EMV

[WBCL08] Lie Wang, Lawrence D. Brown, T. Tony Cai, and Michael Levine. Effect of mean on variance function estimation in nonparametric regression. Annals of Statistics, 36(2):646–664, April 2008. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1205420514.

Wang:2009:ELE

[WC09] Dong Wang and Song Xi Chen. Empirical likelihood for estimating equations with missing values. Annals of Statistics, 37 (1):490–517, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1232115943.

Wegkamp:2003:MSN

[Weg03] Marten Wegkamp. Model selection in nonparametric regression. Annals of Statistics, 31(1):252–273, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1046294464.

Wang:2009:ICQ

[WF09] Huixia Judy Wang and Mendel Fygenson. Inference for censored quantile regression models in longitudinal studies. Annals of Statistics, 37(2):756–781, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1236693149.

Wei:2006:CGC

[WH06a] Ying Wei and Xuming He. Conditional growth charts. Annals of Statistics, 34(5):2069–2097, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571786.

Wei:2006:R

[WH06b] Ying Wei and Xuming He. Rejoinder. Annals of Statistics, 34(5): 2126–2131, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1169571791.

Walker:2007:RCP

[WLP07] Stephen G. Walker, Antonio Lijoi, and Igor Prünster. On rates of convergence for posterior distributions in infinite-dimensional models. Annals of Statistics, 35(2):738–746, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1183667291.

Wu:2002:KDE

[WM02] Wei Biao Wu and Jan Mielniczuk. Kernel density estimation for linear processes. Annals of Statistics, 30(5):1441–1459, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844982.

Wang:2007:OOD

[WM07] Haonan Wang and J. S. Marron. Object oriented data analysis: Sets of trees. Annals of Statistics, 35(5):1849–1873, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1194461714.

Wasserman:2009:HDV

[WR09] Larry Wasserman and Kathryn Roeder. High-dimensional variable selection. Annals of Statistics, 37(5A):2178–2201, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663752.

Wu:2005:BRS

[Wu05] Wei Biao Wu. On the Bahadur representation of sample quantiles for dependent sequences. Annals of Statistics, 33(4):1934–1963, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1123250233.

Wu:2007:MEL

[Wu07] Wei Biao Wu. M-estimation of linear models with dependent errors. Annals of Statistics, 35(2):495–521, April 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1183667282.

Wu:2008:FDC

[Wu08] Wei Biao Wu. On false discovery control under dependence. Annals of Statistics, 36(1):364–380, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1201877305.

Wang:2000:PLC

[WW00] Hsiuying Wang and Michael Woodroofe. The problem of low counts in a signal plus noise model. Annals of Statistics, 28(6): 1561–1569, December 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015957470.

Wu:2002:CTF

[WW02] Huaiqing Wu and C. F. J. Wu. Clear two-factor interactions and minimum aberration. Annals of Statistics, 30(5):1496–1511, October 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1035844985.

Wang:2007:KWC

[WW07a] Xiao Wang and Michael Woodroofe. A Kiefer–Wolfowitz comparison theorem for Wicksell’s problem. Annals of Statistics, 35 (4):1559–1575, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1188405622.

Wu:2007:SST

[WW07b] Samuel S. Wu and Weizhen Wang. Step-up simultaneous tests for identifying active effects in orthogonal saturated designs. Annals of Statistics, 35(1):449–463, February 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1181100194.

Wu:2001:GMAa

[WX01a] C. F. J. Wu and Hongquan Xu. Generalized minimum aberration for asymmetrical fractional factorial designs. Annals of Statistics, 29(2):549–560, April 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1009210552.

Wu:2001:GMAb

[WX01b] C. F. J. Wu and Hongquan Xu. Generalized minimum aberration for asymmetrical fractional factorial designs. Annals of Statistics, 29(4):1066–1077, August 2001. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1013699993.

Wang:2007:SBK

[WY07] Li Wang and Lijian Yang. Spline-backfitted kernel smoothing of nonlinear additive autoregression model. Annals of Statistics, 35(6):2474–2503, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1201012969.

Wellner:2000:TEM

[WZ00] Jon A. Wellner and Ying Zhang. Two estimators of the mean of a counting process with panel count data. Annals of Statistics, 28(3):779–814, June 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1015951998.

Wang:2005:SLW

[WZ05] Xiaogang Wang and James V. Zidek. Selecting likelihood weights by cross-validation. Annals of Statistics, 33(2):463–500, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1117114325.

Wu:2006:DEE

[WZ06] Ying Nian Wu and Song-Chun Zhu. Discussion of “Equi-energy sampler” by Kou, Zhou and Wong. Annals of Statistics, 34(4): 1642–1645, August 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1162567626.

Wellner:2007:TLB

[WZ07] Jon A. Wellner and Ying Zhang. Two likelihood-based semiparametric estimation methods for panel count data with covariates. Annals of Statistics, 35(5):2106–2142, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1194461724.

Wang:2009:QRP

[WZZ09] Huixia Judy Wang, Zhongyi Zhu, and Jianhui Zhou. Quantile regression in partially linear varying coefficient models. Annals of Statistics, 37(6B):3841–3866, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1256303529.

Xu:2008:CDT

[XC08] Hongquan Xu and Ching-Shui Cheng. A complementary design theory for doubling. Annals of Statistics, 36(1):445–457, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201877309.

Xie:2009:SPR

[XH09] Huiliang Xie and Jian Huang. SCAD-penalized regression in high-dimensional partially linear models. Annals of Statistics, 37(2):673–696, April 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1236693146.

Xia:2007:CAE

[Xia07] Yingcun Xia. A constructive approach to the estimation of dimension reduction directions. Annals of Statistics, 35(6):2654– 2690, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201012976.

Xu:2006:BRF

[Xu06] Hongquan Xu. Blocked regular fractional factorial designs with minimum aberration. Annals of Statistics, 34(5):2534–2553, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1169571806.

Xu:2005:COM

[XW05] Hongquan Xu and C. F. J. Wu. Construction of optimal multilevel supersaturated designs. Annals of Statistics, 33(6):2811– 2836, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1140191674.

Xie:2003:AGE

[XY03] Minge Xie and Yaning Yang. Asymptotics for generalized estimating equations with large cluster sizes. Annals of Statistics, 31 (1):310–347, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1046294467.

Yang:2000:MSD

[Yan00] Yuhong Yang. Mixing strategies for density estimation. Annals of Statistics, 28(1):75–87, February 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1016120365.

Yang:2002:LC

[Yan02] Grace L. Yang. Lucien Le Cam 1924–2000. Annals of Statistics, 30(3):617–630, June 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028674835.

Yang:2007:CCV

[Yan07] Yuhong Yang. Consistency of cross validation for comparing regression procedures. Annals of Statistics, 35(6):2450–2473, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201012968.

Ye:2003:IFA

[Ye03] Kenny Q. Ye. Indicator function and its application in twolevel factorial designs. Annals of Statistics, 31(3):984–994, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1056562470.

Yajima:2009:NST

[YM09] Yoshihiro Yajima and Yasumasa Matsuda. On nonparametric and semiparametric testing for multivariate linear time series. Annals of Statistics, 37(6A):3529–3554, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1250515395.

Yao:2005:FLR

[YMW05] Fang Yao, Hans-Georg Müller, and Jane-Ling Wang. Functional linear regression analysis for longitudinal data. Annals of Statistics, 33(6):2873–2903, December 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191677.

Yu:2008:SBG

[YPM08] Kyusang Yu, Byeong U. Park, and Enno Mammen. Smooth backfitting in generalized additive models. Annals of Statistics, 36 (1):228–260, February 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1201877300.

Yang:2009:SPL

[YS09] Min Yang and John Stufken. Support points of locally optimal designs for nonlinear models with two parameters. Annals of Statistics, 37(1):518–541, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1232115944.

Yuan:2009:BFH

[Yua09] Ao Yuan. Bayesian frequentist hybrid inference. Annals of Statistics, 37(5A):2458–2501, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663762.

Yang:2002:RAN

[YZ02] Yuhong Yang and Dan Zhu. Randomized allocation with nonparametric estimation for a multi-armed bandit problem with covariates. Annals of Statistics, 30(1):100–121, February 2002. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1015362186.

Yohai:2004:RNI

[YZ04] Víctor J. Yohai and Ruben H. Zamar. Robust nonparametric inference for the median. Annals of Statistics, 32(5):1841–1857, October 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1098883774.

Zeng:2005:ARM

[ZC05a] Donglin Zeng and Jianwen Cai. Asymptotic results for maximum likelihood estimators in joint analysis of repeated measurements and survival time. Annals of Statistics, 33(5):2132–2163, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1132936559.

Zuo:2005:DWS

[ZC05b] Yijun Zuo and Hengjian Cui. Depth weighted scatter estimators. Annals of Statistics, 33(1):381–413, February 2005. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1112967710.

Zhang:2007:SIF

[ZC07] Jin-Ting Zhang and Jianwei Chen. Statistical inferences for functional data. Annals of Statistics, 35(3):1052–1079, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1185303998.

Zuo:2004:SDE

[ZCH04] Yijun Zuo, Hengjian Cui, and Xuming He. On the Stahel– Donoho estimator and depth-weighted means of multivariate data. Annals of Statistics, 32(1):167–188, February 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1079120132.

Zuo:2004:IFM

[ZCY04] Yijun Zuo, Hengjian Cui, and Dennis Young. Influence function and maximum bias of projection depth based estimators. Annals of Statistics, 32(1):189–218, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1079120133.

Zeng:2004:EMS

[Zen04] Donglin Zeng. Estimating marginal survival function by adjusting for dependent censoring using many covariates. Annals of Statistics, 32(4):1533–1555, August 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1091626178.

Zeng:2005:LAM

[Zen05] Donglin Zeng. Likelihood approach for marginal proportional hazards regression in the presence of dependent censoring. Annals of Statistics, 33(2):501–521, April 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1117114326.

Zhang:2005:MFB

[ZFLS05] Aijun Zhang, Kai-Tai Fang, Runze Li, and Agus Sudjianto. Majorization framework for balanced lattice designs. Annals of Statistics, 33(6):2837–2853, December 2005. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1140191675.

Zhang:2009:SMC

[ZFS09] Wenyang Zhang, Jianqing Fan, and Yan Sun. A semiparametric model for cluster data. Annals of Statistics, 37(5A):2377–2408, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1247663759.

Zhou:2005:MET

[ZH05] Harrison H. Zhou and J. T. Gene Hwang. Minimax estimation with thresholding and its application to wavelet analysis. Annals of Statistics, 33(1):101–125, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1112967700.

Zuo:2006:LDM

[ZH06] Yijun Zuo and Xuming He. On the limiting distributions of multivariate depth-based rank sum statistics and related tests. Annals of Statistics, 34(6):2879–2896, December 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1179935068.

Zhang:2008:SBL

[ZH08a] Cun-Hui Zhang and Jian Huang. The sparsity and bias of the Lasso selection in high-dimensional linear regression. Annals of Statistics, 36(4):1567–1594, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237292.

Zhou:2008:DRB

[ZH08b] Jianhui Zhou and Xuming He. Dimension reduction based on constrained canonical correlation and variable filtering. Annals of Statistics, 36(4):1649–1668, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1216237295.

Zhao:2000:BAS

[Zha00] Linda H. Zhao. Bayesian aspects of some nonparametric problems. Annals of Statistics, 28(2):532–552, April 2000. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1016218229.

Zhang:2002:RBI

[Zha02] Cun-Hui Zhang. Risk bounds in isotonic regression. Annals of Statistics, 30(2):528–555, April 2002. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1021379864.

Zhang:2003:CDT

[Zha03] Cun-Hui Zhang. Compound decision theory and empirical Bayes methods: invited paper. Annals of Statistics, 31(2):379–390, April 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1051027872.

Zhang:2004:SBC

[Zha04] Tong Zhang. Statistical behavior and consistency of classification methods based on convex risk minimization. Annals of Statistics, 32(1):56–85, February 2004. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1079120130.

Zhang:2005:ESR

[Zha05a] Cun-Hui Zhang. Estimation of sums of random variables: Examples and information bounds. Annals of Statistics, 33(5): 2022–2041, October 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1132936555.

Zhang:2005:GEB

[Zha05b] Cun-Hui Zhang. General empirical Bayes wavelet methods and exactly adaptive minimax estimation. Annals of Statistics, 33 (1):54–100, February 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1112967699.

Zhang:2006:SEK

[Zha06] Tong Zhang. From ϵ-entropy to KL-entropy: Analysis of minimum information complexity density estimation. Annals of Statistics, 34(5):2180–2210, October 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1169571794.

Zhang:2008:DOS

[Zha08a] Cun-Hui Zhang. Discussion: One-step sparse estimates in nonconcave penalized likelihood models. Annals of Statistics, 36(4): 1553–1560, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237290.

Zhang:2008:QCS

[Zha08b] Zhengjun Zhang. Quotient correlation: A sample based alternative to Pearson’s correlation. Annals of Statistics, 36(2): 1007–1030, April 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1205420527.

Zhang:2009:SSP

[Zha09] Tong Zhang. Some sharp performance bounds for least squares regression with L_{1} regularization. Annals of Statistics, 37(5A): 2109–2144, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247663750.

Zhang:2007:APC

[ZHCC07] Li-Xin Zhang, Feifang Hu, Siu Hung Cheung, and Wai Sum Chan. Asymptotic properties of covariate-adjusted responseadaptive designs. Annals of Statistics, 35(3):1166–1182, July 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1185304002.

Zou:2007:DFL

[ZHT07] Hui Zou, Trevor Hastie, and Robert Tibshirani. On the “degrees of freedom” of the lasso. Annals of Statistics, 35(5):2173–2192, October 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1194461726.

Zhu:2003:SFA

[Zhu03] Yu Zhu. Structure function for aliasing patterns in 2^{l−n} design with multiple groups of factors. Annals of Statistics, 31 (3):995–1011, June 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1056562471.

Zhu:2004:FBA

[Zhu04] Mu Zhu. On the forward and backward algorithms of projection pursuit. Annals of Statistics, 32(1):233–244, February 2004. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1079120135.

Zhu:2007:PSI

[ZILZ07] Hongtu Zhu, Joseph G. Ibrahim, Sikyum Lee, and Heping Zhang. Perturbation selection and influence measures in local influence analysis. Annals of Statistics, 35(6):2565–2588, December 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1201012972.

Zou:2008:OSS

[ZL08a] Hui Zou and Runze Li. One-step sparse estimates in nonconcave penalized likelihood models. Annals of Statistics, 36(4): 1509–1533, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237287.

Zou:2008:ROS

[ZL08b] Hui Zou and Runze Li. Rejoinder: One-step sparse estimates in nonconcave penalized likelihood models. Annals of Statistics, 36 (4):1561–1566, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237291.

Zhou:2009:SIS

[ZL09] Yong Zhou and Hua Liang. Statistical inference for semiparametric varying-coefficient partially linear models with error-prone linear covariates. Annals of Statistics, 37(1):427–458, February 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1232115941.

Zhao:2009:CAP

[ZRY09] Peng Zhao, Guilherme Rocha, and Bin Yu. The composite absolute penalties family for grouped and hierarchical variable selection. Annals of Statistics, 37(6A):3468–3497, December 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1250515393.

Zhang:2022:CNS

[ZSW22] Yi Zhang, Xiaofeng Shao, and Weibiao Wu. Correction note: Asymptotic spectral theory for nonlinear time series. Annals of Statistics, 50(5):3088–3089, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Correction-note-Asymptotic-
spectral-theory-for-nonlinear-time-series/10.1214/22-
AOS2206.full. See [SW07b].

Zuo:2003:PBD

[Zuo03] Yijun Zuo. Projection-based depth functions and associated medians. Annals of Statistics, 31(5):1460–1490, October 2003. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1065705115.

Zuo:2006:MTB

[Zuo06] Yijun Zuo. Multidimensional trimming based on projection depth. Annals of Statistics, 34(5):2211–2251, October 2006. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1169571795.

Zhang:2003:WTE

[ZW03] Shuanglin Zhang and Man-Yu Wong. Wavelet threshold estimation for additive regression models. Annals of Statistics, 31 (1):152–173, February 2003. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1046294460.

Zhao:2008:CBN

[ZW08] Zhibiao Zhao and Wei Biao Wu. Confidence bands in nonparametric time series regression. Annals of Statistics, 36(4): 1854–1878, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237302.

Zhou:2009:LLQ

[ZW09] Zhou Zhou and Wei Biao Wu. Local linear quantile estimation for nonstationary time series. Annals of Statistics, 37(5B): 2696–2729, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1247836666.

Zhang:2005:BES

[ZY05] Tong Zhang and Bin Yu. Boosting with early stopping: Convergence and consistency. Annals of Statistics, 33(4):1538–1579, August 2005. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1123250222.

Zhang:2008:SDS

[ZY08a] Chunming Zhang and Tao Yu. Semiparametric detection of significant activation for brain fMRI. Annals of Statistics, 36(4): 1693–1725, August 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1216237297.

Zou:2008:CQR

[ZY08b] Hui Zou and Ming Yuan. Composite quantile regression and the oracle model selection theory. Annals of Statistics, 36(3): 1108–1126, June 2008. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1211819558.

Zhu:2006:GST

[ZZ06] Hongtu Zhu and Heping Zhang. Generalized score test of homogeneity for mixed effects models. Annals of Statistics, 34 (3):1545–1569, June 2006. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1152540758.

Zhang:2009:ANN

[ZZ09a] Cun-Hui Zhang and Zhiyi Zhang. Asymptotic normality of a nonparametric estimator of sample coverage. Annals of Statistics, 37(5A):2582–2595, October 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1247663767.

Zou:2009:AEN

[ZZ09b] Hui Zou and Hao Helen Zhang. On the adaptive elastic-net with a diverging number of parameters. Annals of Statistics, 37(4): 1733–1751, August 2009. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1245332831.